### v1

# Data-Driven Chance-Constrained Multiple-Choice Knapsack Problem: Model, Algorithms, and Applications

Abstract
The multiple-choice knapsack problem (MCKP) is a classic NP-hard combinatorial optimization problem. Motivated by several significant practical applications, this work investigates a novel variant of MCKP called data-driven chance-constrained multiple-choice knapsack problem (DDCCMCKP), where the item weight is a random variable with unknown probability distribution. We first present the problem formulation of DDCCMCKP, and then establish two benchmark sets. The first set contains synthetic instances, and the second set is devised to simulate a real-world application scenario of a certain telecommunication company. To solve DDCCMCKP, we propose a data-driven adaptive local search (DDALS) algorithm. The main merit of DDALS lies in evaluating solutions with chance constraints by data-driven methods, under the condition of unknown distributions and only historical sample data being available. The experimental results demonstrate the effectiveness of the proposed algorithm and show that it is superior to other baselines. Additionally, ablation experiments confirm the necessity of each component in the algorithm. Our proposed algorithm can serve as the baseline for future research, and the code and benchmark sets will be open-sourced to further promote research on this challenging problem.
Index Terms:
Combinatorial optimization, data-driven approach, multiple-choice knapsack problem, evolutionary algorithm.
I Introduction
The multiple-choice knapsack problem (MCKP) is a variant of the classic knapsack problem [1]. MCKP finds practical applications in various areas, such as supply chain optimization [2], target market selection [3], investments portfolio selection [4], transportation programming [5], and component selection in the information technology systems [6]. Specifically, MCKP involves a set of items, each possessing a weight and a value. The items are divided into several disjoint classes and only one item can be selected in each class [1, 7]. The objective of MCKP is to select exactly one item from each class such as to maximize the total value while adhering to a predefined weight limit (i.e., knapsack constraint). Over the past several decades, plenty of methods for solving MCKP have been proposed, including greedy methods [8], branch-and-bound [9], dynamic programming [10], approximation algorithms [11], and heuristic search [12].
Despite of significant research success in addressing the MCKP, those advanced methods are incapable of handling problems in real-world scenarios where item weights are randomly distributed instead of fixed. To handle random variables, a common approach is to model deterministic constraints as chance constraints, i.e., the probability (confidence level) of total weight (random variable) of the selected items not exceeding a given deterministic value is no less than a certain threshold [13, 14]. Furthermore, the specific probability distributions of the random variables could be unknown [15, 16]. Instead, one only has historical sample data to characterize the random variables. The collection of sample data for each item is assumed to be independent and uncorrelated. This variant of MCKP is referred to as data-driven chance-constrained multiple-choice knapsack problem (DDCCMCKP).
A typical application of DDCCMCKP is the configuration of fifth-generation (5G) end-to-end (E2E) network. In the area of 5G for business [17], user plane time delay is one of the key performance indicators for low latency communication [18, 19]. As illustrated in Figure 1, the complete packet transmission path in a 5G cellular network consists of several segments, i.e., eNB, backhaul, core network, Internet/cloud, and the transport between core network and Internet/cloud. More specifically, these segments correspond to classes, where for each class there exist different options (items) and only one option can be selected to implement the class. The total time delay of a packet transmission can be calculated as the sum of the delays induced by all the selected options, i.e., $(T_{eNB}+T_{Backhaul}+T_{Core}+T_{Transport}+T_{Internet})*2$. Moreover, the packet transmission delays of the implementation options are random variables with unknown distributions as they are affected by various factors, including workloads, network configurations, broadcast environment, and random failures of software and hardware. The configuration of 5G E2E network is to select an implementation option for each segment. Different configurations have different total costs and time delays, where the costs are fixed values, but time delays are unknown random variables. The objective is to minimize the total cost of network configuration while ensuring that the probability of meeting the delay is less than a given threshold under a certain value.
Another application of DDCCMCKP is supply chain planning and optimization in modern industries [2, 3]. The supply-chain operations reference (SCOR) model, published by the Supply-Chain Council [20], is a useful tool for the upper management of a firm to design and reconfigure the supply chain to achieve better performance. In SCOR, the E2E supply chain consists of four distinct processes: source, make, deliver, and plan (see Figure 2). Specifically, these processes can be further subdivided into supplier selection, R&D testing, production, transportation, and recycling handling, which correspond to different classes along the supply chain. Therefore, each class has various configuration options that can be influenced by random factors, such as market demand, cost fluctuations, and technology configurations.
Figure 1: 5G E2E network service model [19]
Figure 2: SCOR model-based supply chain infrastructure [20]
Since DDCCMCKP has not been paid research attention, it merits thorough investigation due to its extensive and noteworthy applications. Compared to MCKP, DDCCMCKP faces greater challenges due to the complexity of its constraints. There are two main difficulties: it is a chance-constrained optimization problem, and the distributions of the random variables are unknown. Therefore, a data-driven approach that does not rely on specific distribution information but rather available sample data to evaluate solutions must be proposed.
The main contributions are summarized as follows:
  1. 1.
    From the problem perspective, we define DDCCMCKP from several real-world scenarios and formulate the model. To the best of our knowledge, this is the first endeavor to address this specific problem. To further promote research on this problem, we have designed two sets of benchmarks. The first set is a simulated benchmark with a simple mechanism, while the other set is to simulate a real-world business requirement of a telecommunication company. The generated benchmarks will be made publicly available to promote further research.
  2. 2.
    From the algorithmic perspective, we propose a novel algorithm named data-driven adaptive local Search (DDALS). The first novelty of DDALS is that we evaluate solutions with chance constraints by a data-driven method. The second is that we design surrogate weights using statistics of sample data to ensure the performance of the search algorithm. Furthermore, our algorithm can adaptively jump between small and larger search regions for efficient searching of high-quality solutions. Lastly, we improve the proportion of feasible solutions through carefully designed filtering mechanisms.
  3. 3.
    From the computational perspective, we experimentally confirm the effectiveness of DDALS. We adapt several commonly-used approaches as the baselines, including greedy search (Greedy), genetic algorithm (GA), and estimation of distribution algorithm (EDA). Experimental results show that DDALS outperforms other baseline methods. The proposed algorithm can be served as a baseline for comparison experiments on this challenging problem.
The rest of the paper is organized as follows. In Section II, we present a comprehensive review of related works. Section III provides a formal definition of DDCCMCKP. In Section IV, DDALS is introduced and a detailed explanation of its implementation is provided. In Section V, the benchmarks we designed are described in details. The experimental results and analyses are given in Section VI. Finally, we summarize our findings and provide suggestions for future research in Section VII.
II Related works
In this section, we will review the relevant literature. We first review existing research on the traditional MCKP. The second category of research focuses on chance-constrained optimization. Since the methods developed in robust optimization can solve chance-constrained optimization by constructing proper uncertainty sets [21, 22], we further review robust optimization and distributionally robust optimization. While robust optimization approach minimizes the worst-case scenario under the uncertainty of random variables, the goal of distributionally robust optimization is to ensure the robustness of the optimal solution in the distribution-unknown situations.
II-A Algorithms for MCKP
There have been various stochastic variations of the classical 0-1 MCKP that have been extensively studied in the literature. For instance, Dyer et al.[9] proposed a branch and bound method to solve the MCKP. Pisinger [23] presented a minimal algorithm including linear relaxation, greedy algorithm and dynamic programming. Lamanna et al. [24] introduced a new algorithm for solving the multidimensional MCKP based on Kernel search, which divided the solution process of the algorithm into two phases with complementary objectives. Gens and Levner [11] proposed a fast approximation algorithm for the MCKP with a 4/5 approximation ratio of the optimal solution using binary search. He et al. [25] developed an improved approximate binary search algorithm with an approximation ratio of $3+(1/2)^{t}$. Hifi et al. [12] proposed a reactive local search algorithm for the multidimensional MCKP. Mkaouar et al. [26] introduced a method based on the artificial bee colony algorithm to solve the multidimensional MCKP. In general, existing researches on MCKP cannot be applied, since the solutions in DDCCMCKP can not be evaluated by deterministic approaches.
II-B Chance-Constrained Optimization
Chance-constrained optimization plays a critical role in stochastic optimization [27, 13, 14]. Kleinberg et al. [28] focused on packing multiple bursty connections with varying transmission rates on a link and considered stochastic knapsack problems with known Bernoulli distribution. Goel and Indyk [29] studied the knapsack problem when the jobs had stochastic processing requirements or sizes. They proposed polynomial time approximation scheme (PTAS) when the jobs were all known to follow either Poisson or exponential distributions. Goyal and Ravi[30] proposed a PTAS that used a parametric linear programming reformulation and yielded near-optimal profit efficiently. The random variables were assumed to follow a known Gaussian distribution. Shabtai et al. [31] concentrated on the chance-constrained knapsack problem with known normal distributions. They presented a fast PTAS for the problem that run in poly($n/\epsilon$) time. Zhang et al. [32] studied trajectory prediction with chance constraint, assuming that future position of the moving obstacle could be represented as Gaussian distributions based on the predicted statistics. Chai et al. [33] studied trajectory optimization with chance constraint with known probability density function of random variables defined on the measurable set. Yang and Chakraborty [34] exploited the geometry of the non-convex feasible region of the chance-constrained knapsack problem in the variance-mean plane and transformed the problem into a sequence of deterministic knapsack problems to find an optimal solution. All these works mainly focused on theoretical analysis of situations that the distribution of random variables is known, which are not applicable to our problem as well.
II-C Robust Optimization
A robust optimization approach aims to construct an uncertainty set $\mathcal{U}_{\epsilon}\subseteq\mathcal{U}$ such that a feasible solution of the form $g(x,\xi)\leq 0,\forall\xi\in\mathcal{U}_{\epsilon}$ will also be feasible with probability at least $1-\epsilon$ with respect to $P$, i.e., $\mathcal{P}(g(x,\xi)\leq 0)\geq 1-\epsilon,\forall\xi\in\mathcal{U}$, which has the same form with chance constraint. Han [21] considered a chance-constrained knapsack problem where each item had an independent normally distributed random weight with a known mean and variance, and proposed an efficient pseudo-polynomial time algorithm based on the robust optimization approach. Joung and Lee[22] proposed a robust optimization-based heuristic algorithm where the optimal solution could be obtained by solving the ordinary binary knapsack problem iteratively.
Distributionally robust optimization was proposed as an improvement of robust optimization when the distribution of the random variable was unknown and only observable through a finite training dataset. Mohajerin Esfahani and Kuhn [35] showed that distributionally robust optimization problems over Wasserstein balls could be reformulated as finite convex programs. Ji and Lejeune [15] studied distributionally robust chance-constrained programming optimization problems with data-driven Wasserstein ambiguity sets. The proposed algorithmic and reformulation framework were applied to all types of distributionally robust chance-constrained optimization problems, including the multidimensional knapsack problem. Additionally, Xie [16] proposed that the distributionally robust chance-constrained problem could be reformulated as a conditional value-at-risk constrained optimization problem. The problem with binary decision variables admitted a big-M free formulation and could be solved by a branch and cut algorithm. Gao and Kelywegt [36] derived a strong duality reformulation of the corresponding distributionally robust stochastic optimization problem and construct approximate worst-case distributions explicitly. Note that most of these work can not be applied. We have tried to make modifications to Ji’s [15] and Yang’s [34] algorithm to make it applicable to our problem, but the results show that both of them exhibited unsatisfactory performance. (see Section VI-F for details).
III Notations and problem definition
We give the formulation of DDCCMCKP in Eq. III. Consider $m$ mutually disjoint classes $N_{1},\ldots,N_{m}$ of items. Each item $j\in N_{i}$ has an associated cost $c_{ij}$ and weight $w_{ij}$, where $c_{ij},j\in N_{i},\forall i\in\mathcal{M}=\{1,\ldots,m\}$ are fixed values while $w_{ij},j\in N_{i},\forall i\in\mathcal{M}$ are unknown random variables. The weight of item $j\in N_{i}$ is characterized by $L$ sample data, denoted as $d_{ijl},i\in\{1,\ldots,m\},j\in N_{i},l\in\{1,\ldots,L\}$. Note that the sample data of each item is independent and identically distributed (i.i.d.). The goal is to select one item from each class to minimize the total cost, subject to keeping the total weight below the knapsack’s capacity $W$ with probability higher than the given confidence level $P_{0}$. By introducing the binary variables $x_{ij}$ which take on value 1 if the item $j$ is chosen in class $N_{i}$, the problem can be formulated as:
$\min\quad$ $\sum\limits_{i=1}^{m}\sum\limits_{j\in N_{i}}c_{ij}x_{ij}$ (1a)
s.t. $\mathcal{P}(\sum\limits_{i=1}^{m}\sum\limits_{j\in N_{i}}w_{ij}x_{ij}\leq W)\geq P_{0}$ (1b)
$\sum\limits_{j\in N_{i}}x_{ij}=1,\forall i\in\mathcal{M}$ (1c)
$x_{ij}\in\{0,1\},\forall i\in\mathcal{M},j\in N_{i}$ (1d)
To avoid trivial solutions, we assume that the following inequality holds:
$\sum\limits_{i=1}^{m}\min_{1\leq l\leq L,1\leq j\leq n_{i}}d_{ijl}\leq W\leq\sum\limits_{i=1}^{m}\max_{1\leq l\leq L,1\leq j\leq n_{i}}d_{ijl}$ (2)
Note that MCKP is NP-hard [1]. Supposing that all samples of each item’s weight $w_{ij}$ are the same, then DDCCMCKP reduces to traditional MCKP. Therefore, MCKP can be seen as a special case of DDCCMCKP, which indicates that DDCCMCKP is also NP-hard. It motivates us to adopt an algorithm based on a heuristic search framework to solve DDCCMCKP.
IV Data-driven adaptive local search
As mentioned in Section II, although there have been many works in the fields of MCKP, chance-constrained optimization and robust optimization, none of the methods could be directly applied to our problem. In order to tackle the DDCCMCKP, we present a novel approach called data-driven adaptive local search (DDALS) in this work. Specifically, DDALS consists of three main parts, confidence level evaluation, surrogate weight design and local search framework. Finally, we will analyze the complexity of the algorithm.
IV-A Confidence Level Evaluation
Since the sample data is assumed to be i.i.d., it is necessary to make full use of the information contained in the sample data, since we have no information about the true distribution. Given a solution $S$, a natural idea is to calculate all possible combinations of the data and count the proportion of these combinations that meet the requirements. This proportion is an estimate of the real confidence level $P$ of the solution. However, obviously, a brute-force method is not applicable since it requires a calculation complexity of $L^{m}$, which is related to the problem size $m$ and sample size $L$. It may be feasible for small-scale problems, but it becomes impractical for large-scale problems. Therefore, based on the above considerations, we propose an exact evaluation method for small-scale problems that is much faster than brute-force enumeration. For large-scale problems, we propose an accelerated Monte Carlo evaluation method.
IV-A1 Exact evaluation
Exact evaluation is based on a simple idea: counting and heaps [37]. Assume that the sample data of items has been already arranged from largest to smallest. Solution $S=[s_{1},s_{2},\ldots,s_{m}]$ is under evaluated. The details of exact evaluation are listed as follows:
  • Maintain a maximum heap queue $Q$ with the length $m(1-P_{0})L^{m}$. Every time $Q$ pops the largest sum in the current heap queue [37];
  • Initialization: push $Sum_{1}=d_{s_{1},1}+\ldots+d_{s_{m},1}$ into $Q$;
  • Iteration: $Q$ pops the largest sum in $D$:
    $D=$ $[d_{s_{1}l_{1}},d_{s_{2}l_{2}},\ldots,d_{s_{m}l_{m}}].$
    Then push neighbors of the popped $D$ whose sum are smaller than that of $D$ into $Q$, listed as follows:
    $[d_{s_{1}l_{1}+1},d_{s_{2}l_{2}},\ldots,d_{s_{m}l_{m}}],$
    $[d_{s_{1}l_{1}},d_{s_{2}l_{2}+1},\ldots,d_{s_{m}l_{m}}],$
    $\ldots$,… ,
    $[d_{s_{1}l_{1}},d_{s_{2}l_{2}},\ldots,d_{s_{m}l_{m}+1}]$
    Note that duplicates are not allowed in $Q$.
  • Repeat until the pop sum of $D$ is smaller than $W$ or the iteration times $k$ reach $\lceil(1-P_{0})L^{m}\rceil$.
We have the following theorem,
Theorem IV.1
The popped sum in the $k^{th}$ iteration is exactly the $k^{th}$ largest sum.
Detailed proofs can be found in Appendix 2. If the ratio of the number to the total number of combinations is greater than $1-P_{0}$, then we determine that the solution is infeasible.
IV-A2 Accelerated Monte Carlo evaluation
Monte Carlo simulation is a method that uses random numbers to solve many computational problems [38]. In this work, we focus on situations where the knapsack constraints $W$ are tight at a high confidence level $P_{0}$, which means feasible solutions account for a very small proportion of the entire solution space of the problem. This observation motivates us that the evaluation procedure can be greatly accelerated if most of the infeasible solutions can be quickly eliminated. Specifically, assuming that the sample data of the items has been sorted from largest to smallest, we can check the $l^{th}$ largest data combination $d_{1l},d_{2l},...,d_{ml}$ for the combination of items to be solved, i.e.,$D_{l}=\sum_{i=1}^{m}d_{il}$, by comparing with the weight threshold $W$ to directly filter out most of the infeasible solutions. Therefore, the evaluations times required during the algorithm operation can be significantly reduced. The process can be summarized as follows.
  • Construct a filtered list according to the minimum prime factorization for $K=L^{m}\times(1-P_{0})$.
  • Only a limited size of combinations will be selected to construct the filtered list $R\ll m!$;
  • Calculate the sum of the data of each item of the screening list $R$, and check whether there is at least one combination satisfying $\sum_{i=1}^{m}d_{il}\geq W$. If not, the solution will be evaluated by Monte Carlo simulation.
IV-B Surrogate Weight Design
It is not feasible to directly obtain the cost-effectiveness of each item in a data-driven scenario, which impedes the algorithm’s ability to search for better solutions. To overcome this challenge, we construct a surrogate cost-effectiveness based on the statistic of sample data. Specifically, we calculate the sample mean and standard deviation of the item weights, $\mu_{ij}=\mathbb{E}(w_{ij}),\sigma^{2}_{ij}=Var(w_{ij}),i\in\{1,...,m\},j\in N_{i}$. Then item weights are represented by the weighted sum of statistical characteristics,
$\tilde{w}_{ij}:=\mu_{ij}+\lambda\sigma_{ij},\forall{i}=1,...,m,j\in N_{i}.$ (3)
where $\lambda$ is the weighted parameter. This approach provides a statistical measure for guiding the search process towards optimal solutions.
IV-C Local Search Framework
Based on the previously data-driven solution evaluation method and surrogate weight design, we now introduce the overall algorithmic framework of DDALS. Local search is a widely used heuristic search framework for efficiently solving various combinatorial optimization problems [12, 39]. Thus, our algorithm adopts the framework of local search algorithm, which can adaptively jump between small and large regions, achieving efficient search for solutions. The block diagram of DDALS shown in Figure 3 . DDALS is composed of the following five components: constructive procedure (CP) for solution initialization, a local swap search (LSS) operator for improving the current solution, a degradation (Degrade) operator for escaping local optima, and a further swap search (FSS) operator for searching in a larger feasible region. Since limited sample data cannot exactly reflect the true distribution, we additionally incorporate a solution feasibility enhancement (SFE) to improve the feasibility of the output.
Figure 3: The block diagram of DDALS
Input: A DDCCMCKP benchmark; Maximum iteration times, $MaxIter$; weighting factor, $\lambda$
Output: the optimal feasible solution $S_{opt}$
1 $S_{cp}\leftarrow ConstructProcedure(\lambda)$, cost list, MC list;
2 while $t\leq MaxIter$ do
3       $S_{t}\leftarrow LocalSwapSearch(S_{t})$;
4       if $S_{t}$ is a better solution then
5            $S^{*}\leftarrow S_{t}$
6      $S_{t}\leftarrow Degrade(S_{t})$;
7       $t=t+1$;
8
[…]
10 Update $S_{opt}$ through solution feasibility enhancement;
11 Return $S_{opt}$;
Algorithm 1 Data-Driven Adaptive Local Search
Algorithm 1 illustrates the framework of DDALS. The inputs of DDALS include the benchmark, maximum iteration times $MaxIter$, and a user-defined weighting factor $\lambda$. We first initialize the solution $S_{cp}$ using CP (line 1), then iteratively improve the quality of solutions. At iteration $t$, LSS attempts to obtain a better feasible solution by operating on $S_{t}$ (line 3). If LSS finds an improvement, the optimal solution $S^{*}$ is updated (line 4-5). The current solution $S_{t}$ will be degraded by randomly selecting an item for replacement to escape a local optimum (line 6). After iterations end, FSS is employed on $S_{opt}$ to search for a better solution (line 8). The final solution $S_{opt}$ will be output by SFE.
The detailed designs of the components are list as follows:
IV-C1 Constructive procedure
The initialization procedure is based on greedy strategy [7]. Utilizing the item weights defined in Definition 3, we calculate the cost-effective variables $u_{ij}=v_{ij}/\tilde{w}_{ij}$ for each item $j\in N_{i}$, then sort the items in descending order of $u_{ij}$ for each class $N_{i}$. $S_{cp}$ is constructed by selecting the item with the largest $u_{ij}$ in each class. If the current $S_{t}$ is infeasible, the item which has the largest $\tilde{w}_{i_{0}j}$ in $S_{t}$ is replaced by an item with a smaller weight in the same class $i_{0}$. The procedure iterates until a feasible $S_{cp}$ is obtained. If all items in class $i_{0}$ are tried but still infeasible,$S_{cp}$ will be constructed by the smallest $\tilde{w}_{ij}$ in each class.
IV-C2 Local swap search
LSS applies a swapping strategy for picked items to improve the current solution $S_{t}$. If the obtained solution is feasible, each swap is authorized. LSS will try all the items with lower cost in each class $N_{i}$ and selects the item with the lowest cost from all authorized swaps.
IV-C3 Degrade
Degrade randomly replaces an item $j_{i_{0}}$ of class $i_{0}$ in the current solution $S_{t}$. If an infeasible solution is generated, it randomly selects a lighter item in $i_{0}$. If there is no suitable item in the $i_{0}$, it randomly switches to another class and tries again until a new feasible solution $S_{t}$ is obtained.
IV-C4 Further swap search
FSS iterates through all class combinations $(i_{0},i_{1})$ of the current optimal solution $S_{t}$ to find a better and feasible solution. If it finds a pair of items with a smaller cost, $S_{t}$ will be updated. The algorithm ends when all pairs of items have been searched.
IV-C5 Solution feasibility enhancement
In practice, due to the demand for high confidence level $P_{0}$, we tend to minimize the cost of the solution while ensuring its feasibility, rather than blindly pursuing the minimization of the cost, leading to solutions that do not satisfy constraints, which renders them useless in practice. It is a crucial requirement in many practical applications. However, it is difficult to guarantee the real feasibility of the optimal solution due to limited sample size. Note that the sample size to achieve smaller errors for arbitrarily distributed random variables will be very large (See Appendix 3 for theoretical analysis between estimation error and sample size). Therefore, we design three filtering variants to enhance the feasibility of output solution, while not increasing the cost too much. In the iterations of DDALS algorithm, two list of feasible solutions are recorded, each with a capacity of 30. The first records 30 feasible solutions with the lowest cost encountered so far, named Cost List. The second records 30 solutions with the largest confidence level encountered so far, named MC List.
  • Original: Output the optimal solution of FSS.
  • Variant 1: The minimum-cost solution with a higher confidence level than $P_{0}$. For example, it should satisfy $P\geq 0.995$ if $P_{0}=0.99$.
  • Variant 2: All solutions will be scored according to the order in the list they are in. Each candidate solution has two scores, which will be added together. Choose the solution with the highest score as the final output.
  • Variant 3: Take the first 10 solution (if the amount is larger than 10) in Cost List as the final output and choose one manually.
IV-D Complexity Analysis
IV-D1 Exact evaluation
First, if the solution can be directly determined infeasible in the fast screening phase, the time complexity takes $O(1)$. Otherwise, in order to verify whether it is feasible, the maximum number of iterations that the algorithm needs to perform is $T=L^{m}(1-P_{0})-l_{min}^{m}$, where $l_{min}$ is the index that the combination of $l_{min}$ samples of solution is larger than $T_{\max}$. The maximum heap size is $k_{0}+(m-1)T=\sum_{i=1}^{m}\binom{m}{i}(l_{min}+1)^{m-i}+(m-1)L^{m}(1-P_{0})-(m-1)l_{min}^{m}=O(mT)$. Since at each iteration it needs to sort the heap once, the total number of times is $O(\log(mT))=O(\log m+\log T)=O(\log T),m\ll L$. Thus, the total time complexity of the filtered algorithm is $O(mL\log L+\sum_{i=1}^{m}\binom{m}{i}(l_{min}+1)^{m-i}+T\log T)=O(\max\{T\log T,ml_{min}^{m-1},m^{2}l_{min}^{m-2},...,ml_{min}\})$.
IV-D2 Accelerated Monte Carlo evaluation
The complexity of the accelerated Monte Carlo evaluation is $O((|G|+\alpha\times MonteCarlo)\times Eval)$, where $MonteCarlo$ denotes the set number of Monte Carlo simulations and $|G|$ denotes the size of fast screening combination. If the solution is an infeasible solution determined by fast screening, $\alpha=0$, otherwise $\alpha=1$.
IV-D3 DDALS
  • CP includes sorting and solution evaluation. Up to $m\times N$ times of evaluation are required. The complexity is $O(Eval\times mN)$, where $N=\max|N_{i}|$.
  • Similar to CP, LSS require up to $m\times N$ times of evaluation to search a possible better solution. The complexity is $O(Eval\times mN)$.
  • The worst case is that only one combination of items is feasible in the whole solution space. The complexity of Degrade is $O(Eval\times mN)$.
  • FSS iterates through a maximum of combinations of classes $\binom{m}{2}\times N^{2}$. The complexity of FSS is $O(Eval\times m(m-1)N^{2}/2)=O(Eval\times m^{2}N^{2})$.
Note that SFH only takes $O(1)$. Therefore, the total time complexity of DDALS is $O(Eval\times mN+MaxIter\times(Eval\times mN+Eval\times mN)+Eval\times N^{2}m^{2}+1)=O(Eval\times\max\{N^{2}m^{2},MaxIter\times mN\})$.
V Benchmark Design
Since DDCCMCKP is a new problem that has not been solved before, it is necessary to design benchmarks to evaluate the performance of algorithms in different scenarios. Therefore, we designed two sets of benchmarks to cover a wide range of problem sizes and complexities, each of which consists of multiple benchmarks of different scales, in order to comprehensively evaluate the algorithm’s performance in different scenarios.
(a) LAB
(b) APP
Figure 4: Probability density function of items. The horizontal axis represents the time delay and the vertical axis represents the probability density.
The first set of benchmarks aims to simulate the case where the delay of items follows different common continuous distributions, named LAB. The distribution of each item is randomly chosen from the set of uniform, truncated normal, fatigue-life, bimodal and gamma distributions. The mean of items is uniformly distributed in $U(2,8)$, i.e., $\mu\sim U(2,8)$, and the variance is uniformly distributed in $U(1,19)$, i.e., $\sigma^{2}\sim U(1,19)$, where $U(2,8)$ denotes a uniform distribution over the interval [2,8]. In the case of the gamma distribution, the mean value is $\mu\sim U(0.5,2.5)$ and variance $\sigma^{2}\sim U(0.05,0.625)$. The cost, representing the value of taking the item, is generated randomly between $U(1,10)$, corresponding to uncorrelated situation in classic MCKP benchmark[26], which is a widely used benchmark in the literature. The curve of probability density function of different distributions of the items involved in LAB is shown in Figure.4a.
The second set of benchmarks simulates the actual business requirements of a telecommunication company in the 5G E2E network, denoted as APP. In APP, the distribution of each item is also randomly chosen in the set of uniform, truncated normal, fatigue-life, bimodal and gamma distribution, which is similar to the LAB case. However, in the 5G E2E network, each data package transmission is restricted to a time window with a fixed size, indicating that packets must be transmitted successfully within a set interval. The difference between APP and LAB lies in the introduction of the retransmission mechanism [17]. This mechanism implies that each item in all classes has a probability of being retransmitted when a transmission failure occurs. In the APP, the first data package transmission attempt must be completed within $(0,10]$ ms with a success probability of 0.9. If the first attempt fails, a second attempt occurs with a delay of $(10,20]$ ms and a success probability of 0.9. Therefore, the delay of one transmission ranges from $(10,20]$ ms with a probability of $0.1\times 0.9=0.09$. The same applies to the third and fourth retransmissions. The transmission fails if an item requires transmitting data more than four times, which limits the maximum number of retransmissions. Consequently, the time delay for an item to transmit a data package belongs to the intervals $(0,10],(10,20],(20,30],(30,40]$ with probabilities of $0.9,0.09,0.009,0.001$, respectively. The distribution of items involved in the APP is presented in Figure 4b. The cost generation formula for items is $c_{ij}=10/(\mu_{ij}+\sigma_{ij})\times U(0.8,1.2)$. The purpose is to make the cost-effectiveness decrease as the delay decreases, thus simulating the marginal diminishing effect of cost-effectiveness in practical application.
For each benchmark in the LAB and APP, we design small-scale and large-scale problems with different parameter settings, as shown in Table I and II respectively. The scale of benchmarks, i.e., $m$ and $N$, is set similar to [26]. The sample sizes $L$ are set 30 for small-scale benchmarks and 500 for large-scale benchmarks. In Table I, ”{11,14}” of $T_{\max}^{LAB}$ means that $T_{\max}$ is set 11 or 14 as two different settings in the experiments of ss1.
TABLE I: Small-scale benchmarks of DDCCMCKP
Benchmark m N L $T_{\max}^{LAB}$ $T_{\max}^{APP}$
$ss1$ 3 5 30 {11,14} {21,27}
$ss2$ 4 5 30 {18,26} {47,49}
$ss3$ 5 5 30 {11,20} {27,38}
$ss4$ 5 10 30 {10,16} {16,27}
TABLE II: Large-scale benchmarks of DDCCMCKP
Benchmark m N L $T_{\max}^{LAB}$ $T_{\max}^{APP}$
$ls1$ 10 10 500 {19,23} {32,37}
$ls2$ 10 20 500 {12,15} {13,16}
$ls3$ 20 10 500 {25,32} {43,48}
$ls4$ 30 10 500 {43,52} {58,70}
$ls5$ 40 10 500 {55,63} {85,95}
$ls6$ 50 10 500 {63,75} {91,100}
VI Computational Study
Experiments are carefully designed to evaluate the performance of DDALS in solving DDCCMCKP. We also conducted ablation studies, which involves disabling individual components of DDALS, to assess the contribution of the components integrated in DDALS. The source code of algorithms and the benchmark sets, are available at Github  [^1] Github: github.com/eddylxf23/DDCCMCKP. All experiments were run in python on an Intel(R) Xeon(R) CPU E5-2699A v4(2.40GHz,252G RAM) with Ubuntu 16.04 LST.
VI-A Algorithm Settings
We set different parameter combinations of weighted parameter $\lambda$ and iteration times $t$ to evaluate the performance on small-scale and large-scale DDCCMCKP respectively. The performance is evaluated by the cost of final solution when the algorithm reaches the maximum of iteration, and the experiment was repeated for 10 times and take the average value. The value of $\lambda$ is tested from 1 to 10. As show in Fig. 5, there are several points worth noting. In the experiments of LAB-ss2, when $T_{\max}=18$, the current solution of the algorithm remains unchanged. The results show that the lowest cost was obtained when $\lambda=0$. However, when we validate the output solution with real confidence level, we find that the solution was infeasible, which means that the algorithm could not find feasible solutions based on the samples without considering the standard deviation. When $T_{\max}=26$, different lambdas do not affect the algorithm’s convergence to the lowest cost solution. In the experiments of LAB-ls2, when $T_{\max}=25$, the algorithm converges very slowly when we set $\lambda=0$, while other settings can converge quickly. However, when $T_{\max}=32$, different lambda values have little effect on the algorithm. That’s because the constraints are relatively loose, and feasible solutions account for a large proportion of the entire solution space, making the effect of standard deviation less significant. In general, it is necessary to set a non-zero value for $\lambda$, and our algorithm is robust to different $\lambda$, with $t=30$ being appropriate.
(a) LAB-ss2, $T_{\max}=18$, exact
(b) LAB-ss2, $T_{\max}=26$, exact
(c) LAB-ls3, $T_{\max}=25$, MC
(d) LAB-ls3, $T_{\max}=32$, MC
Figure 5: Performance analysis of different $\lambda$. As the algorithm iterates, lower cost can be achieved. The results show that the surrogate weight is necessary for DDALS while a non-zero $\lambda$ is necessary to the algorithm.
VI-B Performance of Evaluation Methods
Since each candidate solution can be seen as an unknown random distribution, concentration inequalities can serve as a reference to evaluate the quality of these candidate solutions. We utilize two kinds of concentration inequality as the lower bounds: Bernstein’s Inequality and $K$-order Hoeffding’s Inequality, which have been widely used in the literature for evaluating the performance of optimization algorithms (Please refer to Appendix 1 for detailed formulas).
TABLE III: Evaluation for LAB-ss2
[…]
Real 68.24 89.41 99.51 99.98 100.00
B’s inequality 23.22 49.16 75.63 89.87 95.85
H’s inequality 22.83 42.20 62.32 74.17 85.12
Exact 81.64 93.90 99.53 99.99 100.00
We test exact method on LAB-ss2, and calculate the real confidence level for solution $S_{1},S_{2},S_{3},S_{4},S_{5}$ through the original generator, denoted as ‘Real’. ‘B’s’ and ‘H’s’ represent Bernstein’s and Hoeffding’s inequality, respectively. From Table III, we found a substantial gap between the evaluation outcome and the real confidence level, indicating that the real distribution of solution of DDCCMCKP is far different from Gaussian distribution. Results show that the estimate of the exact method is closest to Real, and the estimation error becomes smaller as the real confidence level increases.
TABLE IV: Evaluation for APP-ls1
[…]
Real 79.18 90.70 94.48 98.78 98.81
B’s inequality 9.56 11.56 19.82 28.09 48.87
H’s inequality 12.58 25.38 38.30 41.98 53.96
MC(10e2) 81.1$\pm$ 3.78 90.7$\pm$ 2.87 95.8$\pm$ 2.15 99.0$\pm$ 0.82 98.3$\pm$ 1.16
MC(10e3) 79.61$\pm$ 1.53 90.28$\pm$ 0.67 95.14$\pm$ 0.60 99.22$\pm$ 0.25 98.24$\pm$ 0.30
MC(10e4) 79.12$\pm$ 0.53 90.27$\pm$ 0.27 95.13$\pm$ 0.60 99.24$\pm$ 0.09 98.44$\pm$ 0.13
MC(10e5) 79.07$\pm$ 0.17 90.38$\pm$ 0.02 95.13$\pm$ 0.05 99.20$\pm$ 0.03 98.40$\pm$ 0.04
MC(10e6) 79.13$\pm$ 0.05 90.38$\pm$ 0.02 95.10$\pm$ 0.03 99.21$\pm$ 0.00 99.41$\pm$ 0.01
Next, we test Monte Carlo evaluation on APP-ls1 in Table IV. ‘MC(10e2)’ represents that the evaluation is carried by 10e2 times of Monte Carlo estimation. The result is the average of 10 times’ results. ‘81.1 $\pm$ 3.78’ represents that the mean of estimated confidence level is 81.1% and the standard deviation is 3.78%. The result shows that the estimation error of $210e$ 2 times has the largest fluctuations and $510e$ 5 tends to be close to the real confidence value, but still not enough to meet the requirement for larger scale of problem and higher confidence level. Considering the trade-off between accuracy and speed, we set $610e$ 6 as the Monte Carlo simulation times.
VI-C Results on Small-Scale and Large-Scale DDCCMCKP
To evaluate the effectiveness of DDALS, we compare DDALS with several baseline methods on various benchmark sets from the real-world application. The baseline methods include Greedy, GA and EDA. Greedy typically works well on problems where the locally optimal choice leads to an overall optimal solution. GA performs well in solving different optimization problems [40]. While EDA achieves great success in both combinatorial and continuous optimization [41].
For fairness, we first run DDALS then record the value of ET. Next, we let the Greedy, GA, and EDA keep running until convergence or the ET of which exceeds that of DDALS, then output the current solution as the result. The population size is set to 10 and the number of dominant individuals is 6 for both GA and EDA. For GA, the crossover probability of parents is set to 0.1 and the probability of variation is $1/m$. The real confidence level is calculated by using the original generators to generate $710e$ 7 data. Note that a solution is regarded as feasible when its real confidence level is larger than $0.985$, which means that we allow for an error margin of $0.5\%$.
TABLE V: Performance on benchmarks of LAB. The highest average real confidence level and the lowest cost are highlighted through bolding and being underlined.
Benchmark small $T_{\max}$ large $T_{\max}$
C ET ECL RCL FSR C ET ECL RCL FSR
ss3-Greedy 24.4$\pm$0.0 2.0$\pm$0.0 100$\pm$0 99.9$\pm$0.0 1 21.8$\pm$0.0 3.0$\pm$0.0 100$\pm$0 99.9$\pm$0.0 1
ss3-GA 22.6$\pm$1.5 303.4$\pm$2.8 99.9$\pm$0.0 99.7$\pm$0.2 1 12.9$\pm$0.7 262.4$\pm$1.6 99.9$\pm$0.0 99.9$\pm$0.1 1
ss3-EDA 24.1$\pm$2.3 304.1$\pm$2.3 99.9$\pm$0.1 99.6$\pm$0.6 0.9 14.0$\pm$2.0 264.5$\pm$2.9 99.9$\pm$0.0 99.9$\pm$0.1 1
ss3-DDALS(O) 21.4$\pm$0.0 291.9$\pm$6.2 99.9$\pm$0 99.5$\pm$0.0 1 12.6$\pm$0.0 254.4$\pm$4.7 100$\pm$0 99.9$\pm$0 1
ss3-DDALS(V1) 21.4$\pm$0.0 291.9$\pm$6.2 99.9$\pm$0 99.5$\pm$0.0 1 12.6$\pm$0.0 254.4$\pm$4.7 100$\pm$0 99.9$\pm$0 1
ss3-DDALS(V2) 21.4$\pm$0.0 291.9$\pm$6.2 99.9$\pm$0 99.5$\pm$0.0 1 12.6$\pm$0.0 254.4$\pm$4.7 100$\pm$0 99.9$\pm$0 1
ss3-DDALS(V3) 23.9$\pm$1.7 279 99.8$\pm$0.2 99.1$\pm$0.8 0.8 15.7$\pm$1.9 260 99.9$\pm$0.3 99.6$\pm$0.9 0.9
ss4-Greedy 30.1$\pm$0.0 2.0$\pm$0.0 99.9$\pm$0.0 99.9$\pm$0.0 1 27.5$\pm$0.0 3.0$\pm$0.0 99.7$\pm$0.0 99.4$\pm$0.0 1
ss4-GA 23.8$\pm$5.5 768.9$\pm$8.5 99.9$\pm$0.3 99.7$\pm$0.4 1 17.2$\pm$3.0 562.6$\pm$5.7 99.7$\pm$0.3 99.6$\pm$0.3 1
ss4-EDA 17.4$\pm$0.0 763.4$\pm$2.5 99.9$\pm$0.0 99.9$\pm$0.0 1 13.0$\pm$1.0 553.4$\pm$2.7 99.8$\pm$0.2 98.9$\pm$0.4 1
ss4-DDALS(O) 17.4$\pm$0.0 759.9$\pm$25.8 99.9$\pm$0.0 99.9$\pm$0.0 1 11.2$\pm$0.0 525.3$\pm$39.1 99.2$\pm$0.0 99.0$\pm$0.0 1
ss4-DDALS(V1) 17.4$\pm$0.0 759.9$\pm$25.8 99.9$\pm$0.0 99.9$\pm$0.0 1 12.1$\pm$0.0 525.3$\pm$39.1 99.5$\pm$0.0 98.9$\pm$0.0 1
ss4-DDALS(V2) 17.4$\pm$0.0 759.9$\pm$25.8 99.9$\pm$0.0 99.9$\pm$0.0 1 14.0$\pm$1.2 525.3$\pm$39.1 99.9$\pm$0.2 99.0$\pm$0.3 1
ss4-DDALS(V3) 24.6$\pm$4.0 735 99.8$\pm$0.2 99.5$\pm$0.5 1 13.8$\pm$1.3 561 99.8$\pm$0.3 99.5$\pm$0.5 1
ls4-Greedy 169.7$\pm$0.0 5$\pm$0.0 99.8$\pm$0.0 99.8$\pm$0.0 1 162.5$\pm$0.0 6.0$\pm$0.0 99.9$\pm$0.0 99.9$\pm$0.0 1
ls4-GA 123.1$\pm$8.8 11307.2$\pm$8.0 99.1$\pm$0.1 99.1$\pm$0.2 1 104.1$\pm$5.1 8409.1$\pm$5.6 99.1$\pm$0.2 99.2$\pm$0.2 1
ls4-EDA 126.2$\pm$2.3 11304.8$\pm$2.1 99.2$\pm$0.2 99.4$\pm$0.2 1 114.8$\pm$4.5 8403.4$\pm$2.8 99.6$\pm$0.2 99.6$\pm$0.2 1
ls4-DDALS(O) 106.6$\pm$0.8 11289.9$\pm$92.0 99.1$\pm$0.1 99.2$\pm$0.1 1 92.8$\pm$0.7 8376.6$\pm$378.8 99.1$\pm$0.1 99.2$\pm$0.1 1
ls4-DDALS(V1) 109.5$\pm$0.4 11289.9$\pm$92.0 99.7$\pm$0.1 99.7$\pm$0.1 1 97.9$\pm$2.2 8376.6$\pm$378.8 99.6$\pm$0.1 99.7$\pm$0.1 1
ls4-DDALS(V2) 108.9$\pm$1.3 11289.9$\pm$92.0 99.6$\pm$0.2 99.7$\pm$0.2 1 93.8$\pm$1.4 8376.6$\pm$378.8 99.3$\pm$0.2 99.4$\pm$0.2 1
ls4-DDALS(V3) 108.7$\pm$1.1 11171 99.2$\pm$0.2 99.3$\pm$0.2 1 96.5$\pm$1.7 8177 99.1$\pm$0.1 99.2$\pm$0.1 1
The results on LAB benchmarks are presented in Table V. ‘ss3-Greedy’ in the left of table indicates the result of Greedy tested on benchmark ss3. ‘DDALS(O)’, ‘DDALS(V1)’, ‘DDALS(V2)’ and ‘DDALS(V3)’ represents the original, first, second and third version of DDALS. ‘small $T_{\max}$’ and ‘large $T_{\max}$’ are corresponding to small and large $T_{\max}$ in Table I and Table II. Five indicators are utilized to measure the performance. The cost of solution is denoted as ‘C’. The evaluation times ‘ET’ is utilized to compare the convergence speed of algorithms. ‘ECL’ indicates estimated confidence level, while ‘RCL’ is calculated by real generator. ‘FSR’ represents the ratio of actual feasible solution in the output results of the 10 experiments. All results are based on statistical analysis from 10 repetitions of the experiment, denoted as ‘mean $\pm$ standard deviation’. Due to space limitation, parts of results are included in the main text, and the complete results are available in Appendix 5.
TABLE VI: Performance on benchmarks of APP. Due to the introduction of retransmission mechanisms, there is a significant difference in the results between LAB and APP.
Benchmark small $T_{\max}$ large $T_{\max}$
C ET ECL RCL FSR C ET ECL RCL FSR
ss2-Greedy 6.4$\pm$0.0 3$\pm$0.0 99.7$\pm$0.0 98.6$\pm$0.0 1 6.4$\pm$0.0 3.0$\pm$0.0 99.8$\pm$0.0 99.0$\pm$0.0 1
ss2-GA 5.3$\pm$0.0 301.8$\pm$2.2 99.2$\pm$0.0 97.3$\pm$0.4 0 5.3$\pm$0.0 272.6$\pm$2.8 99.5$\pm$0.1 98.1$\pm$0.2 0
ss2-EDA 5.4$\pm$0.2 303.1$\pm$3.0 99.3$\pm$0.2 97.4$\pm$0.7 0.1 5.4$\pm$0.2 274.9$\pm$3.0 99.6$\pm$0.2 98.3$\pm$0.4 0.3
ss2-DDALS(O) 5.2$\pm$0.0 307$\pm$7.6 99.2$\pm$0.0 97.3$\pm$0 0 5.2$\pm$0.0 291$\pm$10.6 99.6$\pm$0.0 98.1$\pm$0.0 0
ss2-DDALS(V1) 5.5$\pm$0.1 307$\pm$7.6 99.8$\pm$0.0 97.5$\pm$0.3 0 5.2$\pm$0 291$\pm$10.1 99.6$\pm$0.0 98.1$\pm$0.0 0
ss2-DDALS(V2) 5.5$\pm$0.1 307$\pm$ 7.6 99.8$\pm$0.0 97.5$\pm$0.3 0 5.4$\pm$0.1 291$\pm$10.1 99.9$\pm$0.1 98.2$\pm$0.1 0
ss2-DDALS(V3) 5.4$\pm$0.1 295 99.4$\pm$0.3 97.7$\pm$0.4 0 5.3$\pm$0.1 272.6$\pm$2.8 99.5$\pm$0.1 98.1$\pm$0.2 0
ss4-Greedy 34.9$\pm$0.0 3.0$\pm$0.0 100.0$\pm$0.0 100.0$\pm$0.0 1 27.2$\pm$0.0 4.0$\pm$0.0 100.0$\pm$0.0 99.0$\pm$0.0 1
ss4-GA 27.3$\pm$1.1 2010$\pm$8.4 99.8$\pm$0.3 90.4$\pm$0.5 0 21.5$\pm$2.9 1614.8$\pm$9.8 99.6$\pm$0.3 93.2$\pm$4.8 0
ss4-EDA 26.1$\pm$0.0 2005.7$\pm$2.3 100$\pm$0 90.1$\pm$0 0 16.4$\pm$0.0 1602.7$\pm$2.1 99.9$\pm$0.0 80.3$\pm$0.0 0
ss4-DDALS(O) 26.1$\pm$0.0 1988.2$\pm$30.4 1$\pm$0 90.1$\pm$0 0 16.7$\pm$0.9 1579.4$\pm$76.8 99.8$\pm$0.2 81.8$\pm$4.4 0
ss4-DDALS(V1) 26.1$\pm$0.0 1988.2$\pm$30.4 1$\pm$0 90.1$\pm$0 0 16.7$\pm$1.0 1579.4$\pm$ 76.8 99.9$\pm$0.0 81.5$\pm$3.6 0
ss4-DDALS(V2) 26.1$\pm$0.0 1988.2$\pm$30.4 1$\pm$0 90.1$\pm$0 0 18.0$\pm$1.9 1579.4$\pm$76.8 99.9$\pm$0.1 86.1$\pm$7.0 0
ss4-DDALS(V3) 29.1$\pm$2.4 2022 99.9$\pm$0.2 91.3$\pm$2.9 0 19.9$\pm$1.7 1589 99.7$\pm$0.3 92.0$\pm$5.3 0
ls2-Greedy 85.7$\pm$0.0 3.0$\pm$0.0 99.9$\pm$0.0 99.9$\pm$0.0 1 85.7$\pm$0.0 3.0$\pm$0.0 100.0$\pm$0.0 100.0$\pm$0.0 1
ls2-GA 76.9$\pm$ 2.7 22617.4$\pm$10.9 99.0$\pm$0.1 98.9$\pm$0.2 1 61.2$\pm$2.1 21327.1$\pm$20.3 99.1$\pm$0.0 99.1$\pm$0.1 1
ls2-EDA 75.9$\pm$0.0 22604.3$\pm$3.3 99.0$\pm$0.0 98.9$\pm$0.0 1 59.2$\pm$0.3 21306.5$\pm$2.5 99.2$\pm$0.0 99.2$\pm$0.1 1
ls2-DDALS(O) 71.9$\pm$0.3 22591.3$\pm$63.6 99.0$\pm$0.0 99.0$\pm$0.0 1 57.2$\pm$0.0 21233.5$\pm$51.9 99.3$\pm$0.1 99.3$\pm$0.1 1
ls2-DDALS(V1) 76.6$\pm$1.1 22591.3$\pm$63.6 99.6$\pm$0.1 99.5$\pm$0.1 1 58.5$\pm$0.8 21233.5$\pm$51.9 99.6$\pm$0.1 99.6$\pm$0.1 1
ls2-DDALS(V2) 72.9$\pm$0.6 22591.3$\pm$63.6 99.3$\pm$0.2 99.2$\pm$0.2 1 57.5$\pm$0.2 21233.5$\pm$51.9 99.4$\pm$0.1 99.5$\pm$0.1 1
ls2-DDALS(V3) 73.1$\pm$0.7 22564 99.2$\pm$0.1 99.1$\pm$0.1 1 57.9$\pm$0.4 21245 99.2$\pm$0.1 99.2$\pm$0.2 1
ls4-Greedy 207.3$\pm$0.0 13.0$\pm$0.0 99.0$\pm$0.0 98.9$\pm$0.0 1 195.3$\pm$0.0 17.0$\pm$0.0 99.4$\pm$0.0 99.4$\pm$0.0 1
ls4-GA 208.5$\pm$8.3 48014.1$\pm$12.1 99.0$\pm$0.0 98.6$\pm$0.3 0.7 175.2$\pm$10.3 45014.6$\pm$10.2 99.0$\pm$0.0 98.5$\pm$0.2 0.6
ls4-EDA 205.3$\pm$2.4 48004.5$\pm$3.3 99.0$\pm$0.0 99.0$\pm$0.0 1 163.8$\pm$1.3 45005.5$\pm$2.6 99.1$\pm$0.0 98.8$\pm$0.0 1
ls4-DDALS(O) 194.9$\pm$0.8 47962.8$\pm$67.7 99.0$\pm$0.0 98.9$\pm$0.0 1 163.3$\pm$3.6 44822.1$\pm$583.5 99.0$\pm$0.0 98.4$\pm$0.1 0.3
ls4-DDALS(V1) 203.2$\pm$2.0 47962.8$\pm$67.7 99.1$\pm$0.0 99.0$\pm$0.1 1 172.8$\pm$4.5 44822.1$\pm$583.5 99.3$\pm$0.1 98.9$\pm$0.3 0.8
ls4-DDALS(V2) 196.5$\pm$1.3 47962.8$\pm$67.7 99.0$\pm$0.0 99.0$\pm$0.0 1 165.3$\pm$4.3 44822.1$\pm$583.5 99.1$\pm$0.1 98.5$\pm$0.2 0.3
ls4-DDALS(V3) 196.9$\pm$2.1 48010 99.0$\pm$0.0 98.9$\pm$0.1 1 168.2$\pm$ 2.3 44174 99.0$\pm$0.0 98.9$\pm$0.1 1
In the case of LAB scenarios, the results of ss3, ss4 and ls4 are shown in Table V. We can intuitively see that the DDALS algorithm outputs the lowest cost solution across all test benchmarks, while Greedy needs only less than 10 times of evaluation but results in the highest cost. The feasibility ratio of all algorithms is close to 1, indicating that our solution evaluation algorithm only needs a very few samples to obtain an evaluation that is close to the true confidence level for small-scale scenarios and 500 samples for large-scale scenarios. From another perspective, for other variations, the original version of DDALS can obtain solutions with lower cost than other variations while guaranteeing a feasible ratio of 1, indicating that the solution evaluation is accurate.
There are significant changes in the scenario of APP, shown in Table V. In small-scale benchmark tests, only the Greedy achieves a feasibility ratio of 1, while the feasibility ratio of other algorithms is almost zero. Although DDALS can obtain lower cost solutions, all of them are found to be infeasible. Such results are practically useless. The main reason is due to the retransmission mechanism in APP, which introduces discontinuities in the item distribution that pose a significant challenge to the solution evaluation algorithm. In a random sampling case, small samples may easily miss information regarding exceptional jumps, leading to performance decline of solution evaluation. Greedy can achieve higher feasibility ratios, but it sacrifices the optimality of the solutions. For the results of large-scale problems, although the difficulty of search increased with the increase of scale, DDALS can still find feasible solutions achieving the lowest cost. Note that in the result of ls4 with large $T_{\max}$, the feasibility ratio of DDALS(O) is 0.3 while that of DDALS(V3) is 1. Given that the average quality of the solutions generated by DDALS(V3) decreases by only 3% relative to the original DDALS, the feasibility ratio increased by 233%. It shows the effectiveness and necessity of filtering the solutions in complex scenarios. Moreover, increasing the sample size $L$ can improve the evaluation performance, which can help capture more information and improve the accuracy of solution evaluation.
VI-D Effectiveness of Components in DDALS
We conduct an ablation study to further assess the effectiveness of four components of DDALS, including CP, LSS, Degrade and FSS. To independently assess their effectiveness, four different variants of DDALS are proposed and tested on benchmarks of LAB:
  1. 1.
    r-CP DDALS: It uses random strategy to select the item in each class instead of utilizing CP. Note that if the initial randomly generated is infeasible solution, query the current solution for each class of item and re-generate randomly in each class of items less than the weight of the current item.
  2. 2.
    no-LSS DDALS: The LSS procedure in the iteration process is removed from DDALS.
  3. 3.
    no-Degrade DDALS: The degrade procedure in the iteration process is removed.
  4. 4.
    no-FSS DDALS: The FSS procedure is removed.
For all the above variants, we compute the performance degradation ratio(PDR) by computing $(C_{2}-C_{1})/C1$ for each benchmark, where $C_{1}$ and $C_{2}$ are the average cost of solution obtained by the variant and original across 10 independent runs. The component being examined is useful for a specific LAB benchmark if the result shows that PDR $>0$. Table VII presents the average PDR averaged on each benchmark of LAB. Based on the results, it can be concluded that none of the examined components has a negative effect on the performance of DDALS. In most cases, these components make LS perform better. Constructing utility for items is helpful for DDALS to obtain a good initial solution on average. LSS provides some kind of exploration, but a negative value on ls3 implies that due to the large size of the sample, the algorithm may still not converge after 30 iterations of running. The operation of degradation has the most significant contribution to the performance of LS, since it greatly improves the problem of the algorithm, as it helps the algorithm to escape local optima. A deeper search is necessary, while FSS has a slightly higher PDR than LSS. However, with the increase of the benchmark’s scale, the contribution of FSS seems to decrease, implying that more efficient and exploratory methods are needed to cope with larger-scale problems.
TABLE VII: Average PDR on each benchmark of LAB. The highest PDR is highlighted through bolding and being underlined.
Benchmark r-CP no-LSS no-Degrade no-FSS
$ss1$ 0.00% 0.00% 0.00% 0.00%
$ss2$ 0.00% 0.00% 4.81% 0.00%
$ls1$ 0.16% 0.16% 0.16% 0.73%
$ls2$ 0.10% 1.07% 0.63% 0.74%
$ls3$0.21% 0.63% 0.70% 0.54%
$ls4$ 1.07% -0.69% 0.09% 0.22%
Avg.PDR 0.19% 0.19% 1.06% 0.69%
VI-E Acceleration Performance of Accelerated Monte Carlo
The acceleration effect of Accelerated Monte Carlo (AMC) evaluation is validated. We randomly selected 10,000 solutions for evaluation, using regular Monte Carlo (MC) and AMC respectively. After recording the running time required for each 1000 evaluations, the result is shown in Figure 6.
The results show that the time consumed increases essentially linearly as the number of evaluation times increases. For different $T_{\max}$ settings, MC takes essentially the same amount of time since the number of simulations needed does not vary with $T_{\max}$, while AMC can obtain better acceleration compared to MC when $T_{\max}$ is small. For example, compared with the running time $526.601s$ of MC, AMC takes $458.449s$ to finish the evaluation when $T_{\max}=11$, while the number of fast-filtered solution is 10000. When $T_{\max}=23$, it takes AMC about $471.395s$, although the number of pre-fast filtered is 9718. As $T_{\max}$ increases, the difficulty of the chance constraint decreases, resulting in the increase of the percentage of feasible solutions in the whole solution space. Thus, the number of solutions that can be fast pre-filtered by sequences decreases, which leads to a decrease in the acceleration effect. The accelerated mechanism may have a negative impact on the overall efficiency of the evaluation as $T_{\max}$ becomes large. When $T_{\max}=27$, AMC and MC take almost the same time, while when $T_{\max}=35$, AMC needs to take $807.435s$ to finish rather than about $500s$ of MC. In conclusion, when applying the accelerated mechanism, a pre-assessment of whether the constraints are tight enough is necessary to determine whether the constraints are tight enough.
Figure 6: Acceleration performance of accelerated Monte Carlo on LAB-ls1. The mechanism has a positive effect when the constraints are tight, but has a negative impact on the overall efficiency as $T_{\max}$ becomes large.
VI-F Performance of Other Works
We also attempted to adjust two similar algorithms, Ji’s [15] and Yang’s [34], to solve our problem (see Appendix 4 for modification details). Due to the incapability of dynamic programming and mixed integer programming for solving large-scale DDCCMCKP, we conducted experiments on small-scale benchmarks to evaluate their performance. Results are shown in Table VIII, and we observed that the majority of the results were either ”infeasible” or ”fail”. ”fail” means that the algorithms fail to output any solution, while ”infeasible” means that the optimal solution is infeasible. The value of ”Feasibility” column is the total cost of solution. For Yang’s method, the poor performance was mainly due to the significant difference between the actual distribution and the Gaussian distribution, and the insufficient data to characterize an unknown distribution using sample mean and variance, leading to incorrect iteration. In Ji’s work, a uniform distribution was used and tested. The performance of their method is poor since it could not handle the long-tailed distribution of the candidate solutions in our problem and led to suboptimal performance.
TABLE VIII: Result of Ji’s [15] and Yang’s [34]
Benchmark Ji Yang
Time(s) Feasiblility Time(s) Feasiblility
LAB-$ss1$ fail 1.0 14.5
LAB-$ss2$ 0.4 infeasible - fail
LAB-$ss3$ 0.8 infeasible 1.4 21.5
LAB-$ss4$ 16 infeasible 43.2 17.4
APP-$ss1$ fail 0.1 infeasible
APP-$ss2$ 0.6 infeasible - fail
APP-$ss3$ fail 0.2 infeasible
APP-$ss4$ fail 13.5 infeasible
VII Conclusion
In this work, derived from practical applications, we study a new variant of MCKP called DDCCMCKP. To address this specific problem, a novel DDALS algorithm has been introduced. Based on local search framework, DDALS enables data-driven solution evaluation for the problem, and incorporates a surrogate weight to ensure the optimality of solutions. Moreover, a solution filtering mechanism is introduced to improve the feasibility of the output solutions. Two sets of benchmarks for DDCCMCKP have been proposed, one of which simulates the real business requirement of 5G E2E network scenario for a communication company. The experimental results demonstrate that DDALS is able to solve DDCCMCKP effectively and outperforms other baseline methods. Although the APP benchmark imposes higher demands on problem-solving abilitiy, the performance of DDALS is not satisfactory in small-scale problems. However, DDALS can solve DDCCMCKP well in most of the scenarios and maintain better performance than other baselines.
Overall, our proposed algorithm holds great promise for addressing this new variant of MCKP in the data-driven scenario. DDALS can serve as a baseline for future research. The code and benchmark sets will be open-sourced for further promoting research on DDCCMCKP. The algorithm framework can be further improved in two aspects. A new surrogate weight design may improve the efficiency and solution’s quality instead of the current design. From the other perspective, more effective filtering strategies that can balance the feasibility and optimality of the solutions could be explored in the future.
Appendix A Details about Concentration Inequalities
A-A Bernstein’s Inequality
Assume that there are $m$ classes at the E2E, each denoted as variable $X_{j},j=1,\ldots,m$. $X_{1},X_{2},\ldots,X_{m}$ are independent and satisfy for given some $\{(a_{j},b_{j}),\forall{j}\in\{1,\ldots,m\}\}$ and constant $C$ there are $P(X_{j}\in[a_{j},b_{j}])=1,\forall{j}\in\{1,\ldots,m\}$ and $b_{j}-a_{j}=c_{j}\leq C$. Denote $S=X_{1}+X_{2}+\ldots+X_{m}$ as the sum of random variables $X_{j}$ and $V=\sum_{i=1}^{m}Var(X_{j})$ is the sum of variances. Then we have Bernstein’s inequality,
$P(S\leq E(S)+\alpha)\geq 1-exp(-\frac{\alpha^{2}/2}{V+C\alpha/3})$ (4)
A-B K-order Hoeffding’s Inequality
Similar to Bernstein’s inequality, we assume that $X_{1},X_{2},\ldots,X_{m}$ are independent and satisfy for given some $a_{j},b_{j}$, there is $P(X_{j}\in[a_{j},b_{j}])=1,\forall{j}\in\{1,\ldots,m\}$. Denote new variable $Y_{j}=X_{j}-a_{j}$, which satisfies $P(Y_{j}\in[0,c_{j}])=1,\forall{j}\in\{1,\ldots,m\}$. Denote $S=Y_{1}+Y_{2}+\ldots+Y_{m}$ and we have $\mathbb{E}(Y_{j}^{k})=\mu_{j}^{k}$. $K$ is the order of origin moment. Thus, we have following $K$-order Hoeffding’s inequality:
$P(S\leq E(S)+\alpha)\geq$ (5)
$1-exp(-\frac{2\alpha^{2}}{\sum_{j=1}^{m}b_{j}^{2}C_{K}(\frac{4\alpha b_{j}}{D},b_{j},\mu_{j}^{1},\ldots,\mu_{j}^{K})})$
where $D=\sum_{j=1}^{m}(\mu_{j}^{2}/\mu_{j}^{1})^{2}$ and $C_{K}(\frac{4\alpha b_{j}}{D},b_{j},\mu_{j}^{1},\ldots,\mu_{j}^{K})$ is a value calculated by all the orders of origin moments of $Y_{j}$. However, the concentration inequality does not make any assumptions about the form of the distribution of the variables; hence, the lower bound provided by the inequality may be imprecise. Additionally, as the inequalities are solely based on data-driven background, they may only serve as rough estimates for the upper and lower bounds of the variables and this may introduce extra errors into the confidence evaluation.
Since only data of items are available, if we manage to calculate the ratio of the combination of data meeting the upper bound $W$, to the total number of data combinations, then we can know whether the solution satisfies the chance constraint. For uncertain solution $S=[s_{1},s_{2},\ldots,s_{m}]$, we count the number of $\sum_{i=1}^{m}d_{i,j_{i},l}\leq W$ and check if it exceeds $P_{0}L^{m}$. Since $P_{0}$ is usually a number very close to 1 under high confidence level requirement, it is more practical to compare the number of $\sum_{i=1}^{m}d_{s_{i}l}\geq W$ with $(1-P_{0})L^{m}$. If the number of $\sum_{i=1}^{m}d_{s_{i}l}\geq W$ is greater than $(1-P_{0})L^{m}$, then we can determine quickly that it is an infeasible solution.
Appendix B Proof
Theorem B.1
The popped sum in the $k^{th}$ iteration is exactly the $k^{th}$ largest sum.
Proof B.1
Suppose the $i^{th}$ largest sum is $S_{r}=d_{s_{1}l_{1}}+d_{s_{2}l_{1}}+\ldots+d_{s_{m}l_{m}}$.
  • For any pair in the combination or queue $Q$ that has already been output, the combinations with indices less than or equal to it must have been popped or exist in $Q$.
  • Construct a chain $S$, then we have the following chain, denoted as $S_{0}\leq S_{1}\leq,\ldots,\leq S_{l}$.
    $d_{s_{1}0}+d_{s_{2}0}+\ldots+d_{s_{m}0},$
    $\geq$ $d_{s_{1}1}+d_{s_{2}0}+\ldots+d_{s_{m}0},$
    $\geq$ $\ldots$,… ,
    $\geq$ $d_{s_{1}l_{1}}+d_{s_{2}0}+\ldots+d_{s_{m}0}$
    $\geq$ $\ldots$,… ,
    $\geq$ $d_{s_{1}l_{1}}+d_{s_{2}l_{2}}+\ldots+d_{s_{m}0}$
    $\geq$ $\ldots$,… ,
    $\geq$ $d_{s_{1}l_{1}}+d_{s_{2}l_{2}}+\ldots+d_{s_{m}l_{m}}$
  • Let $S_{j}$ be the first sum in chain $S$ which was not popped from queue $Q$ before the $j^{th}$ iteration. During the $j^{th}$ iteration, when $S_{j-1}$ is popped, $S_{j}$ is added to $Q$.
  • Thus, during the $k^{th}$ iteration, the popped sum $S_{i}$ satisfies $S_{i}\leq S_{j}\leq S_{l}=d_{s_{1}l_{1}}+d_{s_{2}l_{1}}+\ldots+d_{s_{m}l_{m}}$.
  • By induction, since the $(i-1)^{th}$ largest sum has been popped before the $i^{th}$ iteration, so the sum popped in the $i^{th}$ iteration is exactly the $i^{th}$ largest sum0 $S_{l}=d_{s_{1}l_{1}}+d_{s_{2}l_{1}}+\ldots+d_{s_{m}l_{m}}$.
Then we finish the proof.
A simple example is shown when $m=2$ in Fig.7.
Figure 7: An example of exact evaluation when $m=2$
Appendix C A Theoretical Quantitative Analysis between Estimation Error and Sample Size
To theoretically quantify the relationship between estimation error and sample size of data, we construct a new random variable $S$ that is the sum of multiple random variables such that the variable itself is a random variable. Then we define a Bernoulli variable $Z\in\{0,1\}$, i.e.,
$Z=\left\{\begin{aligned} 1,&\quad\quad S\leq W\\ 0,&\quad\quad S>W\end{aligned}\right.$ (6)
for a combination of items with $L$ sample size, where $p$ is the true confidence level, i.e., the expectation $\mathbb{E}Z$ of the variable $Z$. Hoeffding’s inequality shows that for independent identically distributed random variables $Z_{1},Z_{2},\ldots,Z_{L}$ taking values $[a,b]$, for any $\epsilon>0$, the following inequality holds,
$Pr(|\mathbb{E}[Z]-\frac{1}{L}\sum_{l=1}^{L}Z_{l}|\geq\epsilon)\leq 2exp(-\frac{2L\epsilon^{2}}{(b-a)^{2}})$ (7)
For our problem, we are concerned with the true value is lower than the estimate and greater than the error, i.e., where $\epsilon$ is the estimation error corresponding to the confidence level. For the Bernoulli distribution, we have $\mathbb{E}Z=p$ and $a=0,b=1$. Thus, the inequality becomes
$Pr(p-f(Z_{1},Z_{2},\ldots,Z_{L})\leq-\epsilon)\leq exp(-2L\epsilon^{2})$ (8)
where $f(Z_{1},Z_{2},\ldots,Z_{L})$ denotes the confidence evaluation algorithm on some combination of items. Assuming that at least 5 of the 10 solutions in the output solution set are true feasible solutions, and $f(Z_{1},Z_{2},\ldots,Z_{L})$ is accurate, then we have $101-exp(-2L\epsilon^{2})\geq 5/$ 10. Finally, it leads to
$L\geq\frac{ln2}{2\epsilon^{2}}$ (9)
As this bound is derived from Hoeffding’s inequality, Chernoff bound is proposed in [42] and a more tight one is proposed in [43]. However, all of these inequalities indicate that a large amount of samples is necessary. The curve of the variation of $L$ with $\epsilon$ is shown in Figure 8.
Figure 8: Sample size requirement with different error in theory
The minimum sample sizes required for different estimation error are listed in Table IX.
Overall, Hoeffding provides the tightest bounds, especially suitable for larger errors. However, as error requirements become smaller, the bounds provided by Alamo are more advantageous. Practically, it is verified that the amount of data required to meet the corresponding error requirement for the true confidence will be much smaller than the theoretical value obtained.
TABLE IX: Relationship between EE and sample size
[…]
Hoeffding’s bound 139 13863 1386295 138629437
Chernoff’s bound [42] 278 27726 2772589 277258873
Alamo’s bound [43] 1573 15723 157227 1572269
Appendix D Deterministic method based on Gaussian Assumption and Mix Integer Linear Programming for DDCCMCKP
D-A Risk-averse knapsack problem method
Yang and Chakraborty [34] proposed an algorithm for the chance-constrained knapsack problem (CCKP) assuming known Gaussian distributions of the random variables (mean and variance are known). CCKP was then transformed into an iterative Risk-averse Knapsack Problem (RA-KP), which could be solved via dynamic programming. The implement on DDCCMCKP of this work is easy. We first calculated the sample mean and standard deviation of the item weights, $\mu_{ij}=\mathbb{E}(w_{ij}),\sigma^{2}_{ij}=Var(w_{ij}),i\in\{1,...,m\},j\in N_{i}$, and then plug them into the algorithm of Yang’s,
$\min\quad$ $\sum\limits_{i=1}^{m}\sum\limits_{j\in N_{i}}c_{ij}x_{ij}$ (10a)
s.t. $\sum\limits_{i=1}^{m}\sum\limits_{j\in N_{i}}\mu_{ij}x_{ij}+\Phi^{-1}(P_{0})\sqrt{\sum\limits_{i=1}^{m}\sum\limits_{j\in N_{i}}\sigma^{2}_{ij}x_{ij}}\leq W$ (10b)
$\sum\limits_{j\in N_{i}}x_{ij}=1,\forall i\in\mathcal{M}$ (10c)
$x_{ij}\in\{0,1\},\forall i\in\mathcal{M},j\in N_{i}$ (10d)
where $\Phi$ represents the CDF of a standard normal distribution with zero mean and unit variance.
D-B Mix Integer Linear Programming
Ji and Lejeune [15] studied the distribution-robust chance-constrained planning (DRCCP) optimization problem with data-driven Wasserstein fuzzy sets and proposed an algorithm and reconstruction framework suitable for all types of distribution-robust chance-constrained optimization problems, which was then applied to multidimensional knapsack problems. When the decision variables are binary, the proposed LP (who becomes MILP) formulations are equivalent reformulations of the CC. To fit the CC into [15]’s modelling and solution framework, we first formulate $\text{Prob}(\sum\limits_{i=1}^{m}\sum\limits_{j\in N_{i}}w_{ij}x_{ij}\leq W)\geq P_{0}$ as:
$\text{Prob}(\sum\limits_{i=1}^{m}\sum\limits_{j\in N_{i}}w_{ij}x_{ij}>W)\leq 1-P_{0}$ (11)
Define $b=-W,\epsilon=1-P_{0},\tilde{w}_{ij}=-w_{ij},\forall i\in\mathcal{M},j\in N_{i}$. Let $\bar{x}\in\{0,1\}^{NM}$ as the vectorization of $x\in\{0,1\}^{N\times M}$, $\xi\in\mathbb{R}^{NM}$ is the vectorization of $\hat{w}\in\mathbb{R}^{N\times M}$. Finally, we have the standard form of 11 as [15]:
$\text{Prob}(\xi^{T}\bar{x}<b)\leq\epsilon$ (12)
Then we can utilize Theorem 10 in [15] to reformulate the above inequation 11, and get the MILP formulation of DDCCMCKP:
$(\bar{x},\lambda,\eta,s,\phi)\in\mathcal{Z}_{LP}^{MCKP}$ (13)
$=$ $\left\{\begin{array}[]{lr}\lambda\theta+\frac{1}{L}\sum_{l=1}^{L}s_{l}\leq\epsilon,&\\ \lambda\geq 0,&\\ s_{l}\geq 0,\eta_{l}\geq 0,&\forall l\in\mathcal{L}\\ 1+\eta_{l}b^{{}^{\prime}}-\sum_{q\in\mathcal{Q}}\xi_{ql}^{0}\phi_{ql}\leq s_{l},&\forall l\in\mathcal{L}\\ \phi_{ql}\leq\lambda,&\forall q\in\mathcal{Q},l\in\mathcal{L}\\ \phi_{ql}\leq\eta_{l},&\forall q\in\mathcal{Q},l\in\mathcal{L}\\ \phi_{ql}\leq U_{\eta_{l}}\bar{x}_{q},&\forall q\in\mathcal{Q},l\in\mathcal{L}\\ \phi_{ql}\geq 0,&\forall q\in\mathcal{Q},l\in\mathcal{L}\\ \phi_{ql}\geq\eta_{l}+U_{\eta_{l}}\bar{x}_{q}-U_{\eta_{l}},&\forall q\in\mathcal{Q},l\in\mathcal{L}\\ \sum_{i=Nm-N+1}^{Nm}\bar{x}_{i}=1,&\forall m\in\mathcal{M}\\ \bar{x}_{q}\in\{0,1\},&\forall q\in\mathcal{Q}\\ \end{array}\right.$
where $b^{{}^{\prime}}=b-\delta$, $\delta$ is a user-defined infinite small positive number. $\mathcal{Q}$ is the set of all $N\times M$ items, $\mathcal{M}$ is the set of $M$ classes. Note that $\sum_{i=Nm-N+1}^{Nm}\bar{x}_{i}=1,\forall m\in\mathcal{M}$ is multiple-choice constraint in MCKP. $U_{\eta}$ is the upper bound of $\eta$ and can be determined by:
$U_{\eta_{l}}=\left\{\begin{array}[]{lr}\mathcal{K}^{{}^{\prime}*},&if\exists\bar{x}:\bar{x}^{T}\xi_{l}^{0}-b>0\\ 0,&otherwise\\ \end{array}\right.$ (14)
where $*\mathcal{K}^{{}^{\prime}*}$ is the optimal objective value of the linear programming problem:
$\mathcal{K}^{{}^{\prime}*}=$ $\min\bar{x}^{T}\xi_{l}^{0}-b$ (15)
s.t. $b-\bar{x}^{T}\xi_{l}^{0}\leq 0-\tau^{{}^{\prime}}$
$\bar{x}\in\mathcal{\bar{X}}$
where $\tau^{{}^{\prime}}$ a user-defined small non-negative number.
Overall, we first solve linear programming 15 to get $U_{\eta_{l}},\forall l\in\mathcal{L}$ then solve MILP $\mathcal{Z}_{LP}^{MCKP}$ 13 to find the optimal solution of DDCCMCKP.
Appendix E Complete experimental results
In this section, we show the complete experimental results of DDALS on LAB and APP.
TABLE X: Performance on benchmarks of LAB. The highest average real confidence level and the lowest cost are highlighted through bolding and being underlined.
Benchmark small $T_{\max}$ large $T_{\max}$
C ET ECL RCL FSR C ET ECL RCL FSR
ss1-Greedy 22.2$\pm$0.0 2.0$\pm$0.0 0$\pm$0.0 99.6$\pm$0.0 1 15.1$\pm$0.0 3.0$\pm$0.0 99.9$\pm$0.0 99.9$\pm$0.0 1
ss1-GA 18.7$\pm$0.1 354$\pm$4.2 99.8$\pm$0.3 99.6$\pm$0.8 0.8 10.8$\pm$0.0 253.3$\pm$4.1 99.9$\pm$0.1 99.0$\pm$1.2 0.6
ss1-EDA 21.5$\pm$1.4 354.5$\pm$2.9 - 99.5$\pm$0.5 0.1 13.0$\pm$3.4 234.7$\pm$2.8 99.8$\pm$0.3 99.9$\pm$0.1 1
ss1-DDMLS(O) 18.6$\pm$0.0 351.7$\pm$5.8 99.9$\pm$0.0 99.9$\pm$0.0 1 10.8$\pm$0.0 231$\pm$13.7 99.9$\pm$0.0 99.9$\pm$0.0 1
ss1-DDMLS(V1) 18.6$\pm$0.0 351.7$\pm$5.8 99.9$\pm$0.0 99.9$\pm$0.0 1 10.8$\pm$0.0 231$\pm$13.7 99.9$\pm$0.0 99.9$\pm$0.0 1
ss1-DDMLS(V2) 18.6$\pm$0.0 351.7$\pm$5.8 99.9$\pm$0.0 99.9$\pm$0.0 1 10.8$\pm$0.0 231$\pm$13.7 99.9$\pm$0.0 99.9$\pm$0.0 1
ss1-DDMLS(V3) 20.7$\pm$2.0 360 99.6$\pm$0.3 99.2$\pm$0.9 0.3 13.0$\pm$1.7 218 99.8$\pm$0.3 99.1$\pm$1.0 0.7
ss2-Greedy 19.4$\pm$0.0 2.0$\pm$0.0 99.9$\pm$0.0 99.9$\pm$0.0 1 16.5$\pm$0.0 4.0$\pm$0.0 99.9$\pm$0.0 99.9$\pm$0.0 1
ss2-GA 19.4$\pm$0.0 384.5$\pm$2.6 99.9$\pm$0.0 99.9$\pm$0.0 1 11.1$\pm$1.4 303.6$\pm$3.4 99.7$\pm$0.2 99.7$\pm$0.4 1
ss2-EDA 19.9$\pm$0.4 385.9$\pm$2.7 99.9$\pm$0.0 99.9$\pm$0.1 1 14.8$\pm$2.2 304.5$\pm$2.6 99.7$\pm$0.2 99.7$\pm$0.4 1
ss2-DDMLS(O) 19.4$\pm$0.0 378.8$\pm$8.2 99.9$\pm$0.0 99.9$\pm$0.0 1 10.4$\pm$0.0 298.9$\pm$15.5 99.8$\pm$0.0 99.9$\pm$0.0 1
ss2-DDMLS(V1) 19.4$\pm$0.0 378.8$\pm$8.2 99.9$\pm$0.0 99.9$\pm$0.0 1 10.4$\pm$0.0 298.9$\pm$15.5 99.8$\pm$0.0 99.9$\pm$0.0 1
ss2-DDMLS(V2) 19.4$\pm$0.0 378.8$\pm$8.2 99.9$\pm$0.0 99.9$\pm$0.0 1 12.3$\pm$2.0 298.9$\pm$15.5 99.9$\pm$0.1 99.9$\pm$0.0 1
ss2-DDMLS(V3) 21.8$\pm$2.4 377 99.9$\pm$0.0 99.9$\pm$0 1 13.8$\pm$1.6 319 99.8$\pm$0.3 99.9$\pm$0.1 1
ss3-Greedy 24.4$\pm$0.0 2.0$\pm$0.0 100.0$\pm$0.0 99.9$\pm$0.0 1 21.8$\pm$0.0 3.0$\pm$0.0 100$\pm$0 99.9$\pm$0.0 1
ss3-GA 22.6$\pm$1.5 303.4$\pm$2.8 99.9$\pm$0.0 99.7$\pm$0.2 1 12.9$\pm$0.7 262.4$\pm$1.6 99.9$\pm$0.0 99.9$\pm$0.1 1
ss3-EDA 24.1$\pm$2.3 304.1$\pm$2.3 99.9$\pm$0.1 99.6$\pm$0.6 0.9 14.0$\pm$2.0 264.5$\pm$2.9 99.9$\pm$0.0 99.9$\pm$0.1 1
ss3-DDALS(O) 21.4$\pm$0.0 291.9$\pm$6.2 99.9$\pm$0.0 99.5$\pm$0.0 1 12.6$\pm$0.0 254.4$\pm$4.7 100.0$\pm$0.0 99.9$\pm$0.0 1
ss3-DDALS(V1) 21.4$\pm$0.0 291.9$\pm$6.2 99.9$\pm$0.0 99.5$\pm$0.0 1 12.6$\pm$0.0 254.4$\pm$4.7 100.0$\pm$0.0 99.9$\pm$0.0 1
ss3-DDALS(V2) 21.4$\pm$0.0 291.9$\pm$6.2 99.9$\pm$0.0 99.5$\pm$0.0 1 12.6$\pm$0.0 254.4$\pm$4.7 100.0$\pm$0.0 99.9$\pm$0.0 1
ss3-DDALS(V3) 23.9$\pm$1.7 279 99.8$\pm$0.2 99.1$\pm$0.8 0.8 15.7$\pm$1.9 260 99.9$\pm$0.3 99.6$\pm$0.9 0.9
ss4-Greedy 30.1$\pm$0.0 2.0$\pm$0.0 99.9$\pm$0.0 99.9$\pm$0.0 1 27.5$\pm$0.0 3.0$\pm$0.0 99.7$\pm$0.0 99.4$\pm$0.0 1
ss4-GA 23.8$\pm$5.5 768.9$\pm$8.5 99.9$\pm$0.3 99.7$\pm$0.4 1 17.2$\pm$3.0 562.6$\pm$5.7 99.7$\pm$0.3 99.6$\pm$0.3 1
ss4-EDA 17.4$\pm$0.0 763.4$\pm$2.5 99.9$\pm$0.0 99.9$\pm$0.0 1 13.0$\pm$1.0 553.4$\pm$2.7 99.8$\pm$0.2 98.9$\pm$0.4 1
ss4-DDALS(O) 17.4$\pm$0.0 759.9$\pm$25.8 99.9$\pm$0.0 99.9$\pm$0.0 1 11.2$\pm$0.0 525.3$\pm$39.1 99.2$\pm$0.0 99.0$\pm$0.0 1
ss4-DDALS(V1) 17.4$\pm$0.0 759.9$\pm$25.8 99.9$\pm$0.0 99.9$\pm$0.0 1 12.1$\pm$0.0 525.3$\pm$39.1 99.5$\pm$0.0 98.9$\pm$0.0 1
ss4-DDALS(V2) 17.4$\pm$0.0 759.9$\pm$25.8 99.9$\pm$0.0 99.9$\pm$0.0 1 14.0$\pm$1.2 525.3$\pm$39.1 99.9$\pm$0.2 99.0$\pm$0.3 1
ss4-DDALS(V3) 24.6$\pm$4.0 735 99.8$\pm$0.2 99.5$\pm$0.5 1 13.8$\pm$1.3 561 99.8$\pm$0.3 99.5$\pm$0.5 1
ls1-Greedy 58.8$\pm$0.0 2.0$\pm$0.0 99.8$\pm$0.0 99.8$\pm$0.0 1 56.0$\pm$0.0 3.0$\pm$0.0 99.6$\pm$0.0 99.5$\pm$0.0 1
ls1-GA 50.6$\pm$0.9 3007.2$\pm$4.1 99.1$\pm$0.1 99.2$\pm$0.1 1 41.0$\pm$4.7 1412.6$\pm$8.3 99.3$\pm$0.3 99.2$\pm$0.4 1
ls1-EDA 54.7$\pm$2.1 3003.1$\pm$2.2 99.2$\pm$0.1 99.3$\pm$0.1 1 37.7$\pm$0 1404.6$\pm$2.3 99.3$\pm$0.0 99.3$\pm$0.0 1
ls1-DDMLS(O) 49.8$\pm$0.1 2928.8$\pm$71.5 99.3$\pm$0.0 99.2$\pm$0.0 1 32.0$\pm$0 1371.3$\pm$12.4 99.2$\pm$0.0 99.1$\pm$0.0 1
ls1-DDMLS(V1) 52.2$\pm$0.3 2928.8$\pm$71.5 99.6$\pm$0.1 99.6$\pm$0.1 1 34.6$\pm$0.9 1371.3$\pm$12.4 99.6$\pm$0.0 99.6$\pm$0.1 1
ls1-DDMLS(V2) 49.9$\pm$0.3 2928.8$\pm$71.5 99.3$\pm$0.0 99.3$\pm$0.1 1 33.1$\pm$1.1 1371.3$\pm$12.4 99.4$\pm$0.2 99.3$\pm$0.0 1
ls1-DDMLS(V3) 51.2$\pm$0.9 2919 99.3$\pm$0.2 99.3$\pm$0.3 1 35.1$\pm$1.5 1383 99.3$\pm$0.2 99.2$\pm$0.3 1
ls2-Greedy 66.6$\pm$0.0 2.0$\pm$0.0 99.7$\pm$0.0 99.7$\pm$0.0 1 66.6$\pm$0.0 2.0$\pm$0.0 99.9$\pm$0.0 100.0$\pm$0.0 1
ls2-GA 53.1$\pm$2.9 9623.1$\pm$21.8 99.2$\pm$0.1 99.2$\pm$0.2 1 43.3$\pm$3.1 3317.4$\pm$ 17.2 99.6$\pm$0.3 99.5$\pm$0.3 1
ls2-EDA 53.5$\pm$3.1 9604.4$\pm$2.4 99.1$\pm$0.0 99.2$\pm$0.0 1 38.7$\pm$2.7 3304.6$\pm$ 2.5 99.5$\pm$0.3 99.5$\pm$0.2 1
ls2-DDMLS(O) 47.9$\pm$0.0 9548.9$\pm$46.4 99.0$\pm$0.0 99.0$\pm$0.0 1 27.2$\pm$0.0 3305.6$\pm$342.6 99.3$\pm$0.0 99.3$\pm$0.0 1
ls2-DDMLS(V1) 51.3$\pm$0.0 9548.9$\pm$46.4 99.4$\pm$0.0 99.5$\pm$0.0 1 27.9$\pm$0.4 3305.6$\pm$342.6 99.6$\pm$0.0 99.6$\pm$0.0 1
ls2-DDMLS(V2) 49.1$\pm$0.0 9548.9$\pm$46.4 99.2$\pm$0.0 99.2$\pm$0.0 1 28.0$\pm$0.5 3305.6$\pm$342.6 99.6$\pm$0.1 99.6$\pm$0.1 1
ls2-DDMLS(V3) 51.2$\pm$2.4 9518 99.2$\pm$0.1 99.2$\pm$0.1 1 29.2$\pm$1 3197 99.4$\pm$0.2 99.4$\pm$0.3 1
ls3-Greedy 101.8$\pm$0.0 3.0$\pm$0.0 99.9$\pm$0.0 99.9$\pm$0.0 1 101.8$\pm$0.0 4.0$\pm$0.0 99.8$\pm$0.0 99.9$\pm$0.0 1
ls3-GA 85.7$\pm$5.6 7213.2$\pm$11.3 99.1$\pm$0.1 98.8$\pm$0.1 1 71.3$\pm$4.0 5209.5$\pm$7.2 99.4$\pm$0.3 99.3$\pm$0.3 1
ls3-EDA 82.5$\pm$0.4 7202.6$\pm$2.8 99.1$\pm$0.2 98.8$\pm$0.2 1 73.4$\pm$3.3 5203.7$\pm$2.8 99.7$\pm$0.3 99.6$\pm$0.4 1
ls3-DDMLS(O) 74.9$\pm$0.2 7233.5$\pm$58.3 99.1$\pm$0.0 98.7$\pm$0.0 1 63.4$\pm$0.9 5217.5$\pm$274.1 99.3$\pm$0.2 99.1$\pm$0.2 1
ls3-DDMLS(V1) 79.6$\pm$1.3 7233.5$\pm$58.3 99.6$\pm$0.1 99.4$\pm$0.1 1 65.2$\pm$2.1 5217.5$\pm$274.1 99.6$\pm$0.1 99.5$\pm$0.2 1
ls3-DDMLS(V2) 75.9$\pm$0.7 7233.5$\pm$58.3 99.4$\pm$0.1 99.1$\pm$0.1 1 65.0$\pm$1.9 5217.5$\pm$274.1 99.7$\pm$0.2 99.6$\pm$0.2 1
ls3-DDMLS(V3) 77.3$\pm$1.5 7175 99.3$\pm$0.1 99.0$\pm$0.2 1 65.2$\pm$0.8 5109 99.4$\pm$0.3 99.3$\pm$0.4 1
ls4-Greedy 169.7$\pm$0.0 5$\pm$0.0 99.8$\pm$0.0 99.8$\pm$0.0 1 162.5$\pm$0.0 6.0$\pm$0.0 99.9$\pm$0.0 99.9$\pm$0.0 1
ls4-GA 123.1$\pm$8.8 11307.2$\pm$8.0 99.1$\pm$0.1 99.1$\pm$0.2 1 104.1$\pm$5.1 8409.1$\pm$5.6 99.1$\pm$0.2 99.2$\pm$0.2 1
ls4-EDA 126.2$\pm$2.3 11304.8$\pm$2.1 99.2$\pm$0.2 99.4$\pm$0.2 1 114.8$\pm$4.5 8403.4$\pm$2.8 99.6$\pm$0.2 99.6$\pm$0.2 1
ls4-DDALS(O) 106.6$\pm$0.8 11289.9$\pm$92.0 99.1$\pm$0.1 99.2$\pm$0.1 1 92.8$\pm$0.7 8376.6$\pm$378.8 99.1$\pm$0.1 99.2$\pm$0.1 1
ls4-DDALS(V1) 109.5$\pm$0.4 11289.9$\pm$92.0 99.7$\pm$0.1 99.7$\pm$0.1 1 97.9$\pm$2.2 8376.6$\pm$378.8 99.6$\pm$0.1 99.7$\pm$0.1 1
ls4-DDALS(V2) 108.9$\pm$1.3 11289.9$\pm$92.0 99.6$\pm$0.2 99.7$\pm$0.2 1 93.8$\pm$1.4 8376.6$\pm$378.8 99.3$\pm$0.2 99.4$\pm$0.2 1
ls4-DDALS(V3) 108.7$\pm$1.1 11171 99.2$\pm$0.2 99.3$\pm$0.2 1 96.5$\pm$1.7 8177 99.1$\pm$0.1 99.2$\pm$0.1 1
ls5-Greedy 213.8$\pm$0.0 3.0$\pm$0.0 99.9$\pm$0.0 99.8$\pm$0.0 1 205.5$\pm$0.0 4.0$\pm$0.0 99.9$\pm$0.0 99.9$\pm$0.0 1
ls5-GA 192.6$\pm$6.7 29013.4$\pm$7.6 99.0$\pm$0.0 98.8$\pm$0.0 1 170.2$\pm$7.7 18313$\pm$10.3 99.1$\pm$0.2 99.0$\pm$0.3 0.9
ls5-EDA 195.2$\pm$2.2 29004.2$\pm$3.1 99.1$\pm$0.1 99.0$\pm$0.2 1 165.9$\pm$1.9 18305.5$\pm$2.8 99.3$\pm$0.2 99.3$\pm$0.2 1
ls5-DDMLS(O) 178.1$\pm$0.8 28840.1$\pm$289.5 99.0$\pm$0.0 98.9$\pm$0.0 1 144.5$\pm$1.0 18276.1$\pm$395.4 99.0$\pm$0.0 99.0$\pm$0.1 1
ls5-DDMLS(V1) 184.5$\pm$1.4 28840.1$\pm$289.5 99.3$\pm$0.0 99.2$\pm$0.0 1 155.8$\pm$3.3 18276.1$\pm$395.4 99.6$\pm$0.1 99.5$\pm$0.1 1
ls5-DDMLS(V2) 179.6$\pm$1.9 28840.1$\pm$289.5 99.2$\pm$0.0 99.0$\pm$0.0 1 147.3$\pm$3.4 18276.1$\pm$395.4 99.2$\pm$0.2 99.1$\pm$0.2 1
ls5-DDMLS(V3) 182.9$\pm$1.6 28624 99.1$\pm$0.0 99.0$\pm$0.0 1 147.4$\pm$4.0 17573 99.1$\pm$0.2 99.1$\pm$0.2 1
ls6-Greedy 256.3$\pm$0.0 4.0$\pm$0.0 99.4$\pm$0.0 99.4$\pm$0.0 1 250.0$\pm$0.0 7.0$\pm$0.0 99.8$\pm$0.0 99.7$\pm$0.0 1
ls6-GA 233.5$\pm$14.0 34609.3$\pm$6.0 99.0$\pm$0.0 98.9$\pm$0.0 1 187.2$\pm$9.3 20211.9$\pm$8.3 99.0$\pm$0.0 99.0$\pm$0.2 0.9
ls6-EDA 214.1$\pm$2.8 34604.9$\pm$2.8 99.3$\pm$0.2 99.2$\pm$0.2 1 181.7$\pm$1.8 20205.1$\pm$2.7 99.4$\pm$0.2 99.3$\pm$0.2 1
ls6-DDMLS(O) 198.5$\pm$ 0.5 34570.7$\pm$619.3 99.0$\pm$0.0 98.9$\pm$0.0 1 162.2$\pm$1.1 20178.8$\pm$1031.6 99.1$\pm$0.1 99.0$\pm$0.1 1
ls6-DDMLS(V1) 206.8$\pm$ 2.3 34570.7$\pm$619.3 99.5$\pm$0.0 99.4$\pm$0.1 1 169.6$\pm$1.4 20178.8$\pm$1031.6 99.5$\pm$0.0 99.5$\pm$0.1 1
ls6-DDMLS(V2) 201.5$\pm$ 2.4 34570.7$\pm$619.3 99.3$\pm$0.2 99.2$\pm$0.2 1 163.1$\pm$1.5 20178.8$\pm$1031.6 99.3$\pm$0.1 99.3$\pm$0.1 1
ls6-DDMLS(V3) 202.6$\pm$ 2.0 33681 99.1$\pm$0.1 99.1$\pm$0.1 1 165.0$\pm$2.6 19387 99.2$\pm$0.2 99.1$\pm$0.1 1
TABLE XI: Performance on benchmarks of APP. Due to the introduction of retransmission mechanisms, there is a significant difference in the results between LAB and APP.
Benchmark small $T_{\max}$ large $T_{\max}$
C ET ECL RCL FSR C ET ECL RCL FSR
ss1-Greedy 9.2$\pm$0.0 4.0$\pm$0.0 99.3$\pm$0.0 96.7$\pm$0 0 9.0$\pm$0.0 60.0$\pm$0.0 99.6$\pm$0.0 98.3$\pm$0.0 0
ss1-GA 10.5$\pm$2.3 337.4$\pm$5.2 99.4$\pm$0.3 96.9$\pm$0.7 0 8.6$\pm$2.1 303.7$\pm$2.7 99.5$\pm$0.1 97.2$\pm$0.0 0
ss1-EDA 9.0$\pm$0.0 332.5$\pm$2.3 99.3$\pm$0.0 96.5$\pm$0.0 0 8.0$\pm$0.9 304.2$\pm$2.4 99.6$\pm$0.0 97.7$\pm$0.6 0
ss1-DDMLS(O) 9.0$\pm$0.0 311.4$\pm$10.2 99.2$\pm$0.0 96.5$\pm$0.0 0 7.3$\pm$0.0 285.5$\pm$14.8 99.6$\pm$0.0 97.2$\pm$0.0 0
ss1-DDMLS(V1) 14.0$\pm$0.0 311.4$\pm$10.2 99.9$\pm$0.0 97.2$\pm$0.0 0 7.3$\pm$0.0 285.5$\pm$14.8 99.6$\pm$0.0 97.2$\pm$0.0 0
ss1-DDMLS(V2) 13.5$\pm$1.5 311.4$\pm$10.2 99.8$\pm$0.2 97.1$\pm$0.2 0 7.3$\pm$0.0 285.5$\pm$14.8 99.6$\pm$0.0 97.2$\pm$0.0 0
ss1-DDMLS(V3) 14.8$\pm$3.4 321 99.7$\pm$0.3 98.0$\pm$1.0 0.5 11.1$\pm$2.4 300 99.5$\pm$0.0 97.7$\pm$0.6 0.1
ss2-Greedy 6.4$\pm$0.0 3$\pm$0.0 99.7$\pm$0.0 98.6$\pm$0.0 1 6.4$\pm$0.0 3.0$\pm$0.0 99.8$\pm$0.0 99.0$\pm$0.0 1
ss2-GA 5.3$\pm$0.0 301.8$\pm$2.2 99.2$\pm$0.0 97.3$\pm$0.4 0 5.3$\pm$0.0 272.6$\pm$2.8 99.5$\pm$0.1 98.1$\pm$0.2 0
ss2-EDA 5.4$\pm$0.2 303.1$\pm$3.0 99.3$\pm$0.2 97.4$\pm$0.7 0.1 5.4$\pm$0.2 274.9$\pm$3.0 99.6$\pm$0.2 98.3$\pm$0.4 0.3
ss2-DDALS(O) 5.2$\pm$0.0 307$\pm$7.6 99.2$\pm$0.0 97.3$\pm$0.0 0 5.2$\pm$0.0 291$\pm$10.6 99.6$\pm$0.0 98.1$\pm$0.0 0
ss2-DDALS(V1) 5.5$\pm$0.1 307$\pm$7.6 99.8$\pm$0.0 97.5$\pm$0.3 0 5.2$\pm$0.0 291$\pm$10.1 99.6$\pm$0.0 98.1$\pm$0.0 0
ss2-DDALS(V2) 5.5$\pm$0.1 307$\pm$7.6 99.8$\pm$0.0 97.5$\pm$0.3 0 5.4$\pm$0.1 291$\pm$10.1 99.9$\pm$0.1 98.2$\pm$0.1 0
ss2-DDALS(V3) 5.4$\pm$0.1 295 99.4$\pm$0.3 97.7$\pm$0.4 0 5.3$\pm$0.1 272.6$\pm$2.8 99.5$\pm$0.1 98.1$\pm$0.2 0
ss3-Greedy 18.3$\pm$0.0 6.0$\pm$0.0 99.9$\pm$0.0 96.3$\pm$0 0 13.8$\pm$0.0 8.0$\pm$0.0 99.6$\pm$0.0 97.6$\pm$0.0 0
ss3-GA 17.5$\pm$0.3 634.6$\pm$4.0 99.3$\pm$0.3 95.3$\pm$0.5 0 12.1$\pm$1.0 503.6$\pm$2.5 99.2$\pm$0.2 94.0$\pm$1.6 0
ss3-EDA 17.1$\pm$0.3 634.5$\pm$3.5 99.1$\pm$0.2 93.9$\pm$1.1 0 11.2$\pm$0.2 503.7$\pm$3.1 99.5$\pm$0.2 94.7$\pm$0.8 0
ss3-DDMLS(O) 17.1$\pm$0.2 675.9$\pm$36.2 99.1$\pm$0.1 94.2$\pm$0.7 0 11.1$\pm$0.1 499.7$\pm$21.7 99.5$\pm$0.2 94.6$\pm$0.7 0
ss3-DDMLS(V1) 18.1$\pm$0.2 675.9$\pm$36.2 99.6$\pm$0.1 96.7$\pm$0.6 0 11.5$\pm$1.0 499.7$\pm$ 1.7 99.6$\pm$0.0 95.4$\pm$1.1 0
ss3-DDMLS(V2) 18.1$\pm$0.3 675.9$\pm$36.2 99.8$\pm$0.2 96.7$\pm$0.7 0 11.3$\pm$0.8 499.7$\pm$21.7 99.6$\pm$0.2 94.9$\pm$1.0 0
ss3-DDMLS(V3) 19.0$\pm$0.8 632 99.5$\pm$0.3 96.3$\pm$0.5 0 12.4$\pm$0.8 499 99.5$\pm$0.2 95.5$\pm$0.1 0
ss4-Greedy 34.9$\pm$0.0 3.0$\pm$0.0 100.0$\pm$0.0 100.0$\pm$0.0 1 27.2$\pm$0.0 4.0$\pm$0.0 100.0$\pm$0.0 99.0$\pm$0.0 1
ss4-GA 27.3$\pm$1.1 2010$\pm$8.4 99.8$\pm$0.3 90.4$\pm$0.5 0 21.5$\pm$2.9 1614.8$\pm$9.8 99.6$\pm$0.3 93.2$\pm$4.8 0
ss4-EDA 26.1$\pm$0.0 2005.7$\pm$2.3 100.0$\pm$0.0 90.1$\pm$0.0 0 16.4$\pm$0.0 1602.7$\pm$2.1 99.9$\pm$0.0 80.3$\pm$0.0 0
ss4-DDALS(O) 26.1$\pm$0.0 1988.2$\pm$30.4 100.0$\pm$0.0 90.1$\pm$0.0 0 16.7$\pm$0.9 1579.4$\pm$76.8 99.8$\pm$0.2 81.8$\pm$4.4 0
ss4-DDALS(V1) 26.1$\pm$0.0 1988.2$\pm$30.4 100.0$\pm$0.0 90.1$\pm$0.0 0 16.7$\pm$1.0 1579.4$\pm$76.8 99.9$\pm$0.0 81.5$\pm$3.6 0
ss4-DDALS(V2) 26.1$\pm$0.0 1988.2$\pm$30.4 100.0$\pm$0.0 90.1$\pm$0.0 0 18.0$\pm$1.9 1579.4$\pm$76.8 99.9$\pm$0.1 86.1$\pm$7.0 0
ss4-DDALS(V3) 29.1$\pm$2.4 2022 99.9$\pm$0.2 91.3$\pm$2.9 0 19.9$\pm$1.7 1589 99.7$\pm$0.3 92.0$\pm$5.3 0
ls1-Greedy 84.1$\pm$0.0 2.0$\pm$0.0 99.3$\pm$0.0 99.5$\pm$0.0 1 65.3$\pm$0.0 5.0$\pm$0.0 99.1$\pm$0.0 99.0$\pm$0.0 1
ls1-GA 65.0$\pm$1.9 6432.1$\pm$6.4 99.1$\pm$0.2 98.8$\pm$0.1 1 59.3$\pm$3.9 6332.7$\pm$6.7 99.0$\pm$0.0 98.8$\pm$0.2 1
ls1-EDA 68.9$\pm$0.0 6423.1$\pm$2.9 99.1$\pm$0.0 99.4$\pm$0 1 61.35$\pm$0 6323.2$\pm$3.2 99.2$\pm$0.0 99.1$\pm$0.0 1
ls1-DDMLS(O) 61.5$\pm$0.0 6417.8$\pm$16.3 99.2$\pm$0.0 98.5$\pm$0 1 53.9$\pm$2.3 6318$\pm$65.9 99.0$\pm$0.0 98.6$\pm$0.1 0.9
ls1-DDMLS(V1) 63.4$\pm$1.4 6417.8$\pm$16.3 99.6$\pm$0.0 99.1$\pm$0.1 1 61.8$\pm$0.5 6318$\pm$65.9 99.7$\pm$0.3 99.4$\pm$0.2 1
ls1-DDMLS(V2) 62.3$\pm$ 0.3 6417.8$\pm$16.3 99.4$\pm$0.2 98.9$\pm$0.3 1 54.8$\pm$2.8 6318$\pm$65.9 99.1$\pm$0.1 98.8$\pm$0.1 1
ls1-DDMLS(V3) 62.9$\pm$0.8 6403 99.2$\pm$0.1 98.7$\pm$0.2 1 57.4$\pm$2.6 6340$\pm$65.9 99.1$\pm$0.0 98.7$\pm$0.1 1
ls2-Greedy 85.7$\pm$0.0 3.0$\pm$0.0 99.9$\pm$0.0 99.9$\pm$0.0 1 85.7$\pm$0.0 3.0$\pm$0.0 100.0$\pm$0.0 100.0$\pm$0.0 1
ls2-GA 76.9$\pm$2.7 22617.4$\pm$10.9 99.0$\pm$0.1 98.9$\pm$0.2 1 61.2$\pm$2.1 21327.1$\pm$20.3 99.1$\pm$0.0 99.1$\pm$0.1 1
ls2-EDA 75.9$\pm$0.0 22604.3$\pm$3.3 99.0$\pm$0.0 98.9$\pm$0.0 1 59.2$\pm$0.3 21306.5$\pm$2.5 99.2$\pm$0.0 99.2$\pm$0.1 1
ls2-DDALS(O) 71.9$\pm$0.3 22591.3$\pm$63.6 99.0$\pm$0.0 99.0$\pm$0.0 1 57.2$\pm$0.0 21233.5$\pm$51.9 99.3$\pm$0.1 99.3$\pm$0.1 1
ls2-DDALS(V1) 76.6$\pm$1.1 22591.3$\pm$63.6 99.6$\pm$0.1 99.5$\pm$0.1 1 58.5$\pm$0.8 21233.5$\pm$51.9 99.6$\pm$0.1 99.6$\pm$0.1 1
ls2-DDALS(V2) 72.9$\pm$0.6 22591.3$\pm$63.6 99.3$\pm$0.2 99.2$\pm$0.2 1 57.5$\pm$0.2 21233.5$\pm$51.9 99.4$\pm$0.1 99.5$\pm$0.1 1
ls2-DDALS(V3) 73.1$\pm$0.7 22564 99.2$\pm$0.1 99.1$\pm$0.1 1 57.9$\pm$0.4 21245 99.2$\pm$0.1 99.2$\pm$0.2 1
ls3-Greedy 158.3$\pm$0.0 3.0$\pm$0.0 99.4$\pm$0.0 99.3$\pm$0.0 1 158.3$\pm$0.0 3.0$\pm$0.0 99.6$\pm$0.0 99.9$\pm$0.0 1
ls3-GA 119.4$\pm$3.9 21008.5$\pm$7.5 99.1$\pm$0.1 98.4$\pm$0.4 0.7 122.6$\pm$5.3 21007.5$\pm$4.2 99.0$\pm$0.1 98.4$\pm$0.4 0.5
ls3-EDA 117.0$\pm$0.7 21005.3$\pm$3.0 99.4$\pm$0.0 98.9$\pm$0.0 1 115.0$\pm$2.0 21004.8$\pm$2.6 99.3$\pm$0.2 99.0$\pm$0.1 0.6
ls3-DDMLS(O) 112.0$\pm$1.2 20942.2$\pm$211.2 99.1$\pm$0.0 98.9$\pm$0.1 0.9 111.4$\pm$ 1.5 20956.2$\pm$193.7 99.3$\pm$0.3 98.8$\pm$0.4 0.5
ls3-DDMLS(V1) 119.0$\pm$2.7 20942.2$\pm$211.2 99.5$\pm$0.1 98.9$\pm$0.3 0.9 114.4$\pm$4.3 20956.2$\pm$193.7 99.5$\pm$0.2 99.0$\pm$0.4 0.7
ls3-DDMLS(V2) 115.7$\pm$1.9 20942.2$\pm$211.2 99.4$\pm$0.1 98.8$\pm$0.1 1 111.8$\pm$ 1.7 20956.2$\pm$193.7 99.4$\pm$0.3 98.9$\pm$0.4 0.6
ls3-DDMLS(V3) 115.6$\pm$1.5 20810 99.2$\pm$0.1 97.9$\pm$0.6 0.3 113.7$\pm$0.7 20819 99.0$\pm$0.0 98.3$\pm$0.2 0.1
ls4-Greedy 207.3$\pm$0.0 13.0$\pm$0.0 99.0$\pm$0.0 98.9$\pm$0.0 1 195.3$\pm$0.0 17.0$\pm$0.0 99.4$\pm$0.0 99.4$\pm$0.0 1
ls4-GA 208.5$\pm$8.3 48014.1$\pm$12.1 99.0$\pm$0.0 98.6$\pm$0.3 0.7 175.2$\pm$10.3 45014.6$\pm$10.2 99.0$\pm$0.0 98.5$\pm$0.2 0.6
ls4-EDA 205.3$\pm$2.4 48004.5$\pm$3.3 99.0$\pm$0.0 99.0$\pm$0.0 1 163.8$\pm$1.3 45005.5$\pm$2.6 99.1$\pm$0.0 98.8$\pm$0.0 1
ls4-DDALS(O) 194.9$\pm$0.8 47962.8$\pm$67.7 99.0$\pm$0.0 98.9$\pm$0.0 1 163.3$\pm$3.6 44822.1$\pm$583.5 99.0$\pm$0.0 98.4$\pm$0.1 0.3
ls4-DDALS(V1) 203.2$\pm$2.0 47962.8$\pm$67.7 99.1$\pm$0.0 99.0$\pm$0.1 1 172.8$\pm$4.5 44822.1$\pm$583.5 99.3$\pm$0.1 98.9$\pm$0.3 0.8
ls4-DDALS(V2) 196.5$\pm$1.3 47962.8$\pm$67.7 99.0$\pm$0.0 99.0$\pm$0.0 1 165.3$\pm$4.3 44822.1$\pm$583.5 99.1$\pm$0.1 98.5$\pm$0.2 0.3
ls4-DDALS(V3) 196.9$\pm$2.1 48010 99.0$\pm$0.0 98.9$\pm$0.1 1 168.2$\pm$ 2.3 44174 99.0$\pm$0.0 98.9$\pm$ 0.1 1
ls5-Greedy 267.2$\pm$0.0 11.0$\pm$0.0 99.1$\pm$0.0 98.6$\pm$0.0 1 264.0$\pm$0.0 12.0$\pm$0.0 99.3$\pm$0.0 98.9$\pm$0.0 1
ls5-GA 241.2$\pm$9.4 81015.9$\pm$10.3 99.0$\pm$0.0 98.6$\pm$0.1 0.7 219.9$\pm$10.5 79012.6$\pm$7.5 99.0$\pm$0.0 98.5$\pm$0.2 0.6
ls5-EDA 220.3$\pm$0.3 81002.4$\pm$2.4 99.1$\pm$0.0 98.7$\pm$0.0 1 198.7$\pm$1.8 79005.3$\pm$3.1 99.1$\pm$0.2 98.6$\pm$0.2 1
ls5-DDMLS(O) 218.2$\pm$1.0 81144.5$\pm$283.4 99.0$\pm$0.0 98.6$\pm$0.0 1 199.4$\pm$2.2 78638.8$\pm$555.0 99.0$\pm$0.1 98.6$\pm$0.1 0.9
ls5-DDMLS(V1) 228.3$\pm$2.9 81144.5$\pm$283.4 99.2$\pm$0.1 98.9$\pm$0.2 1 209.3$\pm$5.8 78638.8$\pm$555.0 99.2$\pm$0.1 98.9$\pm$0.2 0.9
ls5-DDMLS(V2) 220.3$\pm$3.0 81144.5$\pm$283.4 99.1$\pm$0.1 98.7$\pm$0.1 1 200.1$\pm$2.7 78638.8$\pm$555.0 99.1$\pm$0.1 98.7$\pm$0.2 1
ls5-DDMLS(V3) 221.5$\pm$1.5 81213 99.0$\pm$0.0 98.7$\pm$0.0 1 202.2$\pm$3.1 78784 99.1$\pm$0.1 98.9$\pm$0.2 0.9
ls6-Greedy 388.8$\pm$0.0 4.0$\pm$0.0 99.1$\pm$0.0 98.6$\pm$0.0 1 374.4$\pm$0.0 12.0$\pm$0.0 99.7$\pm$0.0 99.6$\pm$0.0 1
ls6-GA 354.5$\pm$14.3 125315.6$\pm$10.0 99.0$\pm$0.0 98.2$\pm$0.1 0 325.3$\pm$10.5 121508.9$\pm$4.8 99.0$\pm$0.0 98.2$\pm$0.2 0.1
ls6-EDA 337.8$\pm$0.9 125305.4$\pm$2.4 99.0$\pm$0.0 98.6$\pm$0.0 1 301.6$\pm$1.7 121503.7$\pm$2.6 99.1$\pm$0.1 98.7$\pm$0.1 1
ls6-DDMLS(O) 324.2$\pm$0.9 125232.9$\pm$226.2 99.0$\pm$0.0 98.1$\pm$0.0 0 298.7$\pm$5.6 121483.4$\pm$943.5 99.0$\pm$0.0 98.4$\pm$0.1 0.3
ls6-DDMLS(V1) 333.1$\pm$1.9 125232.9$\pm$226.2 99.2$\pm$0.1 98.6$\pm$0.2 0.6 311.5$\pm$7.0 121483.4$\pm$943.5 99.3$\pm$0.1 98.6$\pm$0.3 0.6
ls6-DDMLS(V2) 325.5$\pm$2.7 125232.9$\pm$226.2 99.0$\pm$0.1 98.2$\pm$0.1 0.1 301.4$\pm$5.2 121483.4$\pm$943.5 99.1$\pm$0.2 98.5$\pm$0.2 0.5
ls6-DDMLS(V3) 328.3$\pm$2.4 124923 99.0$\pm$0.0 98.3$\pm$0.1 0 304.7$\pm$3.2 121648 99.1$\pm$0.0 98.4$\pm$0.1 0.1

### v2

@@
-Data-Driven Chance-Constrained Multiple-Choice Knapsack Problem: Model, Algorithms, and Applications
+Chance-Constrained Multiple-Choice Knapsack Problem: Model, Algorithms, and Applications
 Abstract
 The multiple-choice knapsack problem (MCKP) is a classic NP-hard combinatorial optimization problem.
-Motivated by several significant practical applications, this work investigates a novel variant of MCKP called data-driven chance-constrained multiple-choice knapsack problem (DDCCMCKP), where the item weight is a random variable with unknown probability distribution.
-We first present the problem formulation of DDCCMCKP, and then establish two benchmark sets.
-The first set contains synthetic instances, and the second set is devised to simulate a real-world application scenario of a certain telecommunication company.
-To solve DDCCMCKP, we propose a data-driven adaptive local search (DDALS) algorithm.
-The main merit of DDALS lies in evaluating solutions with chance constraints by data-driven methods, under the condition of unknown distributions and only historical sample data being available.
-The experimental results demonstrate the effectiveness of the proposed algorithm and show that it is superior to other baselines.
-Additionally, ablation experiments confirm the necessity of each component in the algorithm.
-Our proposed algorithm can serve as the baseline for future research, and the code and benchmark sets will be open-sourced to further promote research on this challenging problem.
+Motivated by several significant real-world applications, this work investigates a novel variant of MCKP called chance-constrained multiple-choice knapsack problem (CCMCKP), where the item weights are random variables.
+In particular, we focus on the practical scenario of CCMCKP, where the probability distributions of random weights are unknown but only sample data is available.
+We first present the problem formulation of CCMCKP and then establish two benchmark sets.
+The first set contains synthetic instances and the second set is devised to simulate a real-world application scenario of a certain telecommunication company.
+To solve CCMCKP, we propose a data-driven adaptive local search (DDALS) algorithm.
+The main novelty of DDALS lies in its data-driven solution evaluation approach that can effectively handle unknown probability distributions of item weights.
+Moreover, in cases with unknown distributions, high intensity of chance constraints, limited amount of sample data and large-scale problems, it still exhibits good performance.
+Experimental results demonstrate the superiority of DDALS over other baselines on the two benchmarks.
+Additionally, ablation studies confirm the effectiveness of each component of the algorithm.
+Finally, DDALS can serve as the baseline for future research, and the benchmark sets are open-sourced to further promote research on this challenging problem.
 Index Terms:
-Combinatorial optimization, data-driven approach, multiple-choice knapsack problem, evolutionary algorithm.
+Combinatorial optimization, data-driven optimization, chance-constrained multiple-choice knapsack problem, evolutionary algorithm.
 I Introduction
 The multiple-choice knapsack problem (MCKP) is a variant of the classic knapsack problem [1].
-MCKP finds practical applications in various areas, such as supply chain optimization [2], target market selection [3], investments portfolio selection [4], transportation programming [5], and component selection in the information technology systems [6].
+MCKP finds applications in various areas, such as supply chain optimization [2], target market selection [3], investments portfolio selection [4], transportation programming [5], and component selection in the information technology systems [6].
 Specifically, MCKP involves a set of items, each possessing a weight and a value.
-The items are divided into several disjoint classes and only one item can be selected in each class [1, 7].
-The objective of MCKP is to select exactly one item from each class such as to maximize the total value while adhering to a predefined weight limit (i.e., knapsack constraint).
+The items are divided into several disjoint classes.
+The objective of MCKP is to select exactly one item from each class [1, 7] such as to maximize the total value while adhering to a predefined weight limit (i.e., the knapsack constraint).
 Over the past several decades, plenty of methods for solving MCKP have been proposed, including greedy methods [8], branch-and-bound [9], dynamic programming [10], approximation algorithms [11], and heuristic search [12].
-Despite of significant research success in addressing the MCKP, those advanced methods are incapable of handling problems in real-world scenarios where item weights are randomly distributed instead of fixed.
-To handle random variables, a common approach is to model deterministic constraints as chance constraints, i.e., the probability (confidence level) of total weight (random variable) of the selected items not exceeding a given deterministic value is no less than a certain threshold [13, 14].
-Furthermore, the specific probability distributions of the random variables could be unknown [15, 16].
-Instead, one only has historical sample data to characterize the random variables.
-The collection of sample data for each item is assumed to be independent and uncorrelated.
-This variant of MCKP is referred to as data-driven chance-constrained multiple-choice knapsack problem (DDCCMCKP).
-A typical application of DDCCMCKP is the configuration of fifth-generation (5G) end-to-end (E2E) network.
+Despite the research progress made thus far, advanced methods for solving MCKP still fall short when confronted with scenarios where item weights are not fixed but are random variables.
+In such cases, the constraint is no longer the total weight not exceeding a fixed limit, but rather the probability (confidence level) of the total weight not exceeding the limit is no less than a certain threshold.
+In this work, we refer to this particular variant of MCKP as the chance-constrained multiple-choice knapsack problem (CCMCKP).
+Furthermore, in real-world applications of CCMCKP, the probability distributions of item weights are often unknown [13, 14], and sample data is the only available information for characterizing these random variables.
+The above-described CCMCKP finds applications in various domains.
+A typical example is the configuration of fifth-generation (5G) end-to-end (E2E) network.
 In the area of 5G for business [15], user plane time delay is one of the key performance indicators for low latency communication [16, 17].
-As illustrated in Figure 1, the complete packet transmission path in a 5G cellular network consists of several segments, i.e., eNB, backhaul, core network, Internet/cloud, and the transport between core network and Internet/cloud.
-More specifically, these segments correspond to classes, where for each class there exist different options (items) and only one option can be selected to implement the class.
-The total time delay of a packet transmission can be calculated as the sum of the delays induced by all the selected options, i.e., $(T_{eNB}+T_{Backhaul}+T_{Core}+T_{Transport}+T_{Internet})*2$.
-Moreover, the packet transmission delays of the implementation options are random variables with unknown distributions as they are affected by various factors, including workloads, network configurations, broadcast environment, and random failures of software and hardware.
-The configuration of 5G E2E network is to select an implementation option for each segment.
-Different configurations have different total costs and time delays, where the costs are fixed values, but time delays are unknown random variables.
-The objective is to minimize the total cost of network configuration while ensuring that the probability of meeting the delay is less than a given threshold under a certain value.
-Another application of DDCCMCKP is supply chain planning and optimization in modern industries [2, 3].
-The supply-chain operations reference (SCOR) model, published by the Supply-Chain Council [20], is a useful tool for the upper management of a firm to design and reconfigure the supply chain to achieve better performance.
-In SCOR, the E2E supply chain consists of four distinct processes: source, make, deliver, and plan (see Figure 2).
-Specifically, these processes can be further subdivided into supplier selection, R&D testing, production, transportation, and recycling handling, which correspond to different classes along the supply chain.
-Therefore, each class has various configuration options that can be influenced by random factors, such as market demand, cost fluctuations, and technology configurations.
+As illustrated in Figure 1, the complete packet transmission path in a 5G cellular network consists of several segments, such as eNB, backhaul, and core network.
+Specifically, these segments correspond to the classes in CCMCKP, where for each class there exist different implementation options (items) and only one option can be selected.
+The total time delay of a packet transmission is calculated as the sum of the delays induced by all the selected options.
+Here, packet transmission delays (item weights) are random variables with unknown probability distributions because they are generally affected by various unpredictable factors such as workloads, broadcast environment, as well as random failures of software and hardware.
+The configuration of 5G E2E network is to select an implementation option for each segment, such that the total cost of the network configuration is minimized while ensuring the probability of packet transmission delay not exceeding the limit is no less than a certain threshold.
+Another application of CCMCKP is supply chain planning and optimization in modern industries [2, 3].
+The supply-chain operations reference (SCOR) model, published by the Supply-Chain Council [18], is a useful tool to design and reconfigure the supply chain.
+As illustrated in Figure 2, in SCOR, the E2E supply chain consists of four distinct processes: source, make, deliver, and plan.
+Specifically, these processes can be further subdivided into supplier selection, R&D testing, production, transportation, and recycling handling, corresponding to the classes in CCMCKP.
+Each class has different configuration options and only one option can be selected.
+The processing delays of these options are random variables with unknown probability distributions because they are influenced by various random factors, such as market demand, cost fluctuations, and technology configurations.
+Finally, the goal is to select an option for each class to minimize the total cost, while ensuring the probability of total processing delay not exceeding the limit is no less than a certain threshold.
+Despite the practical significance of CCMCKP, currently there is no direct research on solving this challenging problem.
+On the other hand, there are alternative methods that can be adapted to solve CCMCKP, such as stochastic optimization (SO) [19] that assumes known probability distributions of item weights, and distributionally robust optimization (DRO) [13] that employs the uncertainty set technique.
+Nonetheless, these methods exhibit unsatisfactory performance when the assumed probability distribution does not hold, intensity of the chance constraint is high, or the amount of sample data is not sufficiently large (see Section VI-B).
+Motivated by these observations, we propose a novel data-driven adaptive local search (DDALS) algorithm to effectively address CCMCKP.
+Through rigorous experiments, we demonstrate the superior performance of DDALS in comparison to existing methods.
 Figure 1: 5G E2E network service model [17]
 Figure 2: SCOR model-based supply chain infrastructure [18]
-Since DDCCMCKP has not been paid research attention, it merits thorough investigation due to its extensive and noteworthy applications.
-Compared to MCKP, DDCCMCKP faces greater challenges due to the complexity of its constraints.
-There are two main difficulties: it is a chance-constrained optimization problem, and the distributions of the random variables are unknown.
-Therefore, a data-driven approach that does not rely on specific distribution information but rather available sample data to evaluate solutions must be proposed.
-The main contributions are summarized as follows:
+The main contributions of this work are summarized below.
   1. 1.
-    From the problem perspective, we define DDCCMCKP from several real-world scenarios and formulate the model.
+    From the optimization problem perspective, we formulate the model of CCMCKP.
 To the best of our knowledge, this is the first endeavor to address this specific problem.
-To further promote research on this problem, we have designed two sets of benchmarks.
-The first set is a simulated benchmark with a simple mechanism, while the other set is to simulate a real-world business requirement of a telecommunication company.
-The generated benchmarks will be made publicly available to promote further research.
+Then, we establish two benchmark sets of CCMCKP.
+The first set contains synthetic instances while the second set is devised to simulate a real-world business requirement of a certain telecommunication company.
+The generated benchmarks are available online to promote further research.
   2. 2.
-    From the algorithmic perspective, we propose a novel algorithm named data-driven adaptive local Search (DDALS).
-The first novelty of DDALS is that we evaluate solutions with chance constraints by a data-driven method.
-The second is that we design surrogate weights using statistics of sample data to ensure the performance of the search algorithm.
-Furthermore, our algorithm can adaptively jump between small and larger search regions for efficient searching of high-quality solutions.
-Lastly, we improve the proportion of feasible solutions through carefully designed filtering mechanisms.
+    From the algorithmic perspective, we propose DDALS to solve CCMCKP.
+The main novelty of DDALS lies in its data-driven solution evaluation approach that can effectively handle situations where the distributions of item weights are unknown.
+Besides, DDALS incorporate surrogate weights using statistics of sample data and can adaptively jump between small and large search regions to find high-quality solutions.
+Lastly, DDALS employs a carefully designed mechanism to filter out infeasible solutions, further enhancing its efficiency.
   3. 3.
     From the computational perspective, we experimentally confirm the effectiveness of DDALS.
-We adapt several commonly-used approaches as the baselines, including greedy search (Greedy), genetic algorithm (GA), and estimation of distribution algorithm (EDA).
-Experimental results show that DDALS outperforms other baseline methods.
-The proposed algorithm can be served as a baseline for comparison experiments on this challenging problem.
+First, experimental results show the superiority of DDALS over SO and DRO methods in cases with unknown distributions, high intensity of chance constraints, limited amount of sample data and large-scale problems.
+Then, further experiments show that DDALS outperforms alternative commonly used optimization methods including greedy search (Greedy), genetic algorithm (GA), and estimation of distribution algorithm (EDA).
+Lastly, ablation studies confirm the effectiveness of each component of DDALS.
+For future research, DDALS can serve as a baseline method for comparison experiments on solving CCMCKP.
 The rest of the paper is organized as follows.
-In Section II, we present a comprehensive review of related works.
-Section III provides a formal definition of DDCCMCKP.
-In Section IV, DDALS is introduced and a detailed explanation of its implementation is provided.
-In Section V, the benchmarks we designed are described in details.
-The experimental results and analyses are given in Section VI.
+Section II presents the review of related works.
+Section III provides a formal definition of CCMCKP.
+In Section IV, DDALS is introduced and a detailed explanation is provided.
+In Section V, the benchmarks are described in details.
+Experimental results and analyses are given in Section VI.
 Finally, we summarize our findings and provide suggestions for future research in Section VII.
 II Related works
-In this section, we will review the relevant literature.
-We first review existing research on the traditional MCKP.
-The second category of research focuses on chance-constrained optimization.
-Since the methods developed in robust optimization can solve chance-constrained optimization by constructing proper uncertainty sets [21, 22], we further review robust optimization and distributionally robust optimization.
-While robust optimization approach minimizes the worst-case scenario under the uncertainty of random variables, the goal of distributionally robust optimization is to ensure the robustness of the optimal solution in the distribution-unknown situations.
+In this section, we first review existing research on the classic MCKP and then introduce the research on chance-constrained optimization.
+Finally, we review DRO that solves chance-constrained optimization problems by constructing uncertainty sets.
 II-A Algorithms for MCKP
-There have been various stochastic variations of the classical 0-1 MCKP that have been extensively studied in the literature.
-For instance, Dyer et al.[9] proposed a branch and bound method to solve the MCKP.
+Classic MCKP has been extensively studied in the literature.
+For instance, Dyer et al.
+[9] proposed a branch and bound method to solve MCKP.
 Pisinger [20] presented a minimal algorithm including linear relaxation, greedy algorithm and dynamic programming.
-Lamanna et al.
-[24] introduced a new algorithm for solving the multidimensional MCKP based on Kernel search, which divided the solution process of the algorithm into two phases with complementary objectives.
-Gens and Levner [11] proposed a fast approximation algorithm for the MCKP with a 4/5 approximation ratio of the optimal solution using binary search.
-He et al.
-[25] developed an improved approximate binary search algorithm with an approximation ratio of $3+(1/2)^{t}$.
+Gens and Levner [11] and He et al.
+[21] both proposed approximation algorithms for MCKP.
 Hifi et al.
-[12] proposed a reactive local search algorithm for the multidimensional MCKP.
-Mkaouar et al.
-[26] introduced a method based on the artificial bee colony algorithm to solve the multidimensional MCKP.
-In general, existing researches on MCKP cannot be applied, since the solutions in DDCCMCKP can not be evaluated by deterministic approaches.
+[12] presented a reactive local search algorithm, Mkaouar et al.
+[22] employed the artificial bee colony algorithm, and Lamanna et al.
+[23] introduced a Kernel Search-based approach for solving multidimensional MCKP.
+However, these deterministic methods cannot be applied to CCMCKP as they cannot handle stochastic item weights.
 II-B Chance-Constrained Optimization
-Chance-constrained optimization plays a critical role in stochastic optimization [27, 13, 14].
-Kleinberg et al.
-[28] focused on packing multiple bursty connections with varying transmission rates on a link and considered stochastic knapsack problems with known Bernoulli distribution.
-Goel and Indyk [29] studied the knapsack problem when the jobs had stochastic processing requirements or sizes.
-They proposed polynomial time approximation scheme (PTAS) when the jobs were all known to follow either Poisson or exponential distributions.
-Goyal and Ravi[30] proposed a PTAS that used a parametric linear programming reformulation and yielded near-optimal profit efficiently.
-The random variables were assumed to follow a known Gaussian distribution.
-Shabtai et al.
-[31] concentrated on the chance-constrained knapsack problem with known normal distributions.
-They presented a fast PTAS for the problem that run in poly($n/\epsilon$) time.
-Zhang et al.
-[32] studied trajectory prediction with chance constraint, assuming that future position of the moving obstacle could be represented as Gaussian distributions based on the predicted statistics.
-Chai et al.
-[33] studied trajectory optimization with chance constraint with known probability density function of random variables defined on the measurable set.
-Yang and Chakraborty [34] exploited the geometry of the non-convex feasible region of the chance-constrained knapsack problem in the variance-mean plane and transformed the problem into a sequence of deterministic knapsack problems to find an optimal solution.
-All these works mainly focused on theoretical analysis of situations that the distribution of random variables is known, which are not applicable to our problem as well.
-II-C Robust Optimization
-A robust optimization approach aims to construct an uncertainty set $\mathcal{U}_{\epsilon}\subseteq\mathcal{U}$ such that a feasible solution of the form $g(x,\xi)\leq 0,\forall\xi\in\mathcal{U}_{\epsilon}$ will also be feasible with probability at least $1-\epsilon$ with respect to $P$, i.e., $\mathcal{P}(g(x,\xi)\leq 0)\geq 1-\epsilon,\forall\xi\in\mathcal{U}$, which has the same form with chance constraint.
-Han [21] considered a chance-constrained knapsack problem where each item had an independent normally distributed random weight with a known mean and variance, and proposed an efficient pseudo-polynomial time algorithm based on the robust optimization approach.
-Joung and Lee[22] proposed a robust optimization-based heuristic algorithm where the optimal solution could be obtained by solving the ordinary binary knapsack problem iteratively.
-Distributionally robust optimization was proposed as an improvement of robust optimization when the distribution of the random variable was unknown and only observable through a finite training dataset.
-Mohajerin Esfahani and Kuhn [35] showed that distributionally robust optimization problems over Wasserstein balls could be reformulated as finite convex programs.
-Ji and Lejeune [15] studied distributionally robust chance-constrained programming optimization problems with data-driven Wasserstein ambiguity sets.
-The proposed algorithmic and reformulation framework were applied to all types of distributionally robust chance-constrained optimization problems, including the multidimensional knapsack problem.
-Additionally, Xie [16] proposed that the distributionally robust chance-constrained problem could be reformulated as a conditional value-at-risk constrained optimization problem.
-The problem with binary decision variables admitted a big-M free formulation and could be solved by a branch and cut algorithm.
-Gao and Kelywegt [36] derived a strong duality reformulation of the corresponding distributionally robust stochastic optimization problem and construct approximate worst-case distributions explicitly.
-Note that most of these work can not be applied.
-We have tried to make modifications to Ji’s [15] and Yang’s [34] algorithm to make it applicable to our problem, but the results show that both of them exhibited unsatisfactory performance.
-(see Section VI-F for details).
+Chance-constrained optimization plays a critical role in SO [24, 25, 26].
+Prior works have investigated chance-constrained optimization problems with known probability distributions, including Bernoulli, Poisson, exponential, and Gaussian [27, 28, 29, 30, 31].
+Yang and Chakraborty [19] solved the chance-constrained knapsack problem where item weights follow Gaussian distributions.
+We sought to adapt their method, referred to as the Gaussian-distribution stochastic optimization (GDSO), to address CCMCKP.
+Specifically, a statistical fitting approach is applied to sample data to obtain the mean and variance of the Gaussian distribution, based on which GDSO is run.
+However, its performance on solving CCMCKP is not satisfactory due to deviations between the assumed Gaussian distributions and the unknown underlying distributions.
+Besides, as the problem scale increases, GDSO’s runtime becomes prohibitively long, as the algorithm utilizes dynamic programming.
+The detailed results are presented in Section VI-B.
+II-C Distributionally Robust Optimization (DRO)
+Robust optimization aims to construct an uncertainty set $\mathcal{U}_{\epsilon}\subseteq\mathcal{U}$ such that a feasible solution of the form $g(x,\xi)\leq 0,\forall\xi\in\mathcal{U}_{\epsilon}$ is also feasible with probability at least $1-\epsilon$ with respect to $P$, i.e., $\mathcal{P}(g(x,\xi)\leq 0)\geq 1-\epsilon,\forall\xi\in\mathcal{U}$, which has the same form as the chance constraint.
+DRO is proposed as an improvement of robust optimization when the distributions of the random variables are unknown and only observable through sample data.
+Researchers have explored various reformulations and approaches for DRO, utilizing techniques such as Wasserstein balls, conditional value-at-risk constraints, and strong duality formulations [32, 14, 13, 33].
+Ji and Lejeune [13] studied distributionally robust multidimensional knapsack problem with Wasserstein ambiguity sets.
+We adapted their method, denoted as DRO, to address CCMCKP.
+However, experimental results show that DRO exhibits poor performance, when the sample size is not sufficiently large, or when intensity of chance constraints is high.
+Similar to GDSO, DRO also experiences a significant increase in runtime as the problem scale grows, eventually becoming impractical.
+The detailed results are presented in Section VI-B.
 III Notations and problem definition
-We give the formulation of DDCCMCKP in Eq.
+We give the formulation of CCMCKP in Eq.
 III.
 Consider $m$ mutually disjoint classes $N_{1},\ldots,N_{m}$ of items.
-Each item $j\in N_{i}$ has an associated cost $c_{ij}$ and weight $w_{ij}$, where $c_{ij},j\in N_{i},\forall i\in\mathcal{M}=\{1,\ldots,m\}$ are fixed values while $w_{ij},j\in N_{i},\forall i\in\mathcal{M}$ are unknown random variables.
-The weight of item $j\in N_{i}$ is characterized by $L$ sample data, denoted as $d_{ijl},i\in\{1,\ldots,m\},j\in N_{i},l\in\{1,\ldots,L\}$.
-Note that the sample data of each item is independent and identically distributed (i.i.d.).
+Each item $j\in N_{i}$ has an associated cost $c_{ij}$ and weight $w_{ij}$, where $c_{ij},j\in N_{i},\forall i\in\mathcal{M}=\{1,\ldots,m\}$ are fixed values while $w_{ij}$ are random variables.
 The goal is to select one item from each class to minimize the total cost, subject to keeping the total weight below the knapsack’s capacity $W$ with probability higher than the given confidence level $P_{0}$.
-By introducing the binary variables $x_{ij}$ which take on value 1 if the item $j$ is chosen in class $N_{i}$, the problem can be formulated as:
+By introducing the binary variables $x_{ij}$ which take on value 1 if item $j$ is chosen in class $N_{i}$, the problem is formulated as:
 $\min\quad$ $\sum\limits_{i=1}^{m}\sum\limits_{j\in N_{i}}c_{ij}x_{ij}$ (1a)
@@
 $x_{ij}\in\{0,1\},\forall i\in\mathcal{M},j\in N_{i}$ (1d)
-To avoid trivial solutions, we assume that the following inequality holds:
+In this work, we focus on the practical scenario where the probability distributions of random variables are unknown.
+For each item, $L$ sample data is collected, denoted as $d_{ijl},i\in{1,\ldots,m},j\in N_{i},l\in{1,\ldots,L}$, which is assumed to be independent and identically distributed (i.i.d.).
+To avoid trivial solutions, we also assume that the following inequality holds:
 $\sum\limits_{i=1}^{m}\min_{1\leq l\leq L,1\leq j\leq n_{i}}d_{ijl}\leq W\leq\sum\limits_{i=1}^{m}\max_{1\leq l\leq L,1\leq j\leq n_{i}}d_{ijl}$ (2)
-Note that MCKP is NP-hard [1].
-Supposing that all samples of each item’s weight $w_{ij}$ are the same, then DDCCMCKP reduces to traditional MCKP.
-Therefore, MCKP can be seen as a special case of DDCCMCKP, which indicates that DDCCMCKP is also NP-hard.
-It motivates us to adopt an algorithm based on a heuristic search framework to solve DDCCMCKP.
+Note that MCKP is proven to be NP-hard [1].
+Consider the scenario where item’s weights $w_{ij}$ follow a particular distribution which takes probability 1 at one value, then CCMCKP simplifies to classic MCKP.
+Consequently, MCKP can be regarded as a special case of CCMCKP, implying that CCMCKP is also NP-hard.
 IV Data-driven adaptive local search
-As mentioned in Section II, although there have been many works in the fields of MCKP, chance-constrained optimization and robust optimization, none of the methods could be directly applied to our problem.
-In order to tackle the DDCCMCKP, we present a novel approach called data-driven adaptive local search (DDALS) in this work.
-Specifically, DDALS consists of three main parts, confidence level evaluation, surrogate weight design and local search framework.
-Finally, we will analyze the complexity of the algorithm.
+As outlined in Section II, currently there is no direct research on solving CCMCKP.
+To address this problem, a novel approach termed DDALS is introduced.
+We elaborate three core components of DDALS sequentially: confidence level evaluation, surrogate weight design, and local search framework.
+Finally, a detailed analysis of the computational complexity is conducted.
 IV-A Confidence Level Evaluation
-Since the sample data is assumed to be i.i.d., it is necessary to make full use of the information contained in the sample data, since we have no information about the true distribution.
-Given a solution $S$, a natural idea is to calculate all possible combinations of the data and count the proportion of these combinations that meet the requirements.
-This proportion is an estimate of the real confidence level $P$ of the solution.
-However, obviously, a brute-force method is not applicable since it requires a calculation complexity of $L^{m}$, which is related to the problem size $m$ and sample size $L$.
-It may be feasible for small-scale problems, but it becomes impractical for large-scale problems.
-Therefore, based on the above considerations, we propose an exact evaluation method for small-scale problems that is much faster than brute-force enumeration.
-For large-scale problems, we propose an accelerated Monte Carlo evaluation method.
+As the sample data is assumed to be i.i.d., it is imperative to extract the full informational value from the sample data.
+Given a solution $S$, one approach is to calculate all possible combinations of the sample data and determine the proportion of these combinations that meets the chance constraints.
+This proportion serves as an estimate of the real confidence level of $S$.
+However, it is impractical due to its computational complexity of $L^{m}$, especially for large-scale problems.
+Therefore, we propose an exact evaluation for small-scale problems that is much faster than brute-force enumeration.
+For large-scale problems, we propose an accelerated Monte Carlo evaluation.
 IV-A1 Exact evaluation
-Exact evaluation is based on a simple idea: counting and heaps [37].
-Assume that the sample data of items has been already arranged from largest to smallest.
-Solution $S=[s_{1},s_{2},\ldots,s_{m}]$ is under evaluated.
-The details of exact evaluation are listed as follows:
+Exact evaluation is based on counting and heaps [34].
+Assume that the sample data of each item is already arranged from largest to smallest, and we have a solution $S=[s_{1},s_{2},\ldots,s_{m}]$ under evaluated.
+The details of the algorithm are listed as follows:
   • Maintain a maximum heap queue $Q$ with the length $m(1-P_{0})L^{m}$. Every time $Q$ pops the largest sum in the current heap queue [34];
@@
     $D=$ $[d_{s_{1}l_{1}},d_{s_{2}l_{2}},\ldots,d_{s_{m}l_{m}}].$
-    Then push neighbors of the popped $D$ whose sum are smaller than that of $D$ into $Q$, listed as follows:
+    Then push neighbors of the popped $D$ whose sum is smaller than that of $D$ into $Q$, listed as follows:
     $[d_{s_{1}l_{1}+1},d_{s_{2}l_{2}},\ldots,d_{s_{m}l_{m}}],$
@@
   • Repeat until the pop sum of $D$ is smaller than $W$ or the iteration times $k$ reach $\lceil(1-P_{0})L^{m}\rceil$.
-We have the following theorem,
+More importantly, we have the following theorem,
 Theorem IV.1
 The popped sum in the $k^{th}$ iteration is exactly the $k^{th}$ largest sum.
-Detailed proofs can be found in Appendix 2. If the ratio of the number to the total number of combinations is greater than $1-P_{0}$, then we determine that the solution is infeasible.
+Detailed proofs can be found online in the supplementary material. If the ratio of $k^{th}$ to the total number of combinations exceeds $1-P_{0}$, we conclude that the solution is infeasible.
 IV-A2 Accelerated Monte Carlo evaluation
-Monte Carlo simulation is a method that uses random numbers to solve many computational problems [38].
-In this work, we focus on situations where the knapsack constraints $W$ are tight at a high confidence level $P_{0}$, which means feasible solutions account for a very small proportion of the entire solution space of the problem.
-This observation motivates us that the evaluation procedure can be greatly accelerated if most of the infeasible solutions can be quickly eliminated.
-Specifically, assuming that the sample data of the items has been sorted from largest to smallest, we can check the $l^{th}$ largest data combination $d_{1l},d_{2l},...,d_{ml}$ for the combination of items to be solved, i.e.,$D_{l}=\sum_{i=1}^{m}d_{il}$, by comparing with the weight threshold $W$ to directly filter out most of the infeasible solutions.
-Therefore, the evaluations times required during the algorithm operation can be significantly reduced.
-The process can be summarized as follows.
-  • Construct a filtered list according to the minimum prime factorization for $K=L^{m}\times(1-P_{0})$.
-  • Only a limited size of combinations will be selected to construct the filtered list $R\ll m!$;
+Monte Carlo simulation uses random numbers to solve computational problems [35, 36].
+In this work, our focus is on scenarios where the knapsack constraints $W$ are tightly defined at a high confidence level $P_{0}$, indicating that feasible solutions constitute a small proportion of the entire solution space.
+It motivates us that the evaluation procedure can be greatly accelerated if most of the infeasible solutions can be quickly recognized.
+Assuming that the sample data of items has been sorted in descending order, we check the $l^{th}$ largest data combination $d_{1l},d_{2l},...,d_{ml}$, compare them with the weight threshold $W$, thereby achieving the efficient screening of infeasible solutions.
+The process is summarized as follows.
+  • Construct a screening list according to the minimum prime factorization for $K=L^{m}\times(1-P_{0})$.
+  • A limited number of combinations are selected to construct the screening list $R\ll m!$;
   • Calculate the sum of the data of each item of the screening list $R$, and check whether there is at least one combination satisfying $\sum_{i=1}^{m}d_{il}\geq W$.
-If not, the solution will be evaluated by Monte Carlo simulation.
+If not, $S$ is evaluated by Monte Carlo simulation.
 IV-B Surrogate Weight Design
-It is not feasible to directly obtain the cost-effectiveness of each item in a data-driven scenario, which impedes the algorithm’s ability to search for better solutions.
-To overcome this challenge, we construct a surrogate cost-effectiveness based on the statistic of sample data.
-Specifically, we calculate the sample mean and standard deviation of the item weights, $\mu_{ij}=\mathbb{E}(w_{ij}),\sigma^{2}_{ij}=Var(w_{ij}),i\in\{1,...,m\},j\in N_{i}$.
-Then item weights are represented by the weighted sum of statistical characteristics,
+It is infeasible to directly obtain the cost-effectiveness of each item in a data-driven scenario.
+To overcome this challenge, we construct the surrogate weight of item $\tilde{w}_{ij}$ based on sample data.
+The sample mean and sample standard deviation of item are calculated, $\mu_{ij}=\mathbb{E}(w_{ij}),\sigma^{2}_{ij}=Var(w_{ij}),i\in\{1,...,m\},j\in N_{i}$.
+The item weights are then calculated by the weighted sum of $\mu_{ij}$ and $\sigma^{2}_{ij}$,
 $\tilde{w}_{ij}:=\mu_{ij}+\lambda\sigma_{ij},\forall{i}=1,...,m,j\in N_{i}.$ (3)
-where $\lambda$ is the weighted parameter. This approach provides a statistical measure for guiding the search process towards optimal solutions.
+where $\lambda$ is the weighted parameter. This approach provides a statistical measure for constructing the surrogate weight from sample data.
 IV-C Local Search Framework
-Based on the previously data-driven solution evaluation method and surrogate weight design, we now introduce the overall algorithmic framework of DDALS.
-Local search is a widely used heuristic search framework for efficiently solving various combinatorial optimization problems [12, 39].
-Thus, our algorithm adopts the framework of local search algorithm, which can adaptively jump between small and large regions, achieving efficient search for solutions.
-The block diagram of DDALS shown in Figure 3 .
-DDALS is composed of the following five components: constructive procedure (CP) for solution initialization, a local swap search (LSS) operator for improving the current solution, a degradation (Degrade) operator for escaping local optima, and a further swap search (FSS) operator for searching in a larger feasible region.
-Since limited sample data cannot exactly reflect the true distribution, we additionally incorporate a solution feasibility enhancement (SFE) to improve the feasibility of the output.
+Local search is a widely used heuristic search framework for solving various combinatorial optimization problems [12, 37].
+We integrate local search with data-driven evaluation approach and surrogate weight design, leading to DDALS.
+The block diagram is illustrated in Figure 3.
+DDALS comprises five components: constructive procedure (CP) for solution initialization, a local swap search (LSS) operator for improving the current solution, a degradation (Degrade) operator for escaping local optima, and a further swap search (FSS) operator for searching in a larger region.
+Considering that limited sample data can not precisely mirror the true distribution, we additionally incorporate a solution feasibility enhancement (SFE) to enhance the feasibility of the output solutions.
 Figure 3: The block diagram of DDALS
-Input: A DDCCMCKP benchmark; Maximum iteration times, $MaxIter$; weighting factor, $\lambda$
+Input: Problem intance; maximum iteration times, $MaxIter$; weighting factor, $\lambda$
 Output: the optimal feasible solution $S_{opt}$
@@
 Algorithm 1 illustrates the framework of DDALS.
-The inputs of DDALS include the benchmark, maximum iteration times $MaxIter$, and a user-defined weighting factor $\lambda$.
-We first initialize the solution $S_{cp}$ using CP (line 1), then iteratively improve the quality of solutions.
-At iteration $t$, LSS attempts to obtain a better feasible solution by operating on $S_{t}$ (line 3).
-If LSS finds an improvement, the optimal solution $S^{*}$ is updated (line 4-5).
-The current solution $S_{t}$ will be degraded by randomly selecting an item for replacement to escape a local optimum (line 6).
-After iterations end, FSS is employed on $S_{opt}$ to search for a better solution (line 8).
-The final solution $S_{opt}$ will be output by SFE.
-The detailed designs of the components are list as follows:
+The inputs for DDALS include the problem instances, maximum iteration times $MaxIter$, and a pre-defined weighting factor $\lambda$.
+DDALS first initializes the solution $S_{cp}$ using CP (line 1) and iteratively enhances the quality of solutions.
+At iteration $t$, LSS attempts to improve the feasible solution by operating on $S_{t}$ (line 3).
+If LSS identifies an improvement, it updates the optimal solution $S^{*}$ (lines 4-5).
+The current solution $S_{t}$ is degraded by randomly selecting an item for replacement to escape a local optimum (line 6).
+Once iterations end, FSS is applied to search for a better solution (line 8).
+The final solution $S_{opt}$ is output by SFE.
+The detailed designs of these components are list as follows:
 IV-C1 Constructive procedure
 The initialization procedure is based on greedy strategy [7].
-Utilizing the item weights defined in Definition 3, we calculate the cost-effective variables $u_{ij}=v_{ij}/\tilde{w}_{ij}$ for each item $j\in N_{i}$, then sort the items in descending order of $u_{ij}$ for each class $N_{i}$.
+Utilizing the surrogate weights defined in Definition 3, we calculate the cost-effectiveness variables $u_{ij}=v_{ij}/\tilde{w}_{ij}$ for each item $j\in N_{i}$, then sort the items in descending order of $u_{ij}$ for each class $N_{i}$.
 $S_{cp}$ is constructed by selecting the item with the largest $u_{ij}$ in each class.
-If the current $S_{t}$ is infeasible, the item which has the largest $\tilde{w}_{i_{0}j}$ in $S_{t}$ is replaced by an item with a smaller weight in the same class $i_{0}$.
-The procedure iterates until a feasible $S_{cp}$ is obtained.
-If all items in class $i_{0}$ are tried but still infeasible,$S_{cp}$ will be constructed by the smallest $\tilde{w}_{ij}$ in each class.
+If the current $S_{t}$ is infeasible, the item which has the largest $\tilde{w}_{i_{0}j}$ in $S_{t}$ is replaced by the one with a smaller weight in the same class $i_{0}$.
+The procedure iterates until feasible $S_{cp}$ is obtained.
+If all items in class $i_{0}$ have been tried but $S_{cp}$ is still infeasible, $S_{cp}$ will be constructed by the items with the smallest surrogate weights.
 IV-C2 Local swap search
-LSS applies a swapping strategy for picked items to improve the current solution $S_{t}$.
-If the obtained solution is feasible, each swap is authorized.
-LSS will try all the items with lower cost in each class $N_{i}$ and selects the item with the lowest cost from all authorized swaps.
+LSS applies a swapping strategy to improve the current solution $S_{t}$.
+If the obtained solution is feasible, the swap operation is authorized.
+LSS aims to select items with the smallest cost while ensuring that the constructed solution is feasible.
 IV-C3 Degrade
 Degrade randomly replaces an item $j_{i_{0}}$ of class $i_{0}$ in the current solution $S_{t}$.
-If an infeasible solution is generated, it randomly selects a lighter item in $i_{0}$.
-If there is no suitable item in the $i_{0}$, it randomly switches to another class and tries again until a new feasible solution $S_{t}$ is obtained.
+If an infeasible solution is constructed, it randomly selects an item with smaller surrogate weight in class $i_{0}$.
+If there is no suitable item, Degrade randomly selects another class and tries again until a new feasible solution $S_{t}$ is obtained.
 IV-C4 Further swap search
-FSS iterates through all class combinations $(i_{0},i_{1})$ of the current optimal solution $S_{t}$ to find a better and feasible solution.
-If it finds a pair of items with a smaller cost, $S_{t}$ will be updated.
-The algorithm ends when all pairs of items have been searched.
-IV-C5 Solution feasibility enhancement
-In practice, due to the demand for high confidence level $P_{0}$, we tend to minimize the cost of the solution while ensuring its feasibility, rather than blindly pursuing the minimization of the cost, leading to solutions that do not satisfy constraints, which renders them useless in practice.
-It is a crucial requirement in many practical applications.
-However, it is difficult to guarantee the real feasibility of the optimal solution due to limited sample size.
-Note that the sample size to achieve smaller errors for arbitrarily distributed random variables will be very large (See Appendix 3 for theoretical analysis between estimation error and sample size).
-Therefore, we design three filtering variants to enhance the feasibility of output solution, while not increasing the cost too much.
-In the iterations of DDALS algorithm, two list of feasible solutions are recorded, each with a capacity of 30.
-The first records 30 feasible solutions with the lowest cost encountered so far, named Cost List.
-The second records 30 solutions with the largest confidence level encountered so far, named MC List.
-  • Original: Output the optimal solution of FSS.
-  • Variant 1: The minimum-cost solution with a higher confidence level than $P_{0}$. For example, it should satisfy $P\geq 0.995$ if $P_{0}=0.99$.
-  • Variant 2: All solutions will be scored according to the order in the list they are in.
-Each candidate solution has two scores, which will be added together.
-Choose the solution with the highest score as the final output.
-  • Variant 3: Take the first 10 solution (if the amount is larger than 10) in Cost List as the final output and choose one manually.
+FSS iterates through all the item combinations $(i_{0},i_{1})$ from two different classes of the current optimal solution $S_{t}$.
+When a pair of items with smaller total cost is found, $S_{t}$ is updated.
+The procedure terminates when all possible pairs of items have been tried.
+IV-C5 Solution feasibility enhancement (SFE)
+In practical scenarios, we aim to minimize costs while ensuring that the solution is feasible.
+Notably, achieving smaller errors for arbitrarily distributed random variables necessitates an exceedingly large sample size (see the online supplementary material for theoretical analysis).
+Due to limited sample size, instead of blindly minimizing costs, by allowing for a certain margin of error, the probability of obtaining a solution that is truly feasible in practical scenarios is larger.
+Therefore, we designed the SFE.
+When DDALS starts running, SFE maintains two lists, each recording up to 30 feasible solutions.
+The first list is called MC list, which records the 30 solutions with the highest confidence levels obtained during the DDALS search process.
+The second list is called Cost list, which records the 30 lowest-cost feasible solutions obtained throughout the whole operation of DDALS.
+  • Original: DDALS outputs the optimal solution without enhancement.
+  • Variant 1: The minimum-cost solution with a higher confidence level. For example, $S_{t}$ should satisfy that the confidence level is greater than 0.995 when $P_{0}=0.99$.
+  • Variant 2: All solutions are scored within their respective lists, and the scores in both lists are added together to obtain a total score.
+The solution with the highest total score is selected as the output.
 IV-D Complexity Analysis
 IV-D1 Exact evaluation
-First, if the solution can be directly determined infeasible in the fast screening phase, the time complexity takes $O(1)$.
-Otherwise, in order to verify whether it is feasible, the maximum number of iterations that the algorithm needs to perform is $T=L^{m}(1-P_{0})-l_{min}^{m}$, where $l_{min}$ is the index that the combination of $l_{min}$ samples of solution is larger than $T_{\max}$.
-The maximum heap size is $k_{0}+(m-1)T=\sum_{i=1}^{m}\binom{m}{i}(l_{min}+1)^{m-i}+(m-1)L^{m}(1-P_{0})-(m-1)l_{min}^{m}=O(mT)$.
-Since at each iteration it needs to sort the heap once, the total number of times is $O(\log(mT))=O(\log m+\log T)=O(\log T),m\ll L$.
-Thus, the total time complexity of the filtered algorithm is $O(mL\log L+\sum_{i=1}^{m}\binom{m}{i}(l_{min}+1)^{m-i}+T\log T)=O(\max\{T\log T,ml_{min}^{m-1},m^{2}l_{min}^{m-2},...,ml_{min}\})$.
+When $S$ is fast recognized as infeasible solutions, it takes $O(1)$.
+Otherwise, in order to verify whether $S$ is feasible, the maximum number of iterations that the evaluation process is $T=L^{m}(1-P_{0})-l_{min}^{m}$, where $l_{min}$ is the index that the combination of $l_{min}$ samples of solution is larger than $T_{\max}$.
+The maximum size of heap is $k_{0}+(m-1)T=\sum_{i=1}^{m}\binom{m}{i}(l_{min}+1)^{m-i}+(m-1)L^{m}(1-P_{0})-(m-1)l_{min}^{m}=O(mT)$.
+At each iteration, the evaluation approach needs to sort the heap once, so the total number of evaluation times is $O(\log(mT))=O(\log m+\log T)=O(\log T),m\ll L$.
+The total time complexity is $O(mL\log L+\sum_{i=1}^{m}\binom{m}{i}(l_{min}+1)^{m-i}+T\log T)=O(\max\{T\log T,ml_{min}^{m-1},m^{2}l_{min}^{m-2},...,ml_{min}\})$.
 IV-D2 Accelerated Monte Carlo evaluation
-The complexity of the accelerated Monte Carlo evaluation is $O((|G|+\alpha\times MonteCarlo)\times Eval)$, where $MonteCarlo$ denotes the set number of Monte Carlo simulations and $|G|$ denotes the size of fast screening combination.
-If the solution is an infeasible solution determined by fast screening, $\alpha=0$, otherwise $\alpha=1$.
+The complexity of the accelerated Monte Carlo evaluation is $O((|G|+\alpha\times MonteCarlo)\times Eval)$, where $MonteCarlo$ denotes the simulation times of Monte Carlo simulations and $|G|$ is the size of the filtered list.
+If $S$ is infeasible recognized by filtered list, $\alpha=0$, otherwise $\alpha=1$.
 IV-D3 DDALS
-  • CP includes sorting and solution evaluation. Up to $m\times N$ times of evaluation are required. The complexity is $O(Eval\times mN)$, where $N=\max|N_{i}|$.
-  • Similar to CP, LSS require up to $m\times N$ times of evaluation to search a possible better solution. The complexity is $O(Eval\times mN)$.
-  • The worst case is that only one combination of items is feasible in the whole solution space. The complexity of Degrade is $O(Eval\times mN)$.
-  • FSS iterates through a maximum of combinations of classes $\binom{m}{2}\times N^{2}$. The complexity of FSS is $O(Eval\times m(m-1)N^{2}/2)=O(Eval\times m^{2}N^{2})$.
-Note that SFH only takes $O(1)$.
+  • CP includes sorting and solution evaluation. The times of evaluation are up to $m\times N$ Thus, the complexity is $O(Eval\times mN)$, where $N=\max|N_{i}|$.
+  • LSS requires up to $m\times N$ times of evaluation to search for a better solution. The complexity is $O(Eval\times mN)$.
+  • For Degrade procedure, the worst case is that only one combination of items is feasible. Thus, the complexity is $O(Eval\times mN)$.
+  • FSS iterates through a maximum of combinations of classes $\binom{m}{2}\times N^{2}$. The complexity is $O(Eval\times m(m-1)N^{2}/2)=O(Eval\times m^{2}N^{2})$.
+Note that SFE only takes $O(1)$.
 Therefore, the total time complexity of DDALS is $O(Eval\times mN+MaxIter\times(Eval\times mN+Eval\times mN)+Eval\times N^{2}m^{2}+1)=O(Eval\times\max\{N^{2}m^{2},MaxIter\times mN\})$.
 V Benchmark Design
-Since DDCCMCKP is a new problem that has not been solved before, it is necessary to design benchmarks to evaluate the performance of algorithms in different scenarios.
-Therefore, we designed two sets of benchmarks to cover a wide range of problem sizes and complexities, each of which consists of multiple benchmarks of different scales, in order to comprehensively evaluate the algorithm’s performance in different scenarios.
+As CCMCKP has not been studied before, there is a lack of corresponding benchmarks.
+To evaluate the performance of algorithms for solving CCMCKP, we designed two sets of benchmark with different problem scales and the intensity of chance constraints.
 (a) LAB
@@
 Figure 4: Probability density function of items. The horizontal axis represents the time delay and the vertical axis represents the probability density.
-The first set of benchmarks aims to simulate the case where the delay of items follows different common continuous distributions, named LAB.
-The distribution of each item is randomly chosen from the set of uniform, truncated normal, fatigue-life, bimodal and gamma distributions.
-The mean of items is uniformly distributed in $U(2,8)$, i.e., $\mu\sim U(2,8)$, and the variance is uniformly distributed in $U(1,19)$, i.e., $\sigma^{2}\sim U(1,19)$, where $U(2,8)$ denotes a uniform distribution over the interval [2,8].
-In the case of the gamma distribution, the mean value is $\mu\sim U(0.5,2.5)$ and variance $\sigma^{2}\sim U(0.05,0.625)$.
-The cost, representing the value of taking the item, is generated randomly between $U(1,10)$, corresponding to uncorrelated situation in classic MCKP benchmark[26], which is a widely used benchmark in the literature.
-The curve of probability density function of different distributions of the items involved in LAB is shown in Figure.4a.
-The second set of benchmarks simulates the actual business requirements of a telecommunication company in the 5G E2E network, denoted as APP.
-In APP, the distribution of each item is also randomly chosen in the set of uniform, truncated normal, fatigue-life, bimodal and gamma distribution, which is similar to the LAB case.
+The first set aims to simulate the scenarios where the delay of items follows different common continuous distributions, named LAB.
+The distributions of items are randomly chosen from uniform, truncated normal, fatigue-life, bimodal and gamma distributions.
+The mean of items is uniformly distributed in $U(2,8)$, where $U(2,8)$ denotes a uniform distribution over the interval [2,8].
+In the case of gamma distribution, the mean value is $\mu\sim U(0.5,2.5)$ and the variance is $\sigma^{2}\sim U(0.05,0.625)$.
+The cost, representing the value associated with taking each item, is randomly generated from $U(1,10)$.
+This corresponding to an uncorrelated situation in the widely used benchmark of classic MCKP [22].
+The probability density functions for different distributions in LAB are shown in Figure 4a.
+The second set, denoted as APP, is designed to simulate the real-world business requirements of the 5G E2E network from a certain telecommunications company.
+In APP, the distributions of item are also randomly chosen in the set of uniform, truncated normal, fatigue-life, bimodal and gamma distribution.
 However, in the 5G E2E network, each data package transmission is restricted to a time window with a fixed size, indicating that packets must be transmitted successfully within a set interval.
-The difference between APP and LAB lies in the introduction of the retransmission mechanism [17].
-This mechanism implies that each item in all classes has a probability of being retransmitted when a transmission failure occurs.
-In the APP, the first data package transmission attempt must be completed within $(0,10]$ ms with a success probability of 0.9.
-If the first attempt fails, a second attempt occurs with a delay of $(10,20]$ ms and a success probability of 0.9.
-Therefore, the delay of one transmission ranges from $(10,20]$ ms with a probability of $0.1\times 0.9=0.09$.
-The same applies to the third and fourth retransmissions.
-The transmission fails if an item requires transmitting data more than four times, which limits the maximum number of retransmissions.
-Consequently, the time delay for an item to transmit a data package belongs to the intervals $(0,10],(10,20],(20,30],(30,40]$ with probabilities of $0.9,0.09,0.009,0.001$, respectively.
-The distribution of items involved in the APP is presented in Figure 4b.
-The cost generation formula for items is $c_{ij}=10/(\mu_{ij}+\sigma_{ij})\times U(0.8,1.2)$.
-The purpose is to make the cost-effectiveness decrease as the delay decreases, thus simulating the marginal diminishing effect of cost-effectiveness in practical application.
-For each benchmark in the LAB and APP, we design small-scale and large-scale problems with different parameter settings, as shown in Table I and II respectively.
-The scale of benchmarks, i.e., $m$ and $N$, is set similar to [26].
-The sample sizes $L$ are set 30 for small-scale benchmarks and 500 for large-scale benchmarks.
-In Table I, ”{11,14}” of $T_{\max}^{LAB}$ means that $T_{\max}$ is set 11 or 14 as two different settings in the experiments of ss1.
-TABLE I: Small-scale benchmarks of DDCCMCKP
-Benchmark m N L $T_{\max}^{LAB}$ $T_{\max}^{APP}$
+This distinctive feature is known as retransmission mechanism [15].
+Retransmission mechanism implies that each item in all classes has a probability of being retransmitted when a transmission failure occurs.
+In APP, the first attempt to transmit a data package must be completed within $(0,10]$ ms with a success probability of 0.9.
+If the initial attempt fails, the second attempt occurs with a delay of $(10,20]$ ms and a success probability of 0.9.
+This pattern continues for the third and fourth retransmissions.
+Finally, the transmission is recognized as a failure if the attempt is more than four times.
+The time delay for an item to transmit a data package falls within the intervals $(0,10]$, $(10,20]$, $(20,30]$, $(30,40]$ with probabilities of $0.9,0.09,0.009,0.001$, respectively.
+The distributions of items involved in APP are illustrated in Figure 4b.
+The cost generation formula is defined as $c_{ij}=10/(\mu_{ij}+\sigma_{ij})\times U(0.8,1.2)$.
+This formulation is designed to induce a decrease in cost-effectiveness as the time delay decreases, simulating the marginal diminishing effect of cost-effectiveness in practical scenarios.
+For instances in LAB and APP, we set the values of $m$ and $N$ similar to [22].
+Additionally, we categorize them into small-scale and large-scale instances.
+The sample size is set as 30 for small-scale instances and 500 for large-scale instances.
+The settings of all instances are shown in Table I.
+“11,14” of $T_{\max}^{LAB}$ means that $T_{\max}$ of LAB is set 11 or 14 as two different intensity of chance constraints.
+TABLE I: Instances setting
+Instance $m$ $N$ $L$ $T_{\max}^{LAB}$ $T_{\max}^{APP}$
 $ss1$ 3 5 30 {11,14} {21,27}
@@
 $ss4$ 5 10 30 {10,16} {16,27}
-TABLE II: Large-scale benchmarks of DDCCMCKP
-Benchmark m N L $T_{\max}^{LAB}$ $T_{\max}^{APP}$
 $ls1$ 10 10 500 {19,23} {32,37}
@@
 VI Computational Study
-Experiments are carefully designed to evaluate the performance of DDALS in solving DDCCMCKP.
-We also conducted ablation studies, which involves disabling individual components of DDALS, to assess the contribution of the components integrated in DDALS.
-The source code of algorithms and the benchmark sets, are available at Github  [^1] Github: github.com/eddylxf23/DDCCMCKP.
-All experiments were run in python on an Intel(R) Xeon(R) CPU E5-2699A v4(2.40GHz,252G RAM) with Ubuntu 16.04 LST.
-VI-A Algorithm Settings
-We set different parameter combinations of weighted parameter $\lambda$ and iteration times $t$ to evaluate the performance on small-scale and large-scale DDCCMCKP respectively.
-The performance is evaluated by the cost of final solution when the algorithm reaches the maximum of iteration, and the experiment was repeated for 10 times and take the average value.
-The value of $\lambda$ is tested from 1 to 10.
-As show in Fig.
-5, there are several points worth noting.
-In the experiments of LAB-ss2, when $T_{\max}=18$, the current solution of the algorithm remains unchanged.
+In the experiments, we first analyze the sensitivity of different parameter settings of DDALS.
+Secondly, we adapt the existing methods, GDSO and DRO, to address our problem and compare with DDALS.
+Thirdly, we further conduct experiments to compare and analyze various commonly used search methods, such as greedy, estimation of distribution algorithm (EDA), and genetic algorithm (GA), combined with solution evaluation approaches against DDALS.
+Furthermore, we analyze the impact of the integrated acceleration mechanism, evaluating how it enhances the overall efficiency.
+Finally, we conduct ablation studies, which involves disabling individual components of DDALS, to assess the contribution of the components integrated in DDALS.
+The source code, benchmark sets and supplementary materials are available online  [^1] Github: github.com/eddylxf23/CCMCKP.
+All experiments were executed in python on an Intel(R) Xeon(R) CPU E5-2699A v4(2.40GHz,252G RAM) with Ubuntu 16.04 LST.
+VI-A Sensitivity Analysis and Parameter Setting
+Various parameter combinations of the weighted parameter $\lambda$ and iteration times $t$ to assess the algorithm’s performance are employed.
+The performance metric is the cost of the final solution when LSS reaches its maximum iteration times.
+The experiments are repeated 10 times, and the average value is computed.
+The parameter $\lambda$ is varied from 1 to 10.
+The results are shown in Figure 5.
+In the experiments of LAB-ss2, when $T_{\max}=18$, the current solution remains unchanged.
 The results show that the lowest cost was obtained when $\lambda=0$.
-However, when we validate the output solution with real confidence level, we find that the solution was infeasible, which means that the algorithm could not find feasible solutions based on the samples without considering the standard deviation.
-When $T_{\max}=26$, different lambdas do not affect the algorithm’s convergence to the lowest cost solution.
-In the experiments of LAB-ls2, when $T_{\max}=25$, the algorithm converges very slowly when we set $\lambda=0$, while other settings can converge quickly.
-However, when $T_{\max}=32$, different lambda values have little effect on the algorithm.
-That’s because the constraints are relatively loose, and feasible solutions account for a large proportion of the entire solution space, making the effect of standard deviation less significant.
-In general, it is necessary to set a non-zero value for $\lambda$, and our algorithm is robust to different $\lambda$, with $t=30$ being appropriate.
+However, it is revealed that the solution is infeasible, indicating that the algorithm could not find feasible solutions without considering the standard deviation.
+For $T_{\max}=26$, different values of $\lambda$ had no impact on the algorithm’s convergence to the lowest cost solution.
+In the experiments of LAB-ls2, with $T_{\max}=25$, the algorithm exhibited slow convergence when $\lambda=0$, while other settings allowed for quicker convergence.
+When $T_{\max}=32$, different $\lambda$ settings had small impact on the performance since the intensity of chance constraints are low, making the effect of standard deviation less significant.
+In summary, it is necessary to set a non-zero value for $\lambda$, and our algorithm exhibits robustness across different $\lambda$ values.
+Meanwhile, $t=30$ is an appropriate choice for iteration times.
 (a) LAB-ss2, $T_{\max}=18$, exact
@@
 (d) LAB-ls3, $T_{\max}=32$, MC
-Figure 5: Performance analysis of different $\lambda$.
-As the algorithm iterates, lower cost can be achieved.
-The results show that the surrogate weight is necessary for DDALS while a non-zero $\lambda$ is necessary to the algorithm.
-VI-B Performance of Evaluation Methods
-Since each candidate solution can be seen as an unknown random distribution, concentration inequalities can serve as a reference to evaluate the quality of these candidate solutions.
-We utilize two kinds of concentration inequality as the lower bounds: Bernstein’s Inequality and $K$-order Hoeffding’s Inequality, which have been widely used in the literature for evaluating the performance of optimization algorithms (Please refer to Appendix 1 for detailed formulas).
-TABLE III: Evaluation for LAB-ss2
+Figure 5: Performance analysis of different $\lambda$. The results show that the surrogate weight with a non-zero $\lambda$ is necessary for DDALS.
+TABLE II: Evaluation on LAB-ss2
 […]
 Real 68.24 89.41 99.51 99.98 100.00
-B’s inequality 23.22 49.16 75.63 89.87 95.85
-H’s inequality 22.83 42.20 62.32 74.17 85.12
 Exact 81.64 93.90 99.53 99.99 100.00
-We test exact method on LAB-ss2, and calculate the real confidence level for solution $S_{1},S_{2},S_{3},S_{4},S_{5}$ through the original generator, denoted as ‘Real’.
-‘B’s’ and ‘H’s’ represent Bernstein’s and Hoeffding’s inequality, respectively.
-From Table III, we found a substantial gap between the evaluation outcome and the real confidence level, indicating that the real distribution of solution of DDCCMCKP is far different from Gaussian distribution.
-Results show that the estimate of the exact method is closest to Real, and the estimation error becomes smaller as the real confidence level increases.
-TABLE IV: Evaluation for APP-ls1
+TABLE III: Evaluation on APP-ls1
 […]
 Real 79.18 90.70 94.48 98.78 98.81
-B’s inequality 9.56 11.56 19.82 28.09 48.87
-H’s inequality 12.58 25.38 38.30 41.98 53.96
 MC(10e2) 81.1$\pm$ 3.78 90.7$\pm$ 2.87 95.8$\pm$ 2.15 99.0$\pm$ 0.82 98.3$\pm$ 1.16
@@
 MC(10e6) 79.13$\pm$ 0.05 90.38$\pm$ 0.02 95.10$\pm$ 0.03 99.21$\pm$ 0.00 99.41$\pm$ 0.01
-Next, we test Monte Carlo evaluation on APP-ls1 in Table IV.
-‘MC(10e2)’ represents that the evaluation is carried by 10e2 times of Monte Carlo estimation.
-The result is the average of 10 times’ results.
-‘81.1 $\pm$ 3.78’ represents that the mean of estimated confidence level is 81.1% and the standard deviation is 3.78%.
-The result shows that the estimation error of $210e$ 2 times has the largest fluctuations and $510e$ 5 tends to be close to the real confidence value, but still not enough to meet the requirement for larger scale of problem and higher confidence level.
+In order to assess the performance of the exact evaluation approach, we calculate the real confidence level for specific solutions (denoted as “Real”) of LAB-ss2, namely $S_{1},S_{2},S_{3},S_{4},S_{5}$ through the original generator.
+Results in Table II show that the estimation error becomes smaller as the real confidence level increases.
+Additionally, we conduct Monte Carlo evaluation on APP-ls1 as shown in Table III.
+“MC(10e2)” represents that the simulation times is set 10e2.
+The experimental results are the averages of 10 trials.
+“81.1 $\pm$ 3.78” represents that the mean of estimated confidence level is 81.1% and the standard deviation is 3.78%.
+The results show that the estimation error of $210e$ 2 times has the largest fluctuations and $510e$ 5 tends to be close to the real confidence level, but still not enough to meet the requirement for larger problem scale and higher confidence level.
 Considering the trade-off between accuracy and speed, we set $610e$ 6 as the Monte Carlo simulation times.
-VI-C Results on Small-Scale and Large-Scale DDCCMCKP
-To evaluate the effectiveness of DDALS, we compare DDALS with several baseline methods on various benchmark sets from the real-world application.
-The baseline methods include Greedy, GA and EDA.
-Greedy typically works well on problems where the locally optimal choice leads to an overall optimal solution.
-GA performs well in solving different optimization problems [40].
-While EDA achieves great success in both combinatorial and continuous optimization [41].
-For fairness, we first run DDALS then record the value of ET.
-Next, we let the Greedy, GA, and EDA keep running until convergence or the ET of which exceeds that of DDALS, then output the current solution as the result.
-The population size is set to 10 and the number of dominant individuals is 6 for both GA and EDA.
-For GA, the crossover probability of parents is set to 0.1 and the probability of variation is $1/m$.
-The real confidence level is calculated by using the original generators to generate $710e$ 7 data.
-Note that a solution is regarded as feasible when its real confidence level is larger than $0.985$, which means that we allow for an error margin of $0.5\%$.
-TABLE V: Performance on benchmarks of LAB. The highest average real confidence level and the lowest cost are highlighted through bolding and being underlined.
-Benchmark small $T_{\max}$ large $T_{\max}$
-C ET ECL RCL FSR C ET ECL RCL FSR
-ss3-Greedy 24.4$\pm$0.0 2.0$\pm$0.0 100$\pm$0 99.9$\pm$0.0 1 21.8$\pm$0.0 3.0$\pm$0.0 100$\pm$0 99.9$\pm$0.0 1
-ss3-GA 22.6$\pm$1.5 303.4$\pm$2.8 99.9$\pm$0.0 99.7$\pm$0.2 1 12.9$\pm$0.7 262.4$\pm$1.6 99.9$\pm$0.0 99.9$\pm$0.1 1
-ss3-EDA 24.1$\pm$2.3 304.1$\pm$2.3 99.9$\pm$0.1 99.6$\pm$0.6 0.9 14.0$\pm$2.0 264.5$\pm$2.9 99.9$\pm$0.0 99.9$\pm$0.1 1
-ss3-DDALS(O) 21.4$\pm$0.0 291.9$\pm$6.2 99.9$\pm$0 99.5$\pm$0.0 1 12.6$\pm$0.0 254.4$\pm$4.7 100$\pm$0 99.9$\pm$0 1
-ss3-DDALS(V1) 21.4$\pm$0.0 291.9$\pm$6.2 99.9$\pm$0 99.5$\pm$0.0 1 12.6$\pm$0.0 254.4$\pm$4.7 100$\pm$0 99.9$\pm$0 1
-ss3-DDALS(V2) 21.4$\pm$0.0 291.9$\pm$6.2 99.9$\pm$0 99.5$\pm$0.0 1 12.6$\pm$0.0 254.4$\pm$4.7 100$\pm$0 99.9$\pm$0 1
-ss3-DDALS(V3) 23.9$\pm$1.7 279 99.8$\pm$0.2 99.1$\pm$0.8 0.8 15.7$\pm$1.9 260 99.9$\pm$0.3 99.6$\pm$0.9 0.9
-ss4-Greedy 30.1$\pm$0.0 2.0$\pm$0.0 99.9$\pm$0.0 99.9$\pm$0.0 1 27.5$\pm$0.0 3.0$\pm$0.0 99.7$\pm$0.0 99.4$\pm$0.0 1
-ss4-GA 23.8$\pm$5.5 768.9$\pm$8.5 99.9$\pm$0.3 99.7$\pm$0.4 1 17.2$\pm$3.0 562.6$\pm$5.7 99.7$\pm$0.3 99.6$\pm$0.3 1
-ss4-EDA 17.4$\pm$0.0 763.4$\pm$2.5 99.9$\pm$0.0 99.9$\pm$0.0 1 13.0$\pm$1.0 553.4$\pm$2.7 99.8$\pm$0.2 98.9$\pm$0.4 1
-ss4-DDALS(O) 17.4$\pm$0.0 759.9$\pm$25.8 99.9$\pm$0.0 99.9$\pm$0.0 1 11.2$\pm$0.0 525.3$\pm$39.1 99.2$\pm$0.0 99.0$\pm$0.0 1
-ss4-DDALS(V1) 17.4$\pm$0.0 759.9$\pm$25.8 99.9$\pm$0.0 99.9$\pm$0.0 1 12.1$\pm$0.0 525.3$\pm$39.1 99.5$\pm$0.0 98.9$\pm$0.0 1
-ss4-DDALS(V2) 17.4$\pm$0.0 759.9$\pm$25.8 99.9$\pm$0.0 99.9$\pm$0.0 1 14.0$\pm$1.2 525.3$\pm$39.1 99.9$\pm$0.2 99.0$\pm$0.3 1
-ss4-DDALS(V3) 24.6$\pm$4.0 735 99.8$\pm$0.2 99.5$\pm$0.5 1 13.8$\pm$1.3 561 99.8$\pm$0.3 99.5$\pm$0.5 1
-ls4-Greedy 169.7$\pm$0.0 5$\pm$0.0 99.8$\pm$0.0 99.8$\pm$0.0 1 162.5$\pm$0.0 6.0$\pm$0.0 99.9$\pm$0.0 99.9$\pm$0.0 1
-ls4-GA 123.1$\pm$8.8 11307.2$\pm$8.0 99.1$\pm$0.1 99.1$\pm$0.2 1 104.1$\pm$5.1 8409.1$\pm$5.6 99.1$\pm$0.2 99.2$\pm$0.2 1
-ls4-EDA 126.2$\pm$2.3 11304.8$\pm$2.1 99.2$\pm$0.2 99.4$\pm$0.2 1 114.8$\pm$4.5 8403.4$\pm$2.8 99.6$\pm$0.2 99.6$\pm$0.2 1
-ls4-DDALS(O) 106.6$\pm$0.8 11289.9$\pm$92.0 99.1$\pm$0.1 99.2$\pm$0.1 1 92.8$\pm$0.7 8376.6$\pm$378.8 99.1$\pm$0.1 99.2$\pm$0.1 1
-ls4-DDALS(V1) 109.5$\pm$0.4 11289.9$\pm$92.0 99.7$\pm$0.1 99.7$\pm$0.1 1 97.9$\pm$2.2 8376.6$\pm$378.8 99.6$\pm$0.1 99.7$\pm$0.1 1
-ls4-DDALS(V2) 108.9$\pm$1.3 11289.9$\pm$92.0 99.6$\pm$0.2 99.7$\pm$0.2 1 93.8$\pm$1.4 8376.6$\pm$378.8 99.3$\pm$0.2 99.4$\pm$0.2 1
-ls4-DDALS(V3) 108.7$\pm$1.1 11171 99.2$\pm$0.2 99.3$\pm$0.2 1 96.5$\pm$1.7 8177 99.1$\pm$0.1 99.2$\pm$0.1 1
-The results on LAB benchmarks are presented in Table V.
-‘ss3-Greedy’ in the left of table indicates the result of Greedy tested on benchmark ss3.
-‘DDALS(O)’, ‘DDALS(V1)’, ‘DDALS(V2)’ and ‘DDALS(V3)’ represents the original, first, second and third version of DDALS.
-‘small $T_{\max}$’ and ‘large $T_{\max}$’ are corresponding to small and large $T_{\max}$ in Table I and Table II.
-Five indicators are utilized to measure the performance.
-The cost of solution is denoted as ‘C’.
-The evaluation times ‘ET’ is utilized to compare the convergence speed of algorithms.
-‘ECL’ indicates estimated confidence level, while ‘RCL’ is calculated by real generator.
-‘FSR’ represents the ratio of actual feasible solution in the output results of the 10 experiments.
-All results are based on statistical analysis from 10 repetitions of the experiment, denoted as ‘mean $\pm$ standard deviation’.
-Due to space limitation, parts of results are included in the main text, and the complete results are available in Appendix 5.
-TABLE VI: Performance on benchmarks of APP. Due to the introduction of retransmission mechanisms, there is a significant difference in the results between LAB and APP.
-Benchmark small $T_{\max}$ large $T_{\max}$
-C ET ECL RCL FSR C ET ECL RCL FSR
-ss2-Greedy 6.4$\pm$0.0 3$\pm$0.0 99.7$\pm$0.0 98.6$\pm$0.0 1 6.4$\pm$0.0 3.0$\pm$0.0 99.8$\pm$0.0 99.0$\pm$0.0 1
-ss2-GA 5.3$\pm$0.0 301.8$\pm$2.2 99.2$\pm$0.0 97.3$\pm$0.4 0 5.3$\pm$0.0 272.6$\pm$2.8 99.5$\pm$0.1 98.1$\pm$0.2 0
-ss2-EDA 5.4$\pm$0.2 303.1$\pm$3.0 99.3$\pm$0.2 97.4$\pm$0.7 0.1 5.4$\pm$0.2 274.9$\pm$3.0 99.6$\pm$0.2 98.3$\pm$0.4 0.3
-ss2-DDALS(O) 5.2$\pm$0.0 307$\pm$7.6 99.2$\pm$0.0 97.3$\pm$0 0 5.2$\pm$0.0 291$\pm$10.6 99.6$\pm$0.0 98.1$\pm$0.0 0
-ss2-DDALS(V1) 5.5$\pm$0.1 307$\pm$7.6 99.8$\pm$0.0 97.5$\pm$0.3 0 5.2$\pm$0 291$\pm$10.1 99.6$\pm$0.0 98.1$\pm$0.0 0
-ss2-DDALS(V2) 5.5$\pm$0.1 307$\pm$ 7.6 99.8$\pm$0.0 97.5$\pm$0.3 0 5.4$\pm$0.1 291$\pm$10.1 99.9$\pm$0.1 98.2$\pm$0.1 0
-ss2-DDALS(V3) 5.4$\pm$0.1 295 99.4$\pm$0.3 97.7$\pm$0.4 0 5.3$\pm$0.1 272.6$\pm$2.8 99.5$\pm$0.1 98.1$\pm$0.2 0
-ss4-Greedy 34.9$\pm$0.0 3.0$\pm$0.0 100.0$\pm$0.0 100.0$\pm$0.0 1 27.2$\pm$0.0 4.0$\pm$0.0 100.0$\pm$0.0 99.0$\pm$0.0 1
-ss4-GA 27.3$\pm$1.1 2010$\pm$8.4 99.8$\pm$0.3 90.4$\pm$0.5 0 21.5$\pm$2.9 1614.8$\pm$9.8 99.6$\pm$0.3 93.2$\pm$4.8 0
-ss4-EDA 26.1$\pm$0.0 2005.7$\pm$2.3 100$\pm$0 90.1$\pm$0 0 16.4$\pm$0.0 1602.7$\pm$2.1 99.9$\pm$0.0 80.3$\pm$0.0 0
-ss4-DDALS(O) 26.1$\pm$0.0 1988.2$\pm$30.4 1$\pm$0 90.1$\pm$0 0 16.7$\pm$0.9 1579.4$\pm$76.8 99.8$\pm$0.2 81.8$\pm$4.4 0
-ss4-DDALS(V1) 26.1$\pm$0.0 1988.2$\pm$30.4 1$\pm$0 90.1$\pm$0 0 16.7$\pm$1.0 1579.4$\pm$ 76.8 99.9$\pm$0.0 81.5$\pm$3.6 0
-ss4-DDALS(V2) 26.1$\pm$0.0 1988.2$\pm$30.4 1$\pm$0 90.1$\pm$0 0 18.0$\pm$1.9 1579.4$\pm$76.8 99.9$\pm$0.1 86.1$\pm$7.0 0
-ss4-DDALS(V3) 29.1$\pm$2.4 2022 99.9$\pm$0.2 91.3$\pm$2.9 0 19.9$\pm$1.7 1589 99.7$\pm$0.3 92.0$\pm$5.3 0
-ls2-Greedy 85.7$\pm$0.0 3.0$\pm$0.0 99.9$\pm$0.0 99.9$\pm$0.0 1 85.7$\pm$0.0 3.0$\pm$0.0 100.0$\pm$0.0 100.0$\pm$0.0 1
-ls2-GA 76.9$\pm$ 2.7 22617.4$\pm$10.9 99.0$\pm$0.1 98.9$\pm$0.2 1 61.2$\pm$2.1 21327.1$\pm$20.3 99.1$\pm$0.0 99.1$\pm$0.1 1
-ls2-EDA 75.9$\pm$0.0 22604.3$\pm$3.3 99.0$\pm$0.0 98.9$\pm$0.0 1 59.2$\pm$0.3 21306.5$\pm$2.5 99.2$\pm$0.0 99.2$\pm$0.1 1
-ls2-DDALS(O) 71.9$\pm$0.3 22591.3$\pm$63.6 99.0$\pm$0.0 99.0$\pm$0.0 1 57.2$\pm$0.0 21233.5$\pm$51.9 99.3$\pm$0.1 99.3$\pm$0.1 1
-ls2-DDALS(V1) 76.6$\pm$1.1 22591.3$\pm$63.6 99.6$\pm$0.1 99.5$\pm$0.1 1 58.5$\pm$0.8 21233.5$\pm$51.9 99.6$\pm$0.1 99.6$\pm$0.1 1
-ls2-DDALS(V2) 72.9$\pm$0.6 22591.3$\pm$63.6 99.3$\pm$0.2 99.2$\pm$0.2 1 57.5$\pm$0.2 21233.5$\pm$51.9 99.4$\pm$0.1 99.5$\pm$0.1 1
-ls2-DDALS(V3) 73.1$\pm$0.7 22564 99.2$\pm$0.1 99.1$\pm$0.1 1 57.9$\pm$0.4 21245 99.2$\pm$0.1 99.2$\pm$0.2 1
-ls4-Greedy 207.3$\pm$0.0 13.0$\pm$0.0 99.0$\pm$0.0 98.9$\pm$0.0 1 195.3$\pm$0.0 17.0$\pm$0.0 99.4$\pm$0.0 99.4$\pm$0.0 1
-ls4-GA 208.5$\pm$8.3 48014.1$\pm$12.1 99.0$\pm$0.0 98.6$\pm$0.3 0.7 175.2$\pm$10.3 45014.6$\pm$10.2 99.0$\pm$0.0 98.5$\pm$0.2 0.6
-ls4-EDA 205.3$\pm$2.4 48004.5$\pm$3.3 99.0$\pm$0.0 99.0$\pm$0.0 1 163.8$\pm$1.3 45005.5$\pm$2.6 99.1$\pm$0.0 98.8$\pm$0.0 1
-ls4-DDALS(O) 194.9$\pm$0.8 47962.8$\pm$67.7 99.0$\pm$0.0 98.9$\pm$0.0 1 163.3$\pm$3.6 44822.1$\pm$583.5 99.0$\pm$0.0 98.4$\pm$0.1 0.3
-ls4-DDALS(V1) 203.2$\pm$2.0 47962.8$\pm$67.7 99.1$\pm$0.0 99.0$\pm$0.1 1 172.8$\pm$4.5 44822.1$\pm$583.5 99.3$\pm$0.1 98.9$\pm$0.3 0.8
-ls4-DDALS(V2) 196.5$\pm$1.3 47962.8$\pm$67.7 99.0$\pm$0.0 99.0$\pm$0.0 1 165.3$\pm$4.3 44822.1$\pm$583.5 99.1$\pm$0.1 98.5$\pm$0.2 0.3
-ls4-DDALS(V3) 196.9$\pm$2.1 48010 99.0$\pm$0.0 98.9$\pm$0.1 1 168.2$\pm$ 2.3 44174 99.0$\pm$0.0 98.9$\pm$0.1 1
-In the case of LAB scenarios, the results of ss3, ss4 and ls4 are shown in Table V.
-We can intuitively see that the DDALS algorithm outputs the lowest cost solution across all test benchmarks, while Greedy needs only less than 10 times of evaluation but results in the highest cost.
-The feasibility ratio of all algorithms is close to 1, indicating that our solution evaluation algorithm only needs a very few samples to obtain an evaluation that is close to the true confidence level for small-scale scenarios and 500 samples for large-scale scenarios.
-From another perspective, for other variations, the original version of DDALS can obtain solutions with lower cost than other variations while guaranteeing a feasible ratio of 1, indicating that the solution evaluation is accurate.
-There are significant changes in the scenario of APP, shown in Table V.
-In small-scale benchmark tests, only the Greedy achieves a feasibility ratio of 1, while the feasibility ratio of other algorithms is almost zero.
-Although DDALS can obtain lower cost solutions, all of them are found to be infeasible.
-Such results are practically useless.
-The main reason is due to the retransmission mechanism in APP, which introduces discontinuities in the item distribution that pose a significant challenge to the solution evaluation algorithm.
-In a random sampling case, small samples may easily miss information regarding exceptional jumps, leading to performance decline of solution evaluation.
-Greedy can achieve higher feasibility ratios, but it sacrifices the optimality of the solutions.
-For the results of large-scale problems, although the difficulty of search increased with the increase of scale, DDALS can still find feasible solutions achieving the lowest cost.
-Note that in the result of ls4 with large $T_{\max}$, the feasibility ratio of DDALS(O) is 0.3 while that of DDALS(V3) is 1.
-Given that the average quality of the solutions generated by DDALS(V3) decreases by only 3% relative to the original DDALS, the feasibility ratio increased by 233%.
-It shows the effectiveness and necessity of filtering the solutions in complex scenarios.
-Moreover, increasing the sample size $L$ can improve the evaluation performance, which can help capture more information and improve the accuracy of solution evaluation.
+VI-B Comparison of DDALS with GDSO and DRO
+We adapt two existing algorithms, named GDSO [19] and DRO [13] respectively, to solve our problem.
+GDSO assumes that the distributions of random variable follow Gaussian distributions.
+Thus, we apply a statistical fitting method to sample data to obtain the mean and variance of item weights.
+DRO does not make any assumption about the distributions of random variables.
+Instead, DRO constructs ambiguity sets (such as ellipsoid sets) based on sample data and transform the problem into a deterministic mixed integer linear programming which can be solved by corresponding solvers.
+Details of the adaptation of these methods are available in the online supplementary.
+In the first comparison shown in Table IV, we find that DRO can not find feasible solutions in the LAB-ss2 instance.
+GDSO also fail to find feasible solutions when $T_{max}=14$.
+In the second comparison in Table V, we increase the sample size to 500.
+The result reveals that GDSO could solve the problem normally, while DRO again fail to find feasible solutions when $T_{max}$ is less than or equal to 20.
+In the third comparison in Table VI, we change the distributions of the random variables to be uniform distribution, keeping sample size 500.
+GDSO starts to fail to find feasible solutions when $T_{max}$ is less than 35, while DRO can not find any feasible solutions.
+Actually, the difference between the uniform distribution and the Gaussian distribution is greater than the difference between the mixed distribution and the Gaussian distribution in the LAB.
+Under the same intensity of chance constraints, the performance of GDSO gets worse.
+Although GDSO and DRO have the advantage of short solution times for small-scale problems, as the problem scale increases, their running times show explosive growth.
+The time consumed when the problem scale is greater than $10\times 20$ far exceeds the running time of DDALS, as shown in Table VII.
+TABLE IV: DDALS vs GDSO vs DRO comparison 1
+$T_{\max}$ Brute Force DDALS GDSO DRO
+Time(s) Cost Time(s) Cost Time(s) Cost Time(s) Cost
+27 3.81 8.07 0.99 8.07 0.03 8.07 - fail
+23 3.86 10.05 1.19 10.05 0.06 10.05 - fail
+18 3.94 10.36 1.69 10.36 0.04 10.36 - fail
+16 3.73 14 1.83 14 0.06 14 - fail
+14 3.60 19.41 2.1 19.41 - fail - fail
+TABLE V: DDALS vs GDSO vs DRO comparison 2
+$T_{\max}$ Brute Force DDALS GDSO DRO
+Time(s) Cost Time(s) Cost Time(s) Cost Time(s) Cost
+30 3.46 11.55 0.99 11.55 0.13 11.55 36.5 15.4
+28 3.44 11.55 1.1 11.55 0.13 11.55 22.4 27.1
+20 3.44 15.4 1.7 15.4 0.21 15.4 - fail
+15 3.40 18.91 2.2 18.91 0.16 18.91 - fail
+10 3.49 26.32 2.6 26.32 0.16 26.32 - fail
+TABLE VI: DDALS vs GDSO vs DRO comparison 3
+$T_{\max}$ Brute Force DDALS GDSO DRO
+Time(s) Cost Time(s) Cost Time(s) Cost Time(s) Cost
+37 3.44 4.17 1.0 4.17 0.13 4.17 - fail
+36 3.53 4.17 1.0 4.17 0.13 4.17 - fail
+35.5 3.49 4.17 1.1 4.17 0.13 4.17 - fail
+34.5 3.48 4.35 1.7 4.35 - fail - fail
+TABLE VII: DDALS vs GDSO vs DRO comparison 4
+Instance DDALS GDSO DRO
+Time(s) Cost Time(s) Cost Time(s) Cost
+LAB-ss1 0.75 5.91 0.07 5.91 0.35 7.1
+LAB-ss2 0.82 8.07 0.07 8.07 0.31 8.07
+LAB-ss3 1.10 10.06 0.08 10.06 0.42 10.06
+LAB-ss4 1.36 9.22 0.09 9.22 1.43 10.47
+LAB-ls1 7.73 26.65 2.13 26.68 $>$3600
+LAB-ls2 7.95 17.3 48.37 17.3 $>$3600
+LAB-ls3 60.98 55.67 $>$3600 $>$3600
+LAB-ls4 194.41 97.03 $>$3600 $>$3600
+The comparison results are summarized as follows:
+  • GDSO fails to find any feasible solution under high intensity of chance constraint.
+Its runtime increases rapidly with the growth of problem scale.
+As the difference between the real distribution and the Gaussian distribution increases, its performance gets worse.
+  • DRO fails to find feasible solutions with insufficient sample size ($\leq 500$) and high intensity of chance constraint. Its runtime exponentially increases with problem scale.
+  • Our proposed algorithm, DDALS, proves effective across all these scenarios.
+VI-C Testing Results on LAB and APP
+To evaluate the effectiveness of DDALS, we conduct experiments to compare and analyze various commonly used search methods, named Greedy, GA, EDA, each combined with evaluation approaches.
+Greedy is particularly effective in scenarios where locally optimal choices contribute to globally optimal solutions.
+GA demonstrate proficiency across various combinatorial optimization problems [38, 39, 40], and EDA exhibits great success in both combinatorial and continuous optimization [41].
+For equitable comparisons, we initiate the execution of DDALS, recording its evaluation times.
+Subsequently, Greedy, GA, and EDA start execution until convergence or their respective ET surpasses that of DDALS.
+The population sizes of GA and EDA are both set to 10, with 6 dominant individuals.
+The crossover probability for parents of GA is set to 0.1 and a variation probability of $1/m$.
+The real confidence level (RCL) is computed by generating $710e$ 7 data using the real generators.
+In our experiments, $P_{0}$ is set to 0.99.
+A solution is considered feasible if its real confidence level exceeds $0.985$, allowing for a 0.5% error margin.
+TABLE VIII: Performance on the instances of LAB. The highest average real confidence level and the lowest cost are highlighted through bolding and being underlined.
+Instance small $T_{\max}$ large $T_{\max}$
+C ET ECL RCL FSR Time C ET ECL RCL FSR Time
+ss3-Greedy 21.4$\pm$0.0 3.0$\pm$0.0 99.9$\pm$0.0 99.5$\pm$0.0 1 0.2$\pm$0.0 17.4$\pm$0.0 4.0$\pm$0.0 100.0$\pm$0.0 100.0$\pm$0.0 1 0.4$\pm$0.0
+ss3-GA 22.4$\pm$1.4 306.0$\pm$3.7 99.9$\pm$0.0 99.7$\pm$0.2 1 31.7$\pm$1.9 14.2$\pm$2.4 263.5$\pm$2.9 99.7$\pm$0.4 99.2$\pm$1.3 0.73 37.9$\pm$1.1
+ss3-EDA 24.8$\pm$1.9 305.3$\pm$3.0 99.9$\pm$0.1 99.8$\pm$0.5 0.93 44.1$\pm$3.0 13.7$\pm$1.3 264.5$\pm$2.8 99.9$\pm$0.0 99.9$\pm$0.1 1 42.8$\pm$3.4
+ss3-DDALS(O) 21.4$\pm$0.0 288.5$\pm$9.3 99.9$\pm$0.0 99.5$\pm$0.0 1 9.6$\pm$0.5 12.6$\pm$0.0 254.4$\pm$4.7 100.0$\pm$0.0 99.9$\pm$0.0 1 14.7$\pm$0.9
+ss3-DDALS(V1) 21.4$\pm$0.0 288.5$\pm$9.3 99.9$\pm$0.0 99.5$\pm$0.0 1 9.6$\pm$0.5 12.6$\pm$0.0 254.4$\pm$4.7 100.0$\pm$0.0 99.9$\pm$0.0 1 14.7$\pm$0.9
+ss3-DDALS(V2) 21.4$\pm$0.0 288.5$\pm$9.3 99.9$\pm$0.0 99.5$\pm$0.0 1 9.6$\pm$0.5 12.6$\pm$0.0 254.4$\pm$4.7 100.0$\pm$0.0 99.9$\pm$0.0 1 14.7$\pm$0.9
+ss3-DDALS(V3) 23.9$\pm$1.7 284 99.8$\pm$0.2 99.1$\pm$0.8 0.75 9.1 15.5$\pm$1.7 264 99.9$\pm$0.3 99.5$\pm$0.9 0.9 14.9
+ss4-Greedy 30.1$\pm$0.0 2.0$\pm$0.0 99.9$\pm$0.0 99.9$\pm$0.0 1 0.1$\pm$0.0 14.9$\pm$0.0 5.0$\pm$0.0 99.9$\pm$0.0 99.9$\pm$0.0 1 0.3$\pm$0.0
+ss4-GA 23.1$\pm$4.4 773.5$\pm$8.0 99.9$\pm$0.2 99.7$\pm$0.4 1 65.6$\pm$2.5 16.4$\pm$3.0 559.3$\pm$8.8 99.8$\pm$0.3 99.4$\pm$0.6 0.9 60.9$\pm$3.1
+ss4-EDA 17.4$\pm$0.0 764.5$\pm$3.0 99.9$\pm$0.0 99.9$\pm$0.0 1 98.8$\pm$1.7 13.4$\pm$1.0 554.3$\pm$2.6 99.8$\pm$0.2 99.1$\pm$0.41 0.93 83.5$\pm$3.3
+ss4-DDALS(O) 17.4$\pm$0.0 757.9$\pm$31.9 99.9$\pm$0.0 99.9$\pm$0.0 1 21.4$\pm$1.0 11.2$\pm$0.0 538.8$\pm$31.9 99.2$\pm$0.0 99.0$\pm$0.0 1 23.6$\pm$1.1
+ss4-DDALS(V1) 17.4$\pm$0.0 757.9$\pm$31.9 99.9$\pm$0.0 99.9$\pm$0.0 1 21.4$\pm$1.0 12.3$\pm$0.7 538.8$\pm$31.9 99.6$\pm$0.1 98.9$\pm$0.2 1 23.6$\pm$1.1
+ss4-DDALS(V2) 17.4$\pm$0.0 757.9$\pm$31.9 99.9$\pm$0.0 99.9$\pm$0.0 1 21.4$\pm$1.0 14.1$\pm$0.9 538.8$\pm$31.9 99.9$\pm$0.2 99.0$\pm$0.1 1 23.6$\pm$1.1
+ss4-DDALS(V3) 25.3$\pm$4.1 822 99.8$\pm$0.2 99.6$\pm$0.5 1 22.9 14.3$\pm$1.5 514 99.6$\pm$0.3 99.2$\pm$0.7 0.9 24.1
+ls4-Greedy 167.5$\pm$0.0 7.0$\pm$0.0 99.2$\pm$0.0 99.1$\pm$0.0 1 3.5$\pm$0.0 167.6$\pm$0.0 8.0$\pm$0.0 99.9$\pm$0.0 99.9$\pm$0.0 1 3.7$\pm$0.2
+ls4-GA 124.1$\pm$8.7 11309.8$\pm$7.4 99.1$\pm$0.2 99.1$\pm$0.3 0.97 12310.4$\pm$282.8 107.5$\pm$6.8 8411.4$\pm$9.8 99.1$\pm$0.1 99.0$\pm$0.3 0.93 8437.8$\pm$157.5
+ls4-EDA 123.6$\pm$3.6 11305.1$\pm$3.4 99.3$\pm$0.2 99.4$\pm$0.2 1 9498.8$\pm$243.3 116.6$\pm$3.2 8403.9$\pm$3.1 99.6$\pm$0.2 99.7$\pm$0.2 1 7970.2$\pm$162.7
+ls4-DDALS(O) 107.1$\pm$1.0 11360.9$\pm$206.6 99.1$\pm$0.1 99.2$\pm$0.1 1 5537.5$\pm$95.9 92.6$\pm$1.1 8483.8$\pm$335.8 99.1$\pm$0.1 99.0$\pm$0.3 0.93 4113.9$\pm$147.9
+ls4-DDALS(V1) 109.5$\pm$0.5 11360.9$\pm$206.6 99.7$\pm$0.1 99.7$\pm$0.1 1 5537.5$\pm$95.9 98.8$\pm$2.1 8483.8$\pm$335.8 99.6$\pm$0.1 99.6$\pm$0.1 1 4113.9$\pm$147.9
+ls4-DDALS(V2) 108.9$\pm$1.1 11360.9$\pm$206.6 99.6$\pm$0.2 99.6$\pm$0.2 1 5537.5$\pm$95.9 94.4$\pm$2.0 8483.8$\pm$335.8 99.4$\pm$0.2 99.3$\pm$0.3 1 4113.9$\pm$147.9
+ls4-DDALS(V3) 110.8$\pm$3.5 11576 99.3$\pm$0.2 99.4$\pm$0.2 1 5651.3 95.1$\pm$2.5 7964 99.1$\pm$0.1 99.3$\pm$0.1 1 3883.5
+All presented results are derived from a statistical analysis conducted with 30 repetitions, denoted as “mean $\pm$ standard deviation”.
+Due to space constraints, selected results are included in the main text, while the complete dataset is available in the online supplementary material.
+Take Table VIII as an example.
+“ss3-Greedy” represents the results of the Greedy algorithm applied to the LAB-ss3 instance.
+Variants of the DDALS algorithm are denoted as “DDALS(O)”, “DDALS(V1)”, “DDALS(V2)”, and “DDALS(V3)” for the original, first, second, and third versions, respectively.
+Categories such as “small $T_{\max}$” and “large $T_{\max}$” correspond to the values of $T_{\max}$ specified in Table I.
+For example, in experiments on LAB-ss1, “small $T_{\max}$” is set to 11, and “large $T_{\max}$” is set to 14, corresponding to ${11,14}$ in Table I.
+Five indicators are utilized to measure the performance: “C”represents the cost of the solution, “ET” measures the evaluation times for comparing convergence speeds, “ECL” signifies the estimated confidence level, “RCL” is real confidence level, and “FSR” denotes the ratio of actual feasible solutions in the results of all experiments.
+“Time” represents the actual runtime of the algorithm on the server, measured in seconds.
+TABLE IX: Performance on the instances of APP. Due to the introduction of retransmission mechanisms, there is a significant difference in the results between LAB and APP.
+Instance small $T_{\max}$ large $T_{\max}$
+C ET ECL RCL FSR Time C ET ECL RCL FSR Time
+ss2-Greedy 5.9$\pm$0.0 6.0$\pm$0.0 99.8$\pm$0.0 98.5$\pm$0.0 0 12.6$\pm$0.4 5.9$\pm$0.0 6.0$\pm$0.0 99.9$\pm$0.0 98.9$\pm$0.0 1 8.7$\pm$0.2
+ss2-GA 5.3$\pm$0.1 303.2$\pm$2.3 99.2$\pm$0.0 97.3$\pm$0.5 0 4508.8$\pm$1198.4 5.3$\pm$0.1 272.9$\pm$2.2 99.5$\pm$0.1 98.1$\pm$0.2 0.03 3422.4$\pm$352.0
+ss2-EDA 5.5$\pm$0.3 305.0$\pm$2.7 99.2$\pm$0.2 97.6$\pm$0.7 0.1 3323.2$\pm$1532.5 5.6$\pm$0.3 274.1$\pm$2.4 99.5$\pm$0.2 98.5$\pm$0.5 0.47 1185.3$\pm$1048.5
+ss2-DDALS(O) 5.2$\pm$0.0 300.0$\pm$10.9 99.2$\pm$0.0 97.3$\pm$0.0 0 1010.5$\pm$98.7 5.2$\pm$0.0 297.5$\pm$7.8 99.6$\pm$0.0 98.1$\pm$0.0 0 862.2$\pm$78.7
+ss2-DDALS(V1) 5.4$\pm$0.0 300.0$\pm$10.9 99.8$\pm$0.1 97.4$\pm$0.2 0 1010.5$\pm$98.7 5.2$\pm$0.0 297.5$\pm$7.8 99.6$\pm$0.0 98.1$\pm$0.0 0 862.2$\pm$78.7
+ss2-DDALS(V2) 5.4$\pm$0.0 300.0$\pm$10.9 99.8$\pm$0.0 97.4$\pm$0.2 0 1010.5$\pm$98.7 5.4$\pm$0.1 297$\pm$7.8 99.9$\pm$0.1 98.2$\pm$0.2 0.1 862.2$\pm$78.7
+ss2-DDALS(V3) 5.4$\pm$0.1 306 99.5$\pm$0.3 97.6$\pm$0.4 0 1086.0 5.5$\pm$0.1 298 99.7$\pm$0.2 98.5$\pm$0.3 0.5 858.0
+ss4-Greedy 34.9$\pm$0.0 3.0$\pm$0.0 100.0$\pm$0.0 100.0$\pm$0.0 1 0.2$\pm$0.0 28.5$\pm$0.0 4.0$\pm$0.0 100.0$\pm$0.0 99.0$\pm$0.0 1 0.3$\pm$0.0
+ss4-GA 27.5$\pm$1.4 2008.9$\pm$7.9 99.8$\pm$0.2 90.3$\pm$0.5 0 251.8$\pm$24.5 20.5$\pm$2.7 1608.4$\pm$7.4 99.4$\pm$0.3 92.9$\pm$4.8 0 322.5$\pm$9.2
+ss4-EDA 26.1$\pm$0.0 2005.7$\pm$2.3 100.0$\pm$0.0 90.1$\pm$0.0 0 333.9$\pm$11.8 18.3$\pm$2.7 1604.5$\pm$2.8 99.9$\pm$0.1 85.2$\pm$6.9 0 287.9$\pm$19.6
+ss4-DDALS(O) 26.1$\pm$0.0 1988.9$\pm$27.8 100.0$\pm$0.0 90.1$\pm$0.0 0 33.3$\pm$2.3 16.6$\pm$0.8 1582.8$\pm$54.4 99.8$\pm$0.1 81.3$\pm$3.7 0 61.3$\pm$3.9
+ss4-DDALS(V1) 26.1$\pm$0.0 1988.9$\pm$27.8 100.0$\pm$0.0 90.1$\pm$0.0 0 33.3$\pm$2.3 16.8$\pm$1.1 1582.8$\pm$54.4 99.8$\pm$0.1 81.8$\pm$4.3 0 61.3$\pm$3.9
+ss4-DDALS(V2) 26.1$\pm$0.0 1988.9$\pm$27.8 100.0$\pm$0.0 90.1$\pm$0.0 0 33.3$\pm$2.3 18.3$\pm$2.2 1582.8$\pm$54.4 99.9$\pm$0.1 86.9$\pm$7.3 0 61.3$\pm$3.9
+ss4-DDALS(V3) 29.7$\pm$2.5 1944 99.9$\pm$0.2 92.2$\pm$3.9 0.2 28.3 20.8$\pm$2.9 1586 99.7$\pm$0.3 88.3$\pm$6.3 0 59.3
+ls2-Greedy 89.1$\pm$0.0 2.0$\pm$0.0 99.9$\pm$0.0 99.9$\pm$0.0 1 0.2$\pm$0.0 89.1$\pm$0.0 2.0$\pm$0.0 100.0$\pm$0.0 100.0$\pm$0.0 1 0.3$\pm$0.
+ls2-GA 76.7$\pm$2.8 22616.5$\pm$10.5 99.1$\pm$0.1 99.0$\pm$0.2 1 3925.1$\pm$155.5 60.9$\pm$2.7 21324.5$\pm$13.3 99.1$\pm$0.1 99.1$\pm$0.1 1 3382.2$\pm$64.9
+ls2-EDA 79.3$\pm$1.1 22604.1$\pm$2.6 99.5$\pm$0.1 99.3$\pm$0.1 1 7174.3$\pm$44.2 59.9$\pm$0.8 21304.8$\pm$3.2 99.2$\pm$0.0 99.2$\pm$0.0 1 6970.8$\pm$88.9
+ls2-DDALS(O) 72.0$\pm$0.3 22631.7$\pm$62.4 99.0$\pm$0.0 99.0$\pm$0.0 1 841.3$\pm$7.6 57.2$\pm$0.0 21265.2$\pm$73.3 99.3$\pm$0.1 99.3$\pm$0.0 1 776.4$\pm$6.6
+ls2-DDALS(V1) 76.2$\pm$1.1 22631.7$\pm$62.4 99.5$\pm$0.1 99.5$\pm$0.1 1 841.3$\pm$7.6 59.0$\pm$1.2 21265.2$\pm$73.3 99.6$\pm$0.0 99.6$\pm$0.1 1 776.4$\pm$6.6
+ls2-DDALS(V2) 73.2$\pm$0.4 22631.7$\pm$62.4 99.4$\pm$0.1 99.3$\pm$0.1 1 841.3$\pm$7.6 57.6$\pm$0.5 21265.2$\pm$73.3 99.4$\pm$0.1 99.5$\pm$0.1 1 776.4$\pm$6.6
+ls2-DDALS(V3) 73.7$\pm$1.5 22536 99.2$\pm$0.1 99.0$\pm$0.2 1 846.2 58.3$\pm$0.6 21274 99.3$\pm$0.1 99.3$\pm$0.1 1 769.6
+ls4-Greedy 238.1$\pm$0.0 3.0$\pm$0.0 99.4$\pm$0.0 99.0$\pm$0.0 1 1.5$\pm$0.0 193.7$\pm$0.0 15.0$\pm$0.0 99.2$\pm$0.0 98.8$\pm$0.0 1 7.0$\pm$0.2
+ls4-GA 206.6$\pm$4.7 48018.6$\pm$13.4 99.0$\pm$0.0 98.6$\pm$0.3 0.73 46617.9$\pm$648.4 182.7$\pm$10.7 45015.2$\pm$11.7 99.0$\pm$0.0 98.5$\pm$0.2 0.6 42937.2$\pm$431.6
+ls4-EDA 204.3$\pm$1.9 48004.5$\pm$3.5 99.0$\pm$0.0 99.0$\pm$0.0 1 45488.7$\pm$924.8 163.6$\pm$1.2 45003.5$\pm$2.4 99.1$\pm$0.1 98.8$\pm$0.1 1 42955.7$\pm$613.1
+ls4-DDALS(O) 195.1$\pm$0.5 47999.9$\pm$83.0 99.0$\pm$0.0 98.9$\pm$0.0 1 16357.5$\pm$137.6 161.5$\pm$2.1 44440.5$\pm$344.6 99.0$\pm$0.0 98.5$\pm$0.3 0.4 15160.3$\pm$127.4
+ls4-DDALS(V1) 209.2$\pm$2.8 47999.9$\pm$83.0 99.1$\pm$0.0 98.6$\pm$0.3 0.6 16357.5$\pm$137.6 169.4$\pm$4.2 44440.5$\pm$344.6 99.3$\pm$0.2 98.9$\pm$0.4 0.8 15160.3$\pm$127.4
+ls4-DDALS(V2) 197.4$\pm$1.7 47999.9$\pm$83.0 99.1$\pm$0.0 99.0$\pm$0.1 1 16357.5$\pm$137.6 162.5$\pm$2.4 44440.5$\pm$344.6 99.1$\pm$0.1 98.6$\pm$0.2 0.7 15160.3$\pm$127.4
+ls4-DDALS(V3) 196.8$\pm$0.8 48046 99.0$\pm$0.0 98.9$\pm$0.1 1 16280.8 163.5$\pm$3.5 44648 99.0$\pm$0.0 98.3$\pm$0.1 0.1 15162.6
+Table VIII shows the results of ss3, ss4 and ls4 in the LAB.
+DDALS consistently obtains the lowest-cost solutions across all test instances.
+In contrast, Greedy requires fewer than 10 evaluations but yields the highest cost.
+The feasibility ratio for all algorithms is approximately 1, suggesting that our solution evaluation approach is able to achieve an evaluation close to the true confidence level.
+In comparison with other variations, the original version of DDALS consistently produces lower-cost solutions while maintaining a feasibility ratio of 1, demonstrating the accuracy of the solution evaluation approach.
+Table IX illustrates significant changes in the results of APP.
+In small-scale instance tests, only Greedy achieves a feasibility ratio of 1, while other algorithms nearly approach zero feasibility.
+Although DDALS obtain lower cost solutions than other algorithms, all of them are found to be infeasible.
+The primary challenge arises from the retransmission mechanism in APP, introducing discontinuities in item’s distribution that pose a considerable challenge to the solution evaluation approach.
+In cases of random sampling, small samples may overlook information about exceptional jumps, leading to a decline in solution evaluation performance.
+While Greedy achieves higher feasibility ratios, it sacrifices solution optimality.
+For large-scale problems, despite increased search difficulty with problem scale, DDALS consistently finds feasible solutions with the lowest cost.
+In the result of APP-ls4 with large $T_{\max}$, the feasibility ratio of DDALS(O) is 0.4, while that of DDALS(V1) is 0.8.
+Despite a 4.9% decrease in the average quality of solutions generated by DDALS(V1) relative to the original DDALS, the feasibility ratio increase by 100%.
+This emphasizes the effectiveness and necessity of filtering solutions in practical scenarios.
+Additionally, increasing the sample size $L$ is able to effectively enhances evaluation performance, capturing more information and improving the accuracy of solution evaluation approach.
+The results reveal that the runtime of DDALS is significantly shorter than both EDA and GA.
 VI-D Effectiveness of Components in DDALS
 We conduct an ablation study to further assess the effectiveness of four components of DDALS, including CP, LSS, Degrade and FSS.
-To independently assess their effectiveness, four different variants of DDALS are proposed and tested on benchmarks of LAB:
+Regarding the SFE, we can observe significant distinctions in the results between DDALS(O) and other variants in Table VIII and Table IX.
+To independently assess their effectiveness, four different variants of the original DDALS are tested on the instances of Lab:
   1. 1.
-    r-CP DDALS: It uses random strategy to select the item in each class instead of utilizing CP.
-Note that if the initial randomly generated is infeasible solution, query the current solution for each class of item and re-generate randomly in each class of items less than the weight of the current item.
+    r-CP DDALS: It uses random strategy for solution initialization instead of CP.
   2. 2.
-    no-LSS DDALS: The LSS procedure in the iteration process is removed from DDALS.
+    no-LSS DDALS: LSS is removed.
   3. 3.
-    no-Degrade DDALS: The degrade procedure in the iteration process is removed.
+    no-Degrade DDALS: Degrade is removed.
   4. 4.
-    no-FSS DDALS: The FSS procedure is removed.
-For all the above variants, we compute the performance degradation ratio(PDR) by computing $(C_{2}-C_{1})/C1$ for each benchmark, where $C_{1}$ and $C_{2}$ are the average cost of solution obtained by the variant and original across 10 independent runs.
-The component being examined is useful for a specific LAB benchmark if the result shows that PDR $>0$.
-Table VII presents the average PDR averaged on each benchmark of LAB.
-Based on the results, it can be concluded that none of the examined components has a negative effect on the performance of DDALS.
-In most cases, these components make LS perform better.
-Constructing utility for items is helpful for DDALS to obtain a good initial solution on average.
-LSS provides some kind of exploration, but a negative value on ls3 implies that due to the large size of the sample, the algorithm may still not converge after 30 iterations of running.
-The operation of degradation has the most significant contribution to the performance of LS, since it greatly improves the problem of the algorithm, as it helps the algorithm to escape local optima.
-A deeper search is necessary, while FSS has a slightly higher PDR than LSS.
-However, with the increase of the benchmark’s scale, the contribution of FSS seems to decrease, implying that more efficient and exploratory methods are needed to cope with larger-scale problems.
-TABLE VII: Average PDR on each benchmark of LAB. The highest PDR is highlighted through bolding and being underlined.
-Benchmark r-CP no-LSS no-Degrade no-FSS
+    no-FSS DDALS: FSS is removed.
+Table X presents the results.
+We compute the performance degradation ratio (PDR) by computing $(C_{2}-C_{1})/C1$ for each instance, where $C_{1}$ and $C_{2}$ are the average cost of solution obtained by the variants and original DDALS across 10 independent runs, respectively.
+The effect of the component is considered positive for a specific LAB instance when the PDR is greater than 0.
+The results indicate that none of the examined components adversely impacts the performance of DDALS.
+In most instances, these components enhance the performance of DDALS.
+Constructing utility for items proves beneficial for DDALS in achieving a favorable initial solution on average.
+While the purpose of LSS is to explore the solution space, a negative value on ls3 suggests that, due to large sample size, the algorithm may not converge within 30 iterations.
+Degrade contributes significantly to the performance of DDALS, addressing the problem by aiding its escape from local optima.
+FSS exhibits a slightly higher PDR than LSS.
+However, with the increase in the problem scale, the impact of FSS appears to diminish, indicating the need for more efficient and exploratory methods to address large-scale problems.
+TABLE X: Average PDR on each instance of LAB. The highest PDR is highlighted through bolding and being underlined.
+Instance r-CP no-LSS no-Degrade no-FSS
 $ss1$ 0.00% 0.00% 0.00% 0.00%
@@
 VI-E Acceleration Performance of Accelerated Monte Carlo
-The acceleration effect of Accelerated Monte Carlo (AMC) evaluation is validated.
-We randomly selected 10,000 solutions for evaluation, using regular Monte Carlo (MC) and AMC respectively.
-After recording the running time required for each 1000 evaluations, the result is shown in Figure 6.
-The results show that the time consumed increases essentially linearly as the number of evaluation times increases.
-For different $T_{\max}$ settings, MC takes essentially the same amount of time since the number of simulations needed does not vary with $T_{\max}$, while AMC can obtain better acceleration compared to MC when $T_{\max}$ is small.
-For example, compared with the running time $526.601s$ of MC, AMC takes $458.449s$ to finish the evaluation when $T_{\max}=11$, while the number of fast-filtered solution is 10000.
-When $T_{\max}=23$, it takes AMC about $471.395s$, although the number of pre-fast filtered is 9718.
-As $T_{\max}$ increases, the difficulty of the chance constraint decreases, resulting in the increase of the percentage of feasible solutions in the whole solution space.
-Thus, the number of solutions that can be fast pre-filtered by sequences decreases, which leads to a decrease in the acceleration effect.
-The accelerated mechanism may have a negative impact on the overall efficiency of the evaluation as $T_{\max}$ becomes large.
-When $T_{\max}=27$, AMC and MC take almost the same time, while when $T_{\max}=35$, AMC needs to take $807.435s$ to finish rather than about $500s$ of MC.
-In conclusion, when applying the accelerated mechanism, a pre-assessment of whether the constraints are tight enough is necessary to determine whether the constraints are tight enough.
-Figure 6: Acceleration performance of accelerated Monte Carlo on LAB-ls1.
-The mechanism has a positive effect when the constraints are tight, but has a negative impact on the overall efficiency as $T_{\max}$ becomes large.
-VI-F Performance of Other Works
-We also attempted to adjust two similar algorithms, Ji’s [15] and Yang’s [34], to solve our problem (see Appendix 4 for modification details).
-Due to the incapability of dynamic programming and mixed integer programming for solving large-scale DDCCMCKP, we conducted experiments on small-scale benchmarks to evaluate their performance.
-Results are shown in Table VIII, and we observed that the majority of the results were either ”infeasible” or ”fail”.
-”fail” means that the algorithms fail to output any solution, while ”infeasible” means that the optimal solution is infeasible.
-The value of ”Feasibility” column is the total cost of solution.
-For Yang’s method, the poor performance was mainly due to the significant difference between the actual distribution and the Gaussian distribution, and the insufficient data to characterize an unknown distribution using sample mean and variance, leading to incorrect iteration.
-In Ji’s work, a uniform distribution was used and tested.
-The performance of their method is poor since it could not handle the long-tailed distribution of the candidate solutions in our problem and led to suboptimal performance.
-TABLE VIII: Result of Ji’s [15] and Yang’s [34]
-Benchmark Ji Yang
-Time(s) Feasiblility Time(s) Feasiblility
-LAB-$ss1$ fail 1.0 14.5
-LAB-$ss2$ 0.4 infeasible - fail
-LAB-$ss3$ 0.8 infeasible 1.4 21.5
-LAB-$ss4$ 16 infeasible 43.2 17.4
-APP-$ss1$ fail 0.1 infeasible
-APP-$ss2$ 0.6 infeasible - fail
-APP-$ss3$ fail 0.2 infeasible
-APP-$ss4$ fail 13.5 infeasible
+The effectiveness of accelerated Monte Carlo (AMC) evaluation is confirmed through validation.
+We randomly selected 10,000 solutions of LAB-ls1 for assessment, employing both Monte Carlo (MC) and AMC methods.
+We record the runtime required for every 1000 evaluations, and the results are presented in Figure 6.
+Experiment results reveal a nearly linear increase in the consumed time with the growth in the number of evaluation times.
+For various $T_{\max}$ settings, MC exhibits nearly constant time consumption, since the required number of simulations remains consistent across different $T_{\max}$.
+In contrast, AMC demonstrates superior acceleration, particularly with smaller $T_{\max}$.
+For example, compared to the MC’s runtime of $526.601s$, AMC completes the evaluation in $458.449s$ when $T_{\max}=11$, processing a fast-filtered solution count of 10,000.
+At $T_{\max}=23$, AMC takes approximately $471.395s$, with a pre-fast-filtered count of 9,718.
+As $T_{\max}$ increases, the challenge posed by the intensity of chance constraint diminishes, resulting in a higher percentage of feasible solutions in the entire solution space.
+When $T_{\max}=27$, AMC and MC take nearly the same time, while at $T_{\max}=35$, AMC needs $807.435s$, compared to approximately $500s$ for MC.
+The results show that accelerated mechanism may have a negative impact on the overall efficiency when the intensity of chance constraint is low.
+In conclusion, when implementing the accelerated mechanism, it is necessary to pre-assess whether the intensity of chance constraints is high or not.
+Figure 6: Acceleration performance of AMC on LAB-ls1.
+The mechanism has a positive effect when the intensity of chance constraint is high, but has a negative impact on the overall efficiency as the intensity of chance constraint is lower, i.e., $T_{\max}$ becomes larger.
 VII Conclusion
-In this work, derived from practical applications, we study a new variant of MCKP called DDCCMCKP.
-To address this specific problem, a novel DDALS algorithm has been introduced.
-Based on local search framework, DDALS enables data-driven solution evaluation for the problem, and incorporates a surrogate weight to ensure the optimality of solutions.
+In this work, motivated by practical applications, we study a novel variant of MCKP named chance-constrained MCKP.
+To address CCMCKP, we propose a data-driven adaptive local search algorithm.
+DDALS facilitates the data-driven solution evaluation approach that can effectively handle situations where the distributions of item weights are unknown, and incorporates a surrogate weight mechanism to ensure solution optimality.
 Moreover, a solution filtering mechanism is introduced to improve the feasibility of the output solutions.
-Two sets of benchmarks for DDCCMCKP have been proposed, one of which simulates the real business requirement of 5G E2E network scenario for a communication company.
-The experimental results demonstrate that DDALS is able to solve DDCCMCKP effectively and outperforms other baseline methods.
-Although the APP benchmark imposes higher demands on problem-solving abilitiy, the performance of DDALS is not satisfactory in small-scale problems.
-However, DDALS can solve DDCCMCKP well in most of the scenarios and maintain better performance than other baselines.
-Overall, our proposed algorithm holds great promise for addressing this new variant of MCKP in the data-driven scenario.
-DDALS can serve as a baseline for future research.
-The code and benchmark sets will be open-sourced for further promoting research on DDCCMCKP.
-The algorithm framework can be further improved in two aspects.
-A new surrogate weight design may improve the efficiency and solution’s quality instead of the current design.
-From the other perspective, more effective filtering strategies that can balance the feasibility and optimality of the solutions could be explored in the future.
+Two benchmark sets for CCMCKP have been established, one of which simulates the real business requirement of a 5G E2E network scenario for a communication company.
+Experimental results demonstrate the superiority of DDALS over other baselines on the two benchmarks.
+In cases with high intensity of chance constraints and limited amounts of sample data, DDALS still exhibits good performance.
+Ablation studies also confirm the effectiveness of each component of the algorithm.
+Overall, our proposed algorithm holds great promise for addressing CCMCKP.
+Our proposed algorithm DDALS can serve as a baseline for future research.
+The source code of algorithm, benchmark sets and supplementary materials are available online for promoting further research on solving this problem.
+Furthermore, DDALS is not confined to CCMCKP but can be easily applied to other problems.
+Specifically, it is applicable to any chance-constrained problem that can be addressed via data-driven evaluation approach.
+While the algorithm framework is robust, it can be further improved in two aspects.
+A new surrogate weight design could enhance efficiency and solution quality beyond the current implementation.
+Additionally, it is necessary to further accelerate the evaluation process, since the evaluation of solutions is time-consuming due to extensive Monte Carlo simulations involved in each evaluation process.
 Appendix A Details about Concentration Inequalities
@@
 Figure 8: Sample size requirement with different error in theory
-The minimum sample sizes required for different estimation error are listed in Table IX.
+The minimum sample sizes required for different estimation error are listed in Table XI.
 Overall, Hoeffding provides the tightest bounds, especially suitable for larger errors.
@@
 Practically, it is verified that the amount of data required to meet the corresponding error requirement for the true confidence will be much smaller than the theoretical value obtained.
-TABLE IX: Relationship between EE and sample size
+TABLE XI: Relationship between EE and sample size
 […]
@@
 In this section, we show the complete experimental results of DDALS on LAB and APP.
-TABLE X: Performance on benchmarks of LAB. The highest average real confidence level and the lowest cost are highlighted through bolding and being underlined.
+TABLE XII: Performance on benchmarks of LAB. The highest average real confidence level and the lowest cost are highlighted through bolding and being underlined.
 Benchmark small $T_{\max}$ large $T_{\max}$
@@
 ls6-DDMLS(V3) 202.6$\pm$ 2.0 33681 99.1$\pm$0.1 99.1$\pm$0.1 1 165.0$\pm$2.6 19387 99.2$\pm$0.2 99.1$\pm$0.1 1
-TABLE XI: Performance on benchmarks of APP. Due to the introduction of retransmission mechanisms, there is a significant difference in the results between LAB and APP.
+TABLE XIII: Performance on benchmarks of APP. Due to the introduction of retransmission mechanisms, there is a significant difference in the results between LAB and APP.
 Benchmark small $T_{\max}$ large $T_{\max}$